### v1

# Model-Based Transfer Learning for Contextual Reinforcement Learning

Abstract
Deep reinforcement learning is a powerful approach to complex decision making. However, one issue that limits its practical application is its brittleness, sometimes failing to train in the presence of small changes in the environment. This work is motivated by the empirical observation that directly applying an already trained model to a related task often works remarkably well, also called zero-shot transfer. We take this practical trick one step further to consider how to systematically select good tasks to train, maximizing overall performance across a range of tasks. Given the high cost of training, it is critical to choose a small set of training tasks. The key idea behind our approach is to explicitly model the performance loss (generalization gap) incurred by transferring a trained model. We hence introduce Model-Based Transfer Learning (MBTL) for solving contextual RL problems. In this work, we model the performance loss as a simple linear function of task context similarity. Furthermore, we leverage Bayesian optimization techniques to efficiently model and estimate the unknown training performance of the task space. We theoretically show that the method exhibits regret that is sublinear in the number of training tasks and discuss conditions to further tighten regret bounds. We experimentally validate our methods using urban traffic and standard control benchmarks. Despite the conceptual simplicity, the experimental results suggest that MBTL can achieve greater performance than strong baselines, including exhaustive training on all tasks, multi-task training, and random selection of training tasks. This work lays the foundations for investigating explicit modeling of generalization, thereby enabling principled yet effective methods for contextual RL.
1 Introduction
Deep reinforcement learning (DRL) has made remarkable strides in addressing complex problems across various domains [24, 31, 1, 4, 10, 12, 23]. Despite these successes, DRL models often exhibit brittleness when exposed to small variations in task settings, significantly limiting their scalability and generalizability [28, 40]. Such variations in Markov Decision Processes (MDPs) can be conceptualized as contextual Markov Decision Processes (CMDP), where slight differences in MDP definitions lead to multiple, closely related but different environments [14, 27]. Traditional DRL approaches typically require independent training for each new task variant, a process that is both resource-intensive and impractical for real-time applications. Multi-task training is a relatively inexpensive way to train but still has limited model capacity issues. In light of this, there is a pressing need for more robust approaches that enhance generalization across different tasks. It is crucial to differentiate between training robustness, which seeks resilience of training across a variety of tasks, and model robustness, which focuses on the resistance of a model to environmental perturbations. Contextual reinforcement learning (CRL) offers a framework for modeling changes in tasks [6]. In our exploration of applying DRL to solve multiple related tasks, we identify computational challenges, particularly the high costs and complexities associated with training separate models for each task. While techniques like CARL [6] or multitask RL [36, 2] address these challenges, they still remain computationally expensive and inefficient when applied across a broad context range.
Figure 1: Generalization gap in Cartpole CMDPs. The solid lines illustrate the actual generalized performance after zero-shot transfer across contexts, with the source context indicated by a dotted line.
Figure 2: Model-Based Transfer Learning (MBTL) for solving contextual MDPs. MBTL framework strategically selects optimal tasks to train by modeling the generalization gap that arises when transferring a model to different tasks. For each target task, we deploy the most effective trained model, indicated by solid arrows. We evaluate our framework on a suite of continuous and discrete tasks, including standard control benchmarks [6] and urban traffic benchmarks [34, 19].
We build upon zero-shot transfer, which has shown potential in directly applying a policy trained in one MDP (source task) to another (target task) without adaptation. This often leads to a performance degradation known as the ‘generalization gap’ [16, 20]. For instance, Figure 1 exemplifies the concept of the generalization gap within a Cartpole CMDP, a range of scenarios in which the agent must balance a pole on a moving cart with a different mass of pole. The performance degrades as the target task diverges from the source task, illustrating an increasing generalization gap. Given that training is expensive yet zero-shot transfer is cheap, we are interested in optimally selecting a set of source (training) tasks, such that performance on the target range of tasks is maximized. In this article, we devise strategies to incorporate the structure of the generalization gap to better estimate the value of training. Specifically, we investigate how explicit modeling of the generalization gap structure lends itself to principled and effective algorithms for the selection of training tasks within CMDP.
Hence, this article introduces Model-Based Transfer Learning (MBTL), a novel framework for studying generalization in reinforcement learning (Figure 2). We model the explicit form of the generalization gap in zero-shot transfer and use it in estimating how valuable selecting the source task is. Given a small number of samples in earlier steps, we leverage Bayesian optimization methods to estimate training performance with limited information, thus significantly reducing the necessity for exhaustive training [35, 37].
This paper contributes to the field of reinforcement learning by introducing MBTL framework for efficient source task selection, aiming to solve multiple related tasks with enhanced transferability and fewer required models. Theoretical analyses detail the sublinear cumulative regret of these methods and provide conditions for achieving tighter regret bounds. We validate our methods in simulated traffic management scenarios and standard control benchmarks. Our findings provide foundations for investigating explicit modeling of generalization, enabling efficient and practical model-based solutions for solving multiple related tasks for robust training in DRL.
2 Preliminaries and notation
We leverage the framework of CMDPs [14, 27, 6] to formulate the problem of solving multiple related tasks using reinforcement learning.
Contextual MDP (CMDP). A standard MDP is defined by the tuple $M=(S,A,P,R,\rho)$ where $S$ represents the state space, $A$ is the action space, $P$ denotes the transition dynamics, $R$ is the reward function, and $\rho$ is the distribution over initial states. In a CMDP, denoted by $c\mathcal{M}(x)=(S,A,P_{x},R_{x},\rho_{x})$, these components are parameterized by a context variable $x\in X$, which can influence both dynamics and rewards [14, 27]. Consider zero-shot transfer from a source task (MDP) to solve another target task (MDP) in the CMDP. The notation $x$ is used to represent a context for the source task, with each task $x$ belonging to the finite set of CMDPs denoted as $X$. The context of the target task is similarly denoted by $x^{\prime}$ and is also an element of $X$. In our framework, the expected return of a MDP $c\mathcal{M}(x)$ with context $x$ is denoted by $J(x)$. We differentiate between estimated values $\hat{J}(x)$ and observed outcomes $J(x)$, with the latter measured after training and evaluation.
Generalization gap via zero-shot transfer. Zero-shot transfer involves applying a model trained on a source task $c\mathcal{M}(x)$ to a different target task $c\mathcal{M}(x^{\prime})$, with $x^{\prime}$ also belonging to the set $X$. We observe the generalized performance, denoted by $U(x^{\prime};x)$, by evaluating the target task $x^{\prime}$ based on the model trained using source task $x$ via zero-shot generalization. As the concept of “transferability” has been discussed in literature [32, 3], we define the generalization gap as the absolute performance difference in average reward when transferring from source task $x$ to target task $x^{\prime}$:
$\underbrace{\Delta J(x,x^{\prime})}_{\text{Generalization gap}}=\underbrace{U(x^{\prime};x)}_{\text{Generalized performance}}-\underbrace{J(x)}_{\text{Source task performance}}.$ (1)
3 Problem formulation
Sequential source task selection problem. The selection of source tasks in CMDPs is key to solving the overall CMDP [3]. Specifically, we consider the sequential source task selection (SSTS) problem, which seeks to maximize the expected performance across a dynamically selected set of tasks. This problem is framed as a sequential decision problem, in which the selection of tasks is informed through feedback from the observed task performance of the tasks selected and trained thus far. The notation $x_{k}$ indicates the selected source task at the $k$-th transfer step, where $k$ ranges from 1 to $K$.
Definition 1 (Sequential Source Task Selection Problem).
The sequential source task selection (SSTS) problem seeks to optimize the overall expected performance across a CMDP $c\mathcal{M}(\cdot)$ by selecting one task $x\in X$ at each stage. Formally, at each selection step $k$, we aim to choose a task $x_{k}$ such that the cumulative estimated performance is maximized:
$\max_{x_{k}}\ \ V(x_{k};x_{1},...,x_{k-1})\quad\text{s.t. }x_{k}\in X\setminus\{x_{1},...,x_{k-1}\}.$ (2)
where $V(\cdot)$ is defined in the following. Intuitively, we can keep track of the highest performance achievable by any model trained thus far. Formally, after we train the $k^{\text{th}}$ model at source task $x_{k}$, we update the generalized performance, denoted by $U(x^{\prime};x_{(1:k)})$, by comparing the generalized performance from the model trained using task $x_{k}$. We recursively define $U$ based on the previous observations, as follows:
$U(x^{\prime};x_{(1:k)})=\max\left(U(x^{\prime};x_{k}),U(x^{\prime};x_{(1:k-1)})\right)\quad\forall x^{\prime}\in X\quad\text{if }k>1.$ (3)
For simplicity, we denote the sequence $x_{1},x_{2},...,x_{k}$ by $x_{(1:k)}$. We can calculate the observed expected generalized performance, denoted as $V(x_{(1:k)})=\mathbb{E}_{x^{\prime}\in X}\left[U(x^{\prime};x_{(1:k)})\right]$.
The state at each step $k$ is defined by the best performance for each task, achieved by models trained in earlier stages, represented as $U(x^{\prime};x_{(1:k-1)})$ for each target task $x^{\prime}$. The action at each step involves choosing a new task $x_{k}$, aimed at optimizing generalized performance based on current knowledge. The cumulative reward is quantified by the observed performance of the combined task sequence, $V(x_{(1:k)})$, providing a direct measure of the strategic value of each decision. In general, SSTS exhibits stochastic transitions, for example due to randomness in RL training. For simplicity, in this work, we assume deterministic transitions; that is, training MDP $x$ will always yield the same performance $J(x)$ and generalization gap $\Delta J(x,x^{\prime}),\forall x^{\prime}\in X$. The problem’s horizon is defined by $|X|$, indicating that all models will be trained unless a specific termination condition is met. This termination condition is determined by reaching a predefined level of suboptimality.
Model-based transfer learning (MBTL). All methods that solve SSTS are transfer learning methods, by definition. We additionally designate a method that solves SSTS to be model-based if an explicit structure is imposed upon $\Delta J(\cdot,\cdot)$. To the best of the authors’ knowledge, this is the first work to study such methods in the context of reinforcement learning.
4 Methods for Model-Based Transfer Learning
Algorithm 1 Model-based Transfer Learning (MBTL)
0:CMDPs $c\mathcal{M}(x)$, Task $x\in X$, Transfer budget $K$
0:$\boldsymbol{\pi}$ and $V$ Initialize : $J,V=0\ \forall x\in X$, $\boldsymbol{\pi}=\{\}$, $k=1$
1:while $k\leq K$ do
2:    $x_{k}\leftarrow\textbf{NextTask}(J_{k-1},V_{k-1})$
3:    $\pi_{k}\leftarrow\text{Train}(\mathcal{M}(x_{k}))$
4:    $\boldsymbol{\pi}\leftarrow\boldsymbol{\pi}\cup\{\pi_{k}\}$
5:    Update $J(x),V(x)$
6:    $k\leftarrow k+1$
7:end while
8:return  $\boldsymbol{\pi}$ and $V$
(a) Pseudo-greedy strategy
(b) Equidistant strategy
Figure 3: Naive methods for solving SSTS. Note that the normalized performance of the trained tasks is assumed to always be 1 (known and constant).
4.1 Modeling assumptions
We make key assumptions about the task space, performance functions, and the generalization gap.
Assumption 1 (Continuity of the task space).
The task set $X$ is continuous, meaning that for any two tasks $x,x^{\prime}\in X$ and any $\epsilon>0$, there exists a $\delta>0$ such that $\|x-x^{\prime}\|<\delta$ implies the corresponding tasks $x$ and $x^{\prime}$ are indistinguishably close. Formally,
$\forall x,x^{\prime}\in X,\forall\epsilon>0,\exists\delta>0\text{ such that }\|x-x^{\prime}\|<\delta\implies\|J(x)-J(x^{\prime})\|<\epsilon.$
Assumption 2 (Smoothness of the performance function).
The performance function $J(x)$, $U(x)$, or $V(x)$ is smooth. This implies that these functions are continuously differentiable over $x$. Formally, $J(x),U(x),V(x)\in C^{1}(X)$, where $C^{1}(X)$ denotes the space of continuously differentiable functions on $X$.
Assumption 3 (Linear generalization gap).
A linear model is used to model the generalization gap function, formally $\Delta\hat{J}(x,x^{\prime})\simeq\theta|x-x^{\prime}|$, where $\theta$ is the slope of the linear function and $x$ and $x^{\prime}$ are the context of the source task and target task, respectively.
Assumption 1 and 2 are for implementing MBTL, which relies on the continuous task space and smooth performance function. Moreover, we simplify the generalization gap function as a linear function inspired by empirical observations such as shown in Figure 1. Such structure (Assumption 3) enables the development of provable algorithms for solving SSTS. Algorithm 1 provides a generic psuedocode for a MBTL algorithm. The generalization gap structure is leveraged within Line 2 (function NextTask).
4.2 Naive strategies
To understand what is desirable in a method for solving SSTS, it is instructive to consider some naive strategies. An Equidistant Strategy (ES) samples source tasks so that the tasks are equidistant along the full range of tasks we aim to solve. In this method, source tasks are selected to be equidistant within the context space based on the transfer budget $K$. Mathematically, the selected source tasks are chosen such that the distances between consecutive tasks are approximately equal, i.e., $x_{k}=\frac{2k-1}{2K}|X|\quad\text{for }k=1,2,\ldots,K.$ Note that this strategy requires advanced knowledge of the total number of training tasks (e.g., training budget). A Greedy Strategy (GS) greedily selects the source task that is predicted to have the highest generalized performance in each round, assuming known and constant training performance. Figure 3 depicts the difference between ES and GS; observe that the normalized performance of the trained tasks is assumed to be 1. However, a critical question emerges concerning scenarios where training performance is unknown or stochastic. Furthermore, ES and GS are open-loop strategies; that is, they do not incorporate feedback from actual training outcomes.
4.3 Bayesian optimization
Bayesian optimization (BO) is a powerful strategy for finding the global optimum of an objective function when obtaining the function is costly or the function itself lacks a simple analytical form [26, 7]. BO integrates prior knowledge with observations to efficiently decide the next task to train, aiming to optimize the predicted value.
Gaussian Process. Within the framework of BO, we model the source training performance $J_{\text{pred}}(x)$ using Gaussian Process (GP) regression. Specifically, the function $J_{\text{pred}}(x)$ is assumed to follow a GP $(J_{\text{pred}}(x)\sim\mathcal{GP}(m(x),k(x,x^{\prime})))$, where $m(x)$ represents the mean function, $\mathbb{E}[J_{\text{pred}}(x)]$, and $k(x,x^{\prime})$ is the covariance function, representing the expected product of deviations of $J_{\text{pred}}(x)$ and $J_{\text{pred}}(x^{\prime})$ from their respective means. Let $D_{k-1}$ denote the data observed up to iteration $k-1$, consisting of the pairs $\{(x_{i},V_{\text{obs}}(x_{i}))\}_{i=1,...,k-1}$, where $V_{\text{obs}}(x_{(1:i)})$ are the observed outcomes. The posterior prediction of $J_{\text{pred}}$ at a new point $x$, given the data $D_{k-1}$ and previous inputs $x_{(1:k-1)}$, is normally distributed as $P(J_{\text{pred},k}\mid D_{k-1})=\mathcal{N}(\mu_{k}(x),\sigma_{k}^{2}(x))$. $\mu_{k}(x)$ and $\sigma_{k}^{2}(x)$ are defined as $\mu_{k}(x)=m(x)+\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{y}$ and $\sigma_{k}^{2}(x)=k(x,x)-\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{k}$, with $\mathbf{k}$ being the vector of covariances between $x$ and each $x_{i}$ in the observed data, i.e., $\mathbf{k}=[k(x,x_{1}),\ldots,k(x,x_{k-1})]$, and $\mathbf{K}$ is the covariance matrix for the observed inputs, defined as $\mathbf{K}=[k(x_{i},x_{j})]_{1\leq i,j\leq k-1}$. This formulation enables the GP to update its beliefs about the posterior prediction with every new observation, progressively improving the estimation.
Acquisition function. The acquisition function plays a critical role in BO by guiding the selection of the next source training task. At each decision step $k$, the task $x_{k}$ is chosen by maximizing the acquisition function, as denoted by $x_{k}=\operatorname*{arg\,max}_{x}a(x;x_{(1:k-1)})$. Especially in our case, the acquisition function can be designed as the expected marginal improvement of generalized performance across all tasks. One effective strategy employed in the acquisition function is the upper confidence bound (UCB) acquisition function, which considers the trade-off between the expected performance of a task based on current models (exploitation) and the measure of uncertainty associated with the task’s outcome (exploration). It is defined as follows:
$a(x;x_{(1:k-1)})=\mathbb{E}_{x^{\prime}\in X}[[\mu_{k-1}(x)+\beta_{k}^{1/2}\sigma_{k-1}(x)-\Delta J(x,x^{\prime})-U(x^{\prime};x_{(1:k-1)})]_{+}]$ (4)
where $[\cdot]_{+}$ represents $\max(\cdot,0)$ and we can use various forms of $\beta_{k}$.
Generalized performance. We focus on the generalized performance evaluated across all target tasks, defined as $g(x)=\mathbb{E}_{x^{\prime}\in X}[J_{\text{pred}}(x)-\Delta J(x,x^{\prime})]=J_{\text{pred}}(x)-\mathbb{E}_{x^{\prime}\in X}[\Delta J(x,x^{\prime})]$.
4.4 Regret analysis
We use regret to quantify the effectiveness of our source task selection based on BO. Specifically, we define regret at iteration $k$ as $r_{k}=g(x^{*}_{k})-g(x_{k})$, where $g(x^{*}_{k})$ represents the maximum generalized performance achievable across all tasks, and $g(x_{k})$ is the generalized performance at the current task selection $x_{k}$. Consequently, the cumulative regret after $K$ iterations is given by $R_{K}=\sum_{k=1}^{K}r_{k}$, summing the individual regrets over all iterations. Following the framework presented by Srinivas et al. [35], our goal is to establish that this cumulative regret grows sublinearly with respect to the number of iterations. Mathematically, we aim to prove that $\lim_{K\rightarrow\infty}\frac{R_{K}}{K}=0$, indicating that, on average, the performance of our strategy approaches the optimal performance as the number of iterations increases.
Regret of MBTL-GP. Having established the general framework for regret analysis, we now turn our attention to the specific regret properties of our MBTL-GP algorithm. To analyze the regret of MBTL-GP, consider the scaling factor for the UCB acquisition function given by $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$ in Equation 4. It is designed to achieve sublinear regret with high probability, aligning with the theoretical guarantees outlined in Theorem 1 and 5 from [35].
Theorem 1 (Sublinear Regret).
Given $\delta\in(0,1)$, and with the scaling factor $\beta_{k}$ as defined, the cumulative regret $R_{K}$ is bounded by $\sqrt{KC_{1}\beta_{K}\gamma_{K}}$ with a probability of at least $1-\delta$. The formal expression of this probability is $Pr\left[R_{K}\leq\sqrt{KC_{1}\beta_{K}\gamma_{K}}\right]\geq 1-\delta$, where $C_{1}:=\frac{8}{\log(1+\sigma^{-2})}\geq 8\sigma^{2}$ and $\gamma_{K}=\mathcal{O}(\log K)$ for the squared exponential kernel.
Impact of search space elimination. In this section, we demonstrate that strategic reduction of the possible sets, guided by insights from previous task selections or source task training performance, leads to significantly tighter regret bounds than Theorem 1. By focusing on the most promising regions of the task space, our approach enhances learning efficiency and maximizes the model’s performance and applicability. Given the generalization gap observed in Figure 1, we observe that performance loss decreases as the context similarity increases. We model the degradation from the source task using a simple linear function in Assumption 3. Training on the source task can solve a significant portion of the remaining tasks. Our method progressively eliminates partitions of the task space at a certain rate with each iteration. If the source task selected in the previous steps could solve the remaining target task sufficiently, we can eliminate the search space at a desirable rate. Consequently, at each step, we effectively reduce the search space.
Given the generalization gap observed in Figure 1, we model the degradation from the source task using a simple linear function in Assumption 3. While the figure might not strictly appear linear, the linear approximation simplifies analysis and is supported by empirical observations. Training on the source task can solve a significant portion of the remaining tasks. Our method progressively eliminates partitions of the task space at a certain rate with each iteration. If the source task selection in the previous step sufficiently addresses the remaining target tasks, we can reduce the search space at a desirable rate. Consequently, at each step, we effectively focus on a reduced search space.
Definition 2 (Reduced search space).
We define the reduced search space $X_{k}$ as a subset of $X$, with each element $x^{\prime}\in X_{k}$, such that $U(x^{\prime};x_{(1:k-1)})\leq\hat{J}(x_{k})+\Delta J(x^{\prime},x_{k})$.
We leverage the reduced uncertainty in well-sampled regions to tighten the regret bound while slightly lowering the probability $\delta$ in Theorem 1. For the regret analysis, we propose the following theorem based on the generalization of Lemma 5.2 and 5.4 in [35] to the eliminated search space.
Theorem 2.
For a given $\delta^{\prime}\in(0,1)$ and scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$, the cumulative regret $R_{K}$ is bounded by $\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}$ with probability at least $1-\delta^{\prime}$.
Here, $|X|$ denotes the cardinality of the set $X$, the number of elements in $X$. Theorem 2 matches the Theorem 1 when $X_{k}=X$ for all $k$. This theorem implies that regret has a tighter or equivalent bound if we can design the smaller search space instead of searching the whole space. The comprehensive proof is provided in Appendix A.1.2.
Here are some examples of restricted search space: If we consider an example where $|X_{k}|=\frac{1}{\sqrt{k}}|X|$, the regret can be bounded tighter than that of Theorem 1.
Corollary 2.1.
Consider $|X_{k}|=\frac{1}{\sqrt{k}}|X|$. The regret bound would be $R_{K}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\log K}$ with a probability of at least $1-\delta^{\prime}$.
Figure 4: Empirical results of the restriction of search space by MBTL-GP compared to two examples from Corollaries 2.1 and 2.2.
In cases where the search space is defined using MBTL-GS, the largest segment’s length would reduce geometrically, described by $|X_{k}|\leq 2^{-\lfloor\log_{2}k\rfloor}|X|$.
Corollary 2.2.
The regret bound for the $|X_{k}|\leq 2^{-\lfloor\log_{2}k\rfloor}|X|$ would be $R_{K}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\pi^{2}/6}$ with a probability of at least $1-\delta^{\prime}$.
Proofs for Corollaries 2.1 and 2.2 are provided in Appendix A.1.3 and A.1.4, respectively. Based on our experiments presented in Section 5, the rate of elimination of the largest segment for MBTL-GP is shown in Figure 4.
5 Experiments and analysis
5.1 Setup
Our experiments span a variety of CMDP settings, focusing on transportation tasks as well as classic control benchmarks. We introduce three intelligent transportation tasks that utilize DRL but suffer from brittleness issues, as well as the classic control tasks.
Baselines. Our proposed methods are compared against a range of baselines including (1) Random selection, where each training task is chosen randomly; (2) Exhaustive training, which involves independent training separate models on all tasks exhaustively but is computationally expensive; (3) Multi-task RL, where a single context-conditioned model is trained for all tasks; (4) Oracle transfer, which assumes full knowledge of generalized performance for all policies and selects the best source task for each target task.
Proposed methods. Our proposed methods are: (1) MBTL-ES, selecting source tasks for the context to be equidistance based on the transfer budget; (2) MBTL-GS, selecting the next source task greedily to maximize the marginal improvement of expected generalized performance based on linear generalization gap (Assumption 3); (3) MBTL-GP, selecting source task based on the BO using GP and UCB with scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$.
DRL algorithms and performance measure. We utilize Deep Q-Networks (DQN) for discrete actions [24] and Proximal Policy Optimization (PPO) for continuous actions [30]. We evaluate our methods by the average performance across all target tasks after training 15 source tasks and the number of trained models needed to achieve a certain level of suboptimality. We employ min-max normalization of the rewards for each task.
5.2 Traffic tasks
Table 1: Comparative performance of different methods on traffic CMDP tasks
Benchmark Baselines MBTL (Ours) Oracle
Task Context Variation Random Exhaustive Multitask MBTL-GS MBTL-ES MBTL-GP Oracle Transfer
Number of Trained Models $k$ $N$ $1$ $k$ $K$ $k$ $N$
Traffic Signal Road Length 0.9249 0.9409 0.8242 0.9278 0.9213 0.9371 0.9409
Traffic Signal Inflow 0.8457 0.8646 0.8319 0.8496 0.8700 0.8673 0.8768
Traffic Signal Speed Limit 0.8821 0.8857 0.6083 0.8862 0.8858 0.8854 0.8876
Eco-Driving Penetration Rate 0.5959 0.5260 0.1945 0.5827 0.5934 0.6323 0.6660
Eco-Driving Inflow 0.4774 0.4061 0.2229 0.4673 0.4705 0.5108 0.5528
Eco-Driving Green Phase 0.4406 0.3850 0.4228 0.4431 0.4557 0.4700 0.5027
AA-Ring-Acc Hold Duration 0.8924 0.8362 0.9209 0.8776 0.9057 0.9242 0.9552
AA-Ring-Vel Hold Duration 0.9785 0.9589 0.9720 0.9807 0.9772 0.9816 0.9822
AA-Ramp-Acc Hold Duration 0.6050 0.4276 0.5158 0.6143 0.5956 0.6318 0.7111
AA-Ramp-Vel Hold Duration 0.6690 0.5473 0.5034 0.5907 0.6787 0.7182 0.7686
Average 0.7312 0.6778 0.6017 0.7220 0.7354 0.7559 0.7844
†Higher the better. The bold indicates the best for each benchmark, excluding the Oracle transfer. Detailed results with variance for each method are provided in Appendix A.2.2.
‡AA: Advisory autonomy tasks, Ring: Single lane ring, Ramp: Highway ramp, Acc: Acceleration guidance, Vel: Speed guidance.
We validate our methods on diverse simulated traffic scenarios, focusing on to what extent Our proposed methods can optimize the global objective only with a small number of trained models.First, while most traffic lights operate on fixed schedules, we can design learning-based adaptive (1) Traffic signal control to optimize the traffic [8, 21]. However, considering that every intersection looks different, challenges persist in generalizing across various intersection configurations [18]. Given the significant portion of greenhouse gas emissions in the United States coming from the transportation sector [11], the second traffic task is (2) Eco-driving Lagrangian control, which is critical for climate change mitigation. DRL-based eco-driving strategies have been developed [13, 38, 17] but still have some issues of difficulties in generalization. The last traffic task is (3) Advisory autonomy, which is a real-time speed advisory system that enables human drivers to emulate the system-level performance of autonomous vehicles in mixed autonomy systems [34, 15]. Instead of direct and instantaneous control, human drivers receive periodic guidance, which varies based on the traffic states. In Appendix A.2, we provide comprehensive details about our experiments.
Figure 5: Comparison of normalized generalized performance of all target tasks.
Discussion. Table 1 compares the min-max normalized generalized performance across the full target task range. While exhaustive training is computationally expensive and exhibits suboptimal performance, the Oracle transfer, which assumes perfect knowledge of transferability, demonstrates the potential benefits of zero-shot transfer for solving the complete range of multiple tasks. Our proposed MBTL methods leverage this transferability by intelligently selecting the next source task to learn from. Notably, MBTL-GP shows superior performance, closely aligning with the Oracle transfer, which suggests the Gaussian Process effectively models the underlying task dynamics, facilitating better generalization from learned experiences. For example, MBTL-GP significantly improves generalized performance with a few source tasks for traffic signal control problems varying green phase durations (Figure 5). More results are provided in Appendix A.2.
5.3 Standard control benchmarks
Table 2: Comparative performance of different methods on standard control CMDP tasks
Benchmark Baselines MBTL (Ours) Oracle
Task Context Variation Random Exhaustive Multitask MBTL-GS MBTL-ES MBTL-GP Oracle Transfer
Number of Trained Models $k$ $N$ $1$ $k$ $K$ $k$ $N$
Pendulum Length 0.7270 0.7383 0.6830 0.7327 0.7092 0.7697 0.7969
Pendulum Mass 0.6329 0.6237 0.5793 0.6408 0.6092 0.6827 0.7132
Pendulum Timestep 0.7989 0.8135 0.7247 0.8177 0.7541 0.8141 0.8801
Cartpole Mass of Cart 0.7221 0.9466 0.7153 0.6501 0.7516 0.8212 0.9838
Cartpole Length of Pole 0.8121 0.9110 0.5441 0.8217 0.8428 0.9124 0.9875
Cartpole Mass of Pole 0.8858 0.9560 0.6073 0.8744 0.7909 0.9351 1.0000
BipedalWalker Gravity 0.9330 0.9281 0.7898 0.9359 0.9494 0.9393 0.9674
BipedalWalker Friction 0.9650 0.9317 0.9051 0.9645 0.9664 0.9713 0.9778
BipedalWalker Scale 0.8605 0.8694 0.7452 0.8792 0.8496 0.8886 0.9107
HalfCheetah Gravity 0.8542 0.6679 0.6292 0.8663 0.8634 0.9073 0.9544
HalfCheetah Friction 0.8567 0.6693 0.7242 0.8591 0.8703 0.9274 0.9663
HalfCheetah Stiffness 0.8533 0.6561 0.7007 0.8785 0.7817 0.9146 0.9674
Average 0.8251 0.8093 0.6957 0.8267 0.8116 0.8736 0.9255
†Higher the better. The bold indicates the best for each benchmark, excluding the Oracle transfer. Detailed results with variance for each method are provided in Appendix A.2.2.
Building on the evaluation of a suite of traffic-related tasks, we extend our investigation to classic control tasks to assess the applicability of our methods across a broader spectrum of environments. We utilize context-extended versions of standard RL environments from CARL benchmark library [6] to rigorously test the robustness and adaptability of our methods under varied contexts. For Cartpole tasks, we explored contextual Markov decision processes (CMDPs) with varying cart masses, pole lengths, and pole masses. We vary the timestep duration, pendulum length, and pendulum mass in Pendulum. The BipedalWalker was tested under varying friction, gravity, and scale. In HalfCheetah tasks, we manipulated friction, gravity, and stiffness parameters. These variations critically influence the dynamics and physics of the environments. The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times, enabling a comprehensive analysis of transfer learning under drastically different conditions. We provide more details in Appendix A.2.
Discussion. Table 2 compares the performance across the full target task range. Our experiments demonstrate that MBTL significantly improves performance and sample efficiency in dynamic urban transportation systems compared to the baselines. In control tasks such as Cartpole, Pendulum, BipedalWalker, and HalfCheetah, where physical parameters like mass directly influence the dynamics, MBTL methods, particularly MBTL-GP, excel by effectively transferring knowledge between similar task variations. For example, in Figure 5, MBTL-GP outperforms other baselines, including multitask training. The results underscore the potential of MBTL in solving multiple related tasks.
Figure 6: Sensitivity analysis on different DRL algorithms (Cartpole, length of pole).
Figure 7: Sensitivity analysis on the different acquisition functions.
5.3.1 Sensitivity analysis
DRL algorithms. Figure 6 shows an ablation study comparing different DRL algorithms—DQN, PPO, and Advantage Actor-Critic (A2C) [25]. The heatmap results show differences in transferability, and the results underscore the robustness of MBTL to different algorithms.
Acquisition functions. Figure 7 assesses the role of acquisition functions in Bayesian optimization. While expected improvement (EI) focuses on promising marginal gains beyond the current best, UCB utilizes both mean and variance for balancing exploration and exploitation. For Cartpole and advisory autonomy tasks, UCB with decreasing $\beta_{k}$ as $k$ increases shows the best performance. However, for eco-driving control, the $\beta_{k}$ formulation in Section 4.4 exhibits the highest performance, though the difference is not statistically significant.
6 Related work
Robustness and generalization challenges in DRL are generally addressed by a few common techniques in the literature. The broader umbrella of such methods falls under CRL [6]. CRL utilizes side information about the problem variations to improve the generalization and robustness. In particular, CRL formalizes generalization in DRL using CMDPs [14, 27, 6], which incorporate context-dependent dynamics, rewards, and initial state distributions into the formalism of MDPs. This provides a structured framework to study generalization in DRL [14, 27].
The contexts of CMDPs are not always visible during training [20]. When they are visible, they can be directly used as side information by conditioning the policy on them [33]. A common approach is multi-task learning, which aims to leverage a common structure between tasks. The structure between tasks has been exploited using policy sketches for task structuring [2] and by sharing a distilled policy that captures common behavior across tasks [36]. However, when the context features are not visible, the resulting CMDP becomes a partially observable CMDP [20, 9] and can be challenging for multi-task learning. In this work, we use multi-task learning as a baseline in evaluating our methods.
Zero-shot transfer is another commonly used technique that focuses on adapting models trained in one environment to perform effectively in unseen settings without further training [20]. Rezaei-Shoshtari et al. [29] utilize hypernetworks to approximate an RL algorithm as a mapping from parameterized CMDP to a family of near-optimal solutions to achieve zero-shot transfer across varying task conditions. Sinapov et al. [32] uses meta-data to learn inter-task transferability to learn the expected benefit of transfer given a source-target task pair.
Another popular class of techniques for robust and generalizable DRL is domain adaptation strategies. It is often used in making learned policies adapt to new domains with varying dynamics, observations, or rewards [5]. For instance, Xing et al. [39] first learn a latent unified state representation for different domains and subsequently train DRL agents in the source domain based on that. However, our focus in this work centered around in-distribution generalization, whereas domain adaptation techniques are commonly used for out-of-distribution generalization.
7 Conclusion
Our study refines the robustness of DRL by developing MBTL using zero-shot transfer. MBTL strategically selects a set of source tasks to maximize overall performance and minimize training costs. It models the generalization gap and uses BO to predict task performance efficiently. We theoretically show that the method exhibits regret that is sublinear in the number of training tasks. We validate MBTL with urban traffic and control benchmarks. Limitation is that MBTL relies highly on assumptions that may not hold in highly dynamic or complex environments, potentially limiting its applicability and leading to suboptimal policy convergence. Future work will focus on improving the modeling of the generalization gap and exploring scenarios with multiple context variations where simple strategies like MBTL-ES or MBTL-GS may not be applicable.
Appendix A Appendix
[…]
A.1 Theoretical analysis
A.1.1 Notation
Table 3 describes the notation used in this paper.
Table 3: Notation used in the problem formulation
Symbol Description
$x$ Source task ($x\in X$)
$x^{\prime}$ Target task ($x^{\prime}\in X$)
$x_{k}$ Selected source task at transfer step $k$ ($k=1,...,K$)
$\mathbf{x}$ Vector of source tasks ($\mathbf{x}=[x_{1},...,x_{k}]$ where $x_{i}\in X$ for $1\leq i\leq k$)
$c\mathcal{M}(x)$ Contextual MDP parameterized by $x$
$J(x)$ Performance of task $c\mathcal{M}(x)$
$\Delta\hat{J}(x,x^{\prime})$ Generalization gap (source: $x$, target: $x^{\prime}$)
$U(x^{\prime};x)$ Transferred performance (source: $x$ (or $\mathbf{x}$), target: $x^{\prime}$)
$V(x)$ Expected generalized performance of source model $x$ evaluated on all $x^{\prime}\in X$
Figure 8 helps understand the discrepancy between the observed generalized performance and the predicted one. Figure 9 illustrates how to calculate the marginal improvement of expected generalized performance ($\hat{V}(x;x_{1},...,x_{k-1})-V(x_{1},...,x_{k-1})$).
[…]
Figure 8: Illustration of the discrepancy between observed ($U$) and predicted ($\hat{U}$) generalized performance after training on source task $x_{1}$ and attempting zero-shot transfer to $x^{\prime}$.
[…]
Figure 9: Step for choosing $x_{2}$ that maximizes the estimated marginal improvement ($\hat{V}(x;x_{1})-V(x_{1})$). $\hat{V}(x;x_{1})$ corresponds to the red area under the red line and $V(x_{1})$ as the area under $U(x^{\prime};x_{1})$.
A.1.2 Proof of Theorem 2
See 2
Proof.
The following Lemma 3 and 4 is basically considering the cardinality of restricted search space of $X_{k}$ instead of $X$ upon the lemmas in literature [35].
Lemma 3.
For $t\geq 1$, if $|f(x)-\mu_{k-1}(x)|\leq\beta_{k}^{1/2}\sigma_{k-1}(x)\quad\forall x\in X_{k}$, then the regret $r_{t}$ is bounded by $2|X_{k}|\beta_{k}^{1/2}\sigma_{k-1}(x)/|X|$.
Lemma 4.
Setting $\delta\in(0,1)$, $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$, and $C_{1}:=\frac{8}{\log(1+\sigma^{-2})}\geq 8\sigma^{2}$, we have $Pr\left[\sum_{k=1}^{K}r_{k}\left(\frac{|X|}{|X_{k}|}\right)^{2}\leq C_{1}\beta_{K}\gamma_{K}\quad\forall K\geq 1\right]\geq 1-\delta$.
Using Lemma 3, Lemma 5.3 in [35], Lemma 4, and Cauchy–Schwarz inequality, we can bound the cumulative regret as:
$R_{K}=\sum_{k=1}^{K}r_{k}\leq\sqrt{\sum_{k=1}^{K}r_{k}\left(\frac{|X|}{|X_{k}|}\right)^{2}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}.$ (5)
A.1.3 Proof of Corollary 2.1
See 2.1
Proof.
Recall that $\sum_{k=1}^{K}\frac{1}{k}\leq\log K$.
Calculating the sum of squares for the reduced segments, we have:
$\sum_{k=1}^{K}|X_{k}|^{2}=\sum_{k=1}^{K}\frac{1}{k}|X|^{2}\leq|X|^{2}\log K$ (6)
The cumulative regret can be bounded as below:
$R_{K}=\sum_{k=1}^{K}r_{k}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\log K}.$ (7)
A.1.4 Proof of Corollary 2.2
See 2.2
Proof.
Calculating the sum of squares for the reduced segments, we have:
$\sum_{k=1}^{K}|X_{k}|^{2}=\sum_{k=1}^{K}2^{-2\lfloor\log_{2}k\rfloor}|X|^{2}\leq\frac{1}{k^{2}}|X|^{2}\leq\frac{\pi^{2}}{6}|X|^{2}$ (8)
The cumulative regret can be bounded as below:
$R_{K}=\sum_{k=1}^{K}r_{k}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}\leq\sqrt{\frac{C_{1}\beta_{K}\gamma_{K}\pi^{2}}{6}}.$ (9)
A.2 Experiment details
A.2.1 Details about Gaussian process (GP)
In our study, we conducted hyperparameter tuning experiments for Gaussian Process (GP) regression to optimize its performance. Specifically, we varied the noise standard deviation over the set $\{0.001,0.01,0.1,1\}$ and the number of restarts for the optimizer over the set $\{5,10,15\}$. We explored several kernel configurations, including the combination of a constant kernel (C) and a radial basis function kernel (RBF) with a length scale ranging from 0.01 to 100, as well as more complex kernels incorporating a white noise kernel and an exponential sine squared kernel with varying length scales and periodicities. Each GP model was instantiated with the specified kernel, and the noise level was set to the square of the noise standard deviation. This systematic exploration aimed to identify the optimal set of hyperparameters that could enhance the model’s predictive accuracy.
A.2.2 Results of table with standard deviation
Table 4: Comparative performance of different methods on context-variant traffic and benchmark tasks
Benchmark Baselines Model-Based Transfer Learning Oracle
Task Context Variation Random Exhaustive Multitask MBTL-PS MBTL-ES MBTL-GP Oracle Transfer
Number of Trained Models $k$ $N$ $1$ $k$ $K$ $k$ $N$
Traffic Signal Road Length
0.9249
0.9409
0.8242
0.9278
0.9213
0.9371
0.9409
Traffic Signal Inflow
0.8457
0.8646
0.8319
0.8496
0.8700
0.8673
0.8768
Traffic Signal Speed Limit
0.8821
0.8857
0.6083
0.8862
0.8858
0.8854
0.8876
Eco-Driving Penetration Rate
0.5959
0.5260
0.1945
0.5827
0.5934
0.6323
0.6660
Eco-Driving Inflow
0.4774
0.4061
0.2229
0.4673
0.4705
0.5108
0.5528
Eco-Driving Green Phase
0.4406
0.3850
0.4228
0.4431
0.4557
0.4700
0.5027
AA-Ring-Acc Hold Duration
0.8924
0.8362
0.9209
0.8776
0.9057
0.9242
0.9552
AA-Ring-Vel Hold Duration
0.9785
0.9589
0.9720
0.9807
0.9772
0.9816
0.9822
AA-Ramp-Acc Hold Duration
0.6050
0.4276
0.5158
0.6143
0.5956
0.6318
0.7111
AA-Ramp-Vel Hold Duration
0.6690
0.5473
0.5034
0.5907
0.6787
0.7182
0.7686
Pendulum Length
0.7270
0.7383
0.6830
0.7327
0.7092
0.7697
0.7969
Pendulum Mass
0.6329
0.6237
0.5793
0.6408
0.6092
0.6827
0.7132
Pendulum Timestep
0.7989
0.8135
0.7247
0.8177
0.7541
0.8141
0.8801
Cartpole Mass of Cart
0.7221
0.9466
0.7153
0.6501
0.7516
0.8212
0.9838
Cartpole Length of Pole
0.8121
0.9110
0.5441
0.8217
0.8428
0.9124
0.9875
Cartpole Mass of Pole
0.8858
0.9560
0.6073
0.8744
0.7909
0.9351
1.0000
BipedalWalker Gravity
0.9330
0.9281
0.7898
0.9359
0.9494
0.9393
0.9674
BipedalWalker Friction
0.9650
0.9317
0.9051
0.9645
0.9664
0.9713
0.9778
BipedalWalker Scale
0.8605
0.8694
0.7452
0.8792
0.8496
0.8886
0.9107
HalfCheetah Gravity
0.8542
0.6679
0.6292
0.8663
0.8634
0.9073
0.9544
HalfCheetah Friction
0.8567
0.6693
0.7242
0.8591
0.8703
0.9274
0.9663
HalfCheetah Stiffness
0.8533
0.6561
0.7007
0.8785
0.7817
0.9146
0.9674
Average
0.8068
0.7775
0.6337
0.8037
0.7996
0.8435
0.8833
* Note: The bold values indicate the maximum value for each task excluding the oracle. Standard deviation across multiple runs in the parenthesis.
A.2.3 Details about traffic signal control
Most traffic lights operate on fixed schedules, but adaptive traffic signal control using DRL can optimize the traffic flow using real-time information on the traffic [8, 21], though challenges persist in generalizing across various intersection configurations [18].
Figure 10 showcases the layout of traffic networks used in a traffic signal control task with several lanes and a signalized intersection in the middle. The state space represents the presence of vehicles in discretized lane cells along the incoming roads. Actions determine which lane gets the green phase of the traffic signal, and rewards are based on changes in cumulative stopped time, the period when speed is zero. The global objective is to minimize the average waiting times at the intersection. Different configurations of intersections (e.g., road length, inflow, speed limits) are modeled to represent varying real-world conditions.
Training configuration
We used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [22] v.1.16.0 and PPO for RL algorithm [30]. All experiments are done on a distributed computing cluster equipped with 48 Intel Xeon Platinum 8260 CPUs.
Figure 10: Illustration of the traffic networks in traffic signal control task.
Transferability heatmap
Figure 11 presents heatmaps of transferability for different traffic signal control tasks, each varying a specific aspect: inflow, speed limit, and road length. The heatmaps display the effectiveness of strategy transfer from each source task (vertical axis) to each target task (horizontal axis). In terms of inflow variation, transferability drops when transferring from tasks with lower vehicle inflow to those with higher inflow. In speed limit variation, the transferability shows uniform effectiveness, suggesting less sensitivity to these changes. In road length variation, distinct blocks of high transferability indicate that different road lengths may require significantly tailored strategies.
(a) Inflow variation
(b) Speed limit variation
(c) Road length variation
Figure 11: Examples of transferability heatmap for traffic signal control.
Results
Figure 12 illustrates the normalized generalized performance across various traffic control tasks: inflow, speed limit, and road length. The plots display how different strategies adapt with increasing transfer steps:
  • Inflow: Performance improves as the number of transfer steps increases, with MBTL-GP strategy consistently achieving the highest scores, demonstrating their effectiveness in adapting to changes in inflow conditions.
  • Speed Limit: Here, performance levels are relatively stable across all strategies except for the multitask training.
  • Road Length: There is a general upward trend in performance for all strategies, particularly for MBTL-GP, indicating robustness in adapting to different road lengths.
This data suggests that MBTL-GP and Oracle are particularly effective across varying conditions, maintaining higher levels of performance adaptability.
Figure 12: Comparison of normalized generalized performance of all target tasks: Traffic signal control.
A.2.4 Details about eco-driving Lagrangian control
Given the significant portion of greenhouse gas emissions in the United States coming from the transportation sector [11], eco-driving behaviors are critical for climate change mitigation. Deep reinforcement learning-based eco-driving strategies have been developed [13, 38, 17] but still have some issues of difficulties in generalization. We also extend to various intersection configurations with different traffic inflow rates, penetration rates of eco-driving systems, and durations of green phases at static traffic signals to optimize vehicle behaviors for reduced emissions.
Figure 13 illustrates the traffic road network used in eco-driving Lagrangian control task. The road network is depicted with traffic flowing vertically and horizontally crossing the static phase traffic signal. There are both guided and default vehicles in the system. The state space includes the speed and position of the ego vehicle, the leading vehicle, and the following vehicles, supplemented by the current traffic signal phase and relevant context features, including lane length and green phase durations. The action space specifically focuses on the ego vehicle’s acceleration control. The reward mechanism is designed to optimize the driving strategy by balancing the average speed of the vehicles against penalties for emissions, thereby promoting eco-friendly driving behaviors within the traffic system.
Training configuration
We also used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [22] v.1.16.0 and PPO for RL algorithm [30].
Figure 13: Illustration of the traffic networks in eco-driving control task.
Transferability heatmap
Figure 14 displays heatmaps for the eco-driving control task, with each heatmap varying an aspect such as green phase, inflow, and penetration rate. These visuals illustrate the transferability of strategies from source tasks (vertical axis) to target tasks (horizontal axis), highlighting the impact of traffic light phases, vehicle inflow, and the proportion of guided vehicles on strategy effectiveness. Notably, longer green phases correlate with improved performance and transferability. For inflow variations, reduced inflow typically yields better outcomes. However, variations in the penetration rate of guided vehicles show minimal impact on performance differences.
(a) Green phase variation
(b) Inflow variation
(c) Penetration rate variation
Figure 14: Examples of transferability heatmap for eco-driving Lagrangian control.
Results
Figure 15 illustrates the normalized generalized performance across variants of eco-driving control tasks, specifically looking at variations in green phase time, inflow, and penetration rate. The graphs depict performance enhancement over transfer steps for different strategies. Notably, MBTL-GP consistently demonstrates superior performance across all variations, indicating robust adaptability to changing task parameters.
Figure 15: Comparison of normalized generalized performance of all target tasks: Eco-driving Lagrangian control.
A.2.5 Details about advisory autonomy task
Advisory autonomy involves a real-time speed advisory system that enables human drivers to emulate the system-level performance of autonomous vehicles in mixed autonomy systems [34, 15]. Instead of direct and instantaneous control, human drivers receive periodic guidance, which varies based on road type and guidance strategy. Here, we consider the different frequencies of this periodic guidance as contextual MDPs since the zero-order hold action affects the transition function.
Figure 16 illustrates two distinct traffic network configurations used in the advisory autonomy task: a single-lane ring and a highway ramp. The single-lane ring features 22 vehicles circulating the ring, with only one being actively controlled, presenting a relatively controlled environment for testing vehicle guidance systems. The highway ramp scenario introduces a more complex dynamic, where vehicles not only travel along the highway but also merge from ramps, creating potential stop-and-go traffic patterns that challenge the adaptability of autonomous guidance systems.
Problem Definition: In a single-lane scenario, the state space includes the speeds of the ego and leading vehicles, along with the headway. For highway ramp scenarios, additional states cover the relative positions and speeds of adjacent vehicles. Actions vary by guidance type: for acceleration guidance, the action space is continuous, ranging from $-1$ to $1$; for speed guidance, it has ten discrete actions compared to the speed limit. Rewards are based on system throughput or average speed of all vehicles in the system.
Context Variations: We explore different durations of coarse-grained guidance holds to test various levels of human compatibility, adjusting the model based on observed driver behaviors and system performance.
Figure 16: Illustration of the traffic networks in advisory autonomy task.
Transferability heatmap
Figure 17 showcases heatmaps of transferability for advisory autonomy tasks, each varying in specific aspects: acceleration guidance and speed guidance across a single lane ring and a highway ramp. These heatmaps demonstrate the effectiveness of strategy transfer from each source task (vertical axis) to each target task (horizontal axis), capturing how variations in task conditions influence adaptability. For acceleration guidance in a ring setup (a), transferability is generally higher among tasks with similar acceleration demands. In contrast, speed guidance on a ramp (d) reveals more variability in transferability, potentially due to the complexity of speed adjustments in ramp scenarios.
(a) Ring with acceleration guide
(b) Ring with speed guide
(c) Ramp with acceleration guide
(d) Ramp with speed guide
Figure 17: Examples of transferability heatmap for advisory autonomy.
Results
Figure 18 illustrates the comparison of normalized generalized performance for advisory autonomy tasks, specifically acceleration and speed guidance in a ring and acceleration guidance on a ramp. The graphs demonstrate that MBTL-GP consistently exhibits higher performance across all tasks. Particularly, acceleration guidance in both ring and ramp scenarios shows significant performance improvements over transfer steps, with MBTL-GP closely matching in some instances.
Figure 18: Comparison of normalized generalized performance of all target tasks: Advisory autonomy.
A.2.6 Details of control tasks
For this experimental phase, we selected context-extended versions of standard RL environments from the CARL benchmark library, including Cartpole, Pendulum, BipedalWalker, and Halfcheetah. These environments were chosen to rigorously test the robustness and adaptability of our MBTL algorithm under varied conditions that mirror the complexity encountered in real-world scenarios.
Context Variations: In the Cartpole tasks, we explored CMDPs with varying cart masses, pole lengths, and pole masses. For the Pendulum, the experiments involved adjusting the timestep duration, pendulum length, and pendulum mass. The BipedalWalker was tested under different settings of friction, gravity, and scale. Similarly, in the Halfcheetah tasks, we manipulated parameters such as friction, gravity, and stiffness to simulate different physical conditions. These variations critically influence the dynamics and physics of the environments, thereby presenting unique challenges that test the algorithm’s capacity to generalize from previous learning experiences without the need for extensive retraining. The range of context variations was established by scaling the default values specified in the CARL framework from 0.1 to 10 times, enabling a comprehensive examination of each model’s performance under drastically different conditions.
A.2.7 Results of control tasks
Figure 19 compares our proposed methods with baselines and oracle method in different variations of context in classic control tasks.
Figure 19: Comparison on control benchmarks: Cartpole, Pendulum, BipedalWalker, and Halfcheetah.
A.2.8 Details about Cartpole
Transferability heatmap
Figure 20 presents transferability heatmaps for the Cartpole task with variations in three physical properties: mass of the cart, length of the pole, and mass of the pole. Each heatmap illustrates how well strategies transfer from source tasks (vertical axis) to target tasks (horizontal axis), depicting the influence of each parameter on control strategy effectiveness. For the mass of the cart variation (a), transferability decreases as the mass difference increases. In the length of the pole variation (b), strategies are less transferable between significantly different pole lengths. Similarly, for the mass of the pole variation (c), variations show divergent transferability depending on the extent of mass change.
(a) Mass of cart variation
(b) Length of the pole variation
(c) Mass of pole variation
Figure 20: Examples of transferability heatmap for Cartpole.
Results
Figure 21 presents a comparison of normalized generalized performance for the Cartpole task across different strategies when varying the mass of the cart, length of the pole, and mass of the pole.
In the mass of cart variation, the performance generally increases with transfer steps, with MBTL-GP strategies achieving the highest scores, indicating robust adaptability to changes in cart mass. Similar trends are observed with length variation and mass of pole variation. MBTL-GP shows close to oracle performance.
Figure 21: Comparison of normalized generalized performance of all target tasks: Cartpole.
A.2.9 Details about Pendulum
Transferability heatmap
Figure 22 presents transferability heatmaps for the Pendulum task with variations in three physical properties: timestep, length of the pendulum, and mass of the pendulum. Each heatmap illustrates how effectively strategies transfer from source tasks (vertical axis) to target tasks (horizontal axis), highlighting the impact of each parameter on control strategy effectiveness. For the timestep variation (a), there appears to be high consistency in transferability across different timesteps, especially around the diagonal axis. In the length of the pendulum variation (b), transferability decreases with greater length differences. Similarly, for the mass of the pendulum variation (c), transferability shows variability dependent on the extent of mass changes.
(a) Timestep variation
(b) Length of the pendulum variation
(c) Mass of pendulum variation
Figure 22: Examples of transferability heatmap for Pendulum.
Results
Figure 23 shows a comparison of normalized generalized performance for the Pendulum task across different strategies when varying the timestep, length of the pendulum, and mass of the pendulum. For the length of the pendulum variation and mass of the pendulum one, MBTL-GP strategies demonstrate the highest scores, suggesting robust adaptability to changes in pendulum dynamics. MBTL-GP shows performance close to that of the Oracle across all variations, indicating its effectiveness in handling dynamic changes in system parameters.
Figure 23: Comparison of normalized generalized performance of all target tasks: Pendulum.
A.2.10 Details about BipedalWalker
Transferability heatmap
Figure 24 presents transferability heatmaps for the BipedalWalker task, focusing on three variations: friction, gravity, and scale. Each heatmap illustrates the effectiveness of strategy transfer from source tasks (vertical axis) to target tasks (horizontal axis), highlighting how each parameter influences control strategy adaptability. For friction variation (a), strategies show uniform transferability across different friction levels. In gravity variation (b), transferability is highly variable, suggesting that strategies need specific tuning for different gravity levels. For scale variation (c), the heatmap indicates variable transferability, reflecting the challenges of scaling control strategies.
(a) Friction variation
(b) Gravity variation
(c) Scale variation
Figure 24: Examples of transferability heatmap for BipedalWalker.
Results
Figure 25 shows the comparison of normalized generalized performance for all variations within the BipedalWalker task. There is no huge difference in performance for all three cases, but if we look into the tabualr results in Table 2, MBTL-GP shows the highest performance across varying conditions, indicating their robustness in adapting to changes in physical parameters of the model. This suggests that these strategies are more effective in handling the complexities introduced by different frictions, gravities, and scales compared to other baselines.
Figure 25: Comparison of normalized generalized performance of all target tasks: BipedalWalker.
A.2.11 Details about HalfCheetah
Transferability heatmap
Figure 26 displays transferability heatmaps for the HalfCheetah task, focusing on three physical properties: friction, gravity, and stiffness. Each heatmap demonstrates the transferability of strategies from source tasks (vertical axis) to target tasks (horizontal axis). For friction variation (a), there is uniform high transferability across different friction levels, indicating that strategies are robust to changes in friction. Gravity variation (b) shows less consistent transferability, suggesting a sensitivity to gravity changes that might require adaptation of strategies. Stiffness variation (c) similarly demonstrates variable transferability, highlighting the challenges of adapting to different stiffness levels in control strategies.
(a) Friction variation
(b) Gravity variation
(c) Stiffness variation
Figure 26: Examples of transferability heatmap for HalfCheetah.
Results
Figure 27 presents a comparison of normalized generalized performance across various strategies for the HalfCheetah task with respect to the varied physical properties. The results indicate that the MBTL-GP generally outperforms others, particularly in managing variations in gravity and stiffness, suggesting the superior adaptability of these models to physical changes in the task environment. The trends across different parameters confirm the critical impact of task-specific dynamics on the effectiveness of the tested strategies.
Figure 27: Comparison of normalized generalized performance of all target tasks: HalfCheetah.
A.3 Potential impacts
Our work has the potential to reduce the computational effort needed to solve complex real-world problems, offering scalable solutions for implementing deep reinforcement learning in dynamic environments. While there are no immediate negative societal impacts identified, ongoing research will continue to assess the broader implications of deploying these technologies in urban settings.

### v2

@@
 Abstract
-Deep reinforcement learning is a powerful approach to complex decision making.
+Deep reinforcement learning (RL) is a powerful approach to complex decision making.
 However, one issue that limits its practical application is its brittleness, sometimes failing to train in the presence of small changes in the environment.
-This work is motivated by the empirical observation that directly applying an already trained model to a related task often works remarkably well, also called zero-shot transfer.
-We take this practical trick one step further to consider how to systematically select good tasks to train, maximizing overall performance across a range of tasks.
-Given the high cost of training, it is critical to choose a small set of training tasks.
-The key idea behind our approach is to explicitly model the performance loss (generalization gap) incurred by transferring a trained model.
-We hence introduce Model-Based Transfer Learning (MBTL) for solving contextual RL problems.
-In this work, we model the performance loss as a simple linear function of task context similarity.
-Furthermore, we leverage Bayesian optimization techniques to efficiently model and estimate the unknown training performance of the task space.
-We theoretically show that the method exhibits regret that is sublinear in the number of training tasks and discuss conditions to further tighten regret bounds.
-We experimentally validate our methods using urban traffic and standard control benchmarks.
-Despite the conceptual simplicity, the experimental results suggest that MBTL can achieve greater performance than strong baselines, including exhaustive training on all tasks, multi-task training, and random selection of training tasks.
+Motivated by the success of zero-shot transfer—where pre-trained models perform well on related tasks—we consider the problem of selecting a good set of training tasks to maximize generalization performance across a range of tasks.
+Given the high cost of training, it is critical to select training tasks strategically, but not well understood how to do so.
+We hence introduce Model-Based Transfer Learning (MBTL), which layers on top of existing RL methods to effectively solve contextual RL problems.
+MBTL models the generalization performance in two parts: 1) the performance set point, modeled using Gaussian processes, and 2) performance loss (generalization gap), modeled as a linear function of contextual similarity.
+MBTL combines these two pieces of information within a Bayesian optimization (BO) framework to strategically select training tasks.
+We show theoretically that the method exhibits sublinear regret in the number of training tasks and discuss conditions to further tighten regret bounds.
+We experimentally validate our methods using urban traffic and standard continuous control benchmarks.
+The experimental results suggest that MBTL can achieve up to 50x improved sample efficiency compared with canonical independent training and multi-task training.
+Further experiments demonstrate the efficacy of BO and the insensitivity to the underlying RL algorithm and hyperparameters.
 This work lays the foundations for investigating explicit modeling of generalization, thereby enabling principled yet effective methods for contextual RL.
+Code is available at https://github.com/jhoon-cho/MBTL/.
 1 Introduction
 Deep reinforcement learning (DRL) has made remarkable strides in addressing complex problems across various domains [29, 36, 1, 4, 11, 13, 27].
-Despite these successes, DRL models often exhibit brittleness when exposed to small variations in task settings, significantly limiting their scalability and generalizability [28, 40].
-Such variations in Markov Decision Processes (MDPs) can be conceptualized as contextual Markov Decision Processes (CMDP), where slight differences in MDP definitions lead to multiple, closely related but different environments [14, 27].
-Traditional DRL approaches typically require independent training for each new task variant, a process that is both resource-intensive and impractical for real-time applications.
-Multi-task training is a relatively inexpensive way to train but still has limited model capacity issues.
-In light of this, there is a pressing need for more robust approaches that enhance generalization across different tasks.
-It is crucial to differentiate between training robustness, which seeks resilience of training across a variety of tasks, and model robustness, which focuses on the resistance of a model to environmental perturbations.
-Contextual reinforcement learning (CRL) offers a framework for modeling changes in tasks [6].
-In our exploration of applying DRL to solve multiple related tasks, we identify computational challenges, particularly the high costs and complexities associated with training separate models for each task.
-While techniques like CARL [6] or multitask RL [36, 2] address these challenges, they still remain computationally expensive and inefficient when applied across a broad context range.
-Figure 1: Generalization gap in Cartpole CMDPs.
-The solid lines illustrate the actual generalized performance after zero-shot transfer across contexts, with the source context indicated by a dotted line.
-Figure 2: Model-Based Transfer Learning (MBTL) for solving contextual MDPs.
-MBTL framework strategically selects optimal tasks to train by modeling the generalization gap that arises when transferring a model to different tasks.
-For each target task, we deploy the most effective trained model, indicated by solid arrows.
-We evaluate our framework on a suite of continuous and discrete tasks, including standard control benchmarks [6] and urban traffic benchmarks [34, 19].
-We build upon zero-shot transfer, which has shown potential in directly applying a policy trained in one MDP (source task) to another (target task) without adaptation.
-This often leads to a performance degradation known as the ‘generalization gap’ [16, 20].
-For instance, Figure 1 exemplifies the concept of the generalization gap within a Cartpole CMDP, a range of scenarios in which the agent must balance a pole on a moving cart with a different mass of pole.
-The performance degrades as the target task diverges from the source task, illustrating an increasing generalization gap.
-Given that training is expensive yet zero-shot transfer is cheap, we are interested in optimally selecting a set of source (training) tasks, such that performance on the target range of tasks is maximized.
-In this article, we devise strategies to incorporate the structure of the generalization gap to better estimate the value of training.
-Specifically, we investigate how explicit modeling of the generalization gap structure lends itself to principled and effective algorithms for the selection of training tasks within CMDP.
-Hence, this article introduces Model-Based Transfer Learning (MBTL), a novel framework for studying generalization in reinforcement learning (Figure 2).
-We model the explicit form of the generalization gap in zero-shot transfer and use it in estimating how valuable selecting the source task is.
-Given a small number of samples in earlier steps, we leverage Bayesian optimization methods to estimate training performance with limited information, thus significantly reducing the necessity for exhaustive training [35, 37].
-This paper contributes to the field of reinforcement learning by introducing MBTL framework for efficient source task selection, aiming to solve multiple related tasks with enhanced transferability and fewer required models.
-Theoretical analyses detail the sublinear cumulative regret of these methods and provide conditions for achieving tighter regret bounds.
-We validate our methods in simulated traffic management scenarios and standard control benchmarks.
-Our findings provide foundations for investigating explicit modeling of generalization, enabling efficient and practical model-based solutions for solving multiple related tasks for robust training in DRL.
+Despite these successes, DRL algorithms often exhibit brittleness when exposed to small variations like different number of lanes, weather conditions, or flow density in traffic benchmarks [20], significantly limiting their scalability and generalizability [19].
+Such variations can be modeled using the framework of contextual Markov Decision Processes (CMDP), where task variations can be parameterized within a context space [15, 32, 6].
+There are two predominant solution modalities for CMDPs [22]: independent training and multi-task training.
+Independent training constructs a separate model for each task variant (say, $N\gg 1$), which is compute-intensive.
+At the other extreme, multi-task training constructs a single “universal” policy, and thus can be more compute-efficient, but suffers from model capacity and negative transfer issues [21, 44, 2, 41].
+There is thus a need for more reliable training methodologies for generalization across tasks variants.
+In this work, we consider training an intermediate set of $K$ models, where $N>K>1$, in an effort to balance performance and efficiency; we refer to this strategy as multi-policy training.
+A note on terminology.
+For brevity, we refer to task variants as tasks in the remainder of this article.
+We also use the language of task and context interchangeably.
+We emphasize that this work focuses on within-domain generalization (e.g., traffic signal control for intersection scenario variants) rather than across-domain generalization (e.g., distinct traffic control tasks).
+Additionally, it is crucial to differentiate between training reliability, which concerns the ability to reliably train models across tasks, and model reliability (or robustness), which concerns the resistance of a trained model to differences in tasks.
+This article is concerned with training reliability.
+Figure 1: Example generalization gap depicted for Cartpole CMDP.
+The solid lines show the true zero-shot transfer generalization performance across contexts.
+Source tasks are indicated by dotted lines.
+We build upon zero-shot transfer, a widely-used practical technique which directly applies a policy trained in one context (source task) to another (target task) without adaptation.
+This experiences performance degradation, also called generalization gap [17, 22].
+For instance, Figure 1 depicts that the performance degrades as the target task diverges from the source task, corresponding to an increasing generalization gap.
+Nonetheless, leveraging the notion that training is expensive but zero-shot transfer is cheap, we are interested in optimally selecting a set of source (training) tasks, such that the generalization performance on the target range of tasks is maximized.
+In this article, we strategically select source tasks by explicitly modeling the generalization performance to estimate the value of training a new source task.
+The main contributions of this work are:
+  • We introduce Model-Based Transfer Learning (MBTL), a novel framework for studying generalization in reinforcement learning (Figure 2).
+To the best of our knowledge, this is the first work to explicitly model generalization performance for contextual RL.
+As such, our work opens the door for further investigation into reliable model-based methodologies for contextual RL.
+  • We provide theoretical analysis for the sublinear regret of MBTL and give conditions for achieving tighter regret bounds.
+  • We empirically validate our methods in urban traffic and standard continuous control benchmarks for contextual RL, observing up to 50x improvements in sample efficiency.
+We further include ablations on the components of the algorithm.
+Figure 2: Overview illustration for Model-based Transfer Learning.
+(a) Gaussian process regression is used to estimate the training performance across tasks using existing policies; (b) marginal generalization performance (red area) is calculated using upper confidence bound of estimated training performance, generalization gap, and generalization performance; (c) selects the next training task that maximizes the acquisition function (marginal generalization performance); (d) once the selected task is trained, calculate generalization performance using zero-shot transfer.
+The remainder of the paper is organized as follows.
+After introducing notation in Section 2, we formally define the problem in Section 3.
+A key contribution of our work is the introduction of a Gaussian process model acquisition function specifically tailored to the source task selection problem, which is detailed in Section 4.
+In Section 4.3, we provide a theoretical analysis of the regret bounds of our method, followed by an empirical evaluation across diverse applications in Section 5.
 2 Preliminaries and notation
-We leverage the framework of CMDPs [14, 27, 6] to formulate the problem of solving multiple related tasks using reinforcement learning.
-Contextual MDP (CMDP).
-A standard MDP is defined by the tuple $M=(S,A,P,R,\rho)$ where $S$ represents the state space, $A$ is the action space, $P$ denotes the transition dynamics, $R$ is the reward function, and $\rho$ is the distribution over initial states.
-In a CMDP, denoted by $c\mathcal{M}(x)=(S,A,P_{x},R_{x},\rho_{x})$, these components are parameterized by a context variable $x\in X$, which can influence both dynamics and rewards [14, 27].
-Consider zero-shot transfer from a source task (MDP) to solve another target task (MDP) in the CMDP.
-The notation $x$ is used to represent a context for the source task, with each task $x$ belonging to the finite set of CMDPs denoted as $X$.
-The context of the target task is similarly denoted by $x^{\prime}$ and is also an element of $X$.
-In our framework, the expected return of a MDP $c\mathcal{M}(x)$ with context $x$ is denoted by $J(x)$.
-We differentiate between estimated values $\hat{J}(x)$ and observed outcomes $J(x)$, with the latter measured after training and evaluation.
+Contextual MDP.
+A standard MDP is defined by the tuple ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}=(S,A,P,R,\rho)$ where $S$ represents the state space, $A$ is the action space, $P$ denotes the transition dynamics, $R$ is the reward function, and $\rho$ is the distribution over initial states [43].
+A contextual MDP (CMDP), denoted by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}=}(S,A,P_{x},R_{x},\rho_{x})_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x\in X}}$, is a collection of context-MDPs ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x}}$ parameterized by a context variable $x$ within a context space $X$ (assumed bounded).
+The context variable $x$ can influence dynamics, rewards, and initial state distributions [15, 32, 6].
+We define source task performance $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0};\text{Alg}})$ as follows: we train a policy $\pi_{x}$ on a task with the context $x\in X$ using RL algorithm Alg (e.g., PPO, SAC) and evaluate the policy by the expected return in the same MDP $\mathcal{M}_{x}$ with context $x$.
+For brevity, we will write it as $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)$.
+We distinguish between estimated values $\hat{J}(x)$ and observed outcomes $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)$, with the latter measured after training and evaluation.
 Generalization gap via zero-shot transfer.
-Zero-shot transfer involves applying a model trained on a source task $c\mathcal{M}(x)$ to a different target task $c\mathcal{M}(x^{\prime})$, with $x^{\prime}$ also belonging to the set $X$.
-We observe the generalized performance, denoted by $U(x^{\prime};x)$, by evaluating the target task $x^{\prime}$ based on the model trained using source task $x$ via zero-shot generalization.
-As the concept of “transferability” has been discussed in literature [32, 3], we define the generalization gap as the absolute performance difference in average reward when transferring from source task $x$ to target task $x^{\prime}$:
-$\underbrace{\Delta J(x,x^{\prime})}_{\text{Generalization gap}}=\underbrace{U(x^{\prime};x)}_{\text{Generalized performance}}-\underbrace{J(x)}_{\text{Source task performance}}.$ (1)
+Consider zero-shot transfer from the trained policy $\pi_{x}$ from a source task (context-MDP) to solve another target task (context-MDP) with the context $x^{\prime}\in X$ in the CMDP.
+Zero-shot transfer involves applying a policy trained on a source task ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x}}$ to a different target task ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x^{\prime}}}$, with $x,x^{\prime}\in X$.
+We observe the generalization performance, denoted by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{x},x^{\prime}})$, by evaluating the target task $x^{\prime}$ based on the policy trained using source task $x$ via zero-shot generalization.
+We define the generalization gap as the absolute performance difference in average reward when transferring from source task $x$ to target task $x^{\prime}$:
+$\underbrace{\Delta J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{Generalization gap}}=\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)}_{\text{Source task performance}}-\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Generalization} performance}}.$ (1)
 3 Problem formulation
 Sequential source task selection problem.
-The selection of source tasks in CMDPs is key to solving the overall CMDP [3].
-Specifically, we consider the sequential source task selection (SSTS) problem, which seeks to maximize the expected performance across a dynamically selected set of tasks.
-This problem is framed as a sequential decision problem, in which the selection of tasks is informed through feedback from the observed task performance of the tasks selected and trained thus far.
+The selection of source MDPs from the CMDPs is key to solving the overall CMDP [3].
+We therefore introduce the sequential source task selection (SSTS) problem, which seeks to maximize the expected performance across a dynamically selected set of tasks.
+This problem is cast as a sequential decision problem, in which the selection of tasks is informed through feedback from the observed task performance of the tasks selected and trained thus far.
 The notation $x_{k}$ indicates the selected source task at the $k$-th transfer step, where $k$ ranges from 1 to $K$.
+For brevity, we will denote $\pi_{x_{k}}$ as $\pi_{k}$.
+We denote the sequence $x_{1},x_{2},...,x_{k}$ by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k}}$ and $\pi_{1},\pi_{2},...,\pi_{k}$ by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k}}$.
 Definition 1 (Sequential Source Task Selection Problem).
-The sequential source task selection (SSTS) problem seeks to optimize the overall expected performance across a CMDP $c\mathcal{M}(\cdot)$ by selecting one task $x\in X$ at each stage.
-Formally, at each selection step $k$, we aim to choose a task $x_{k}$ such that the cumulative estimated performance is maximized:
-$\max_{x_{k}}\ \ V(x_{k};x_{1},...,x_{k-1})\quad\text{s.t. }x_{k}\in X\setminus\{x_{1},...,x_{k-1}\}.$ (2)
-where $V(\cdot)$ is defined in the following.
-Intuitively, we can keep track of the highest performance achievable by any model trained thus far.
-Formally, after we train the $k^{\text{th}}$ model at source task $x_{k}$, we update the generalized performance, denoted by $U(x^{\prime};x_{(1:k)})$, by comparing the generalized performance from the model trained using task $x_{k}$.
-We recursively define $U$ based on the previous observations, as follows:
-$U(x^{\prime};x_{(1:k)})=\max\left(U(x^{\prime};x_{k}),U(x^{\prime};x_{(1:k-1)})\right)\quad\forall x^{\prime}\in X\quad\text{if }k>1.$ (3)
-For simplicity, we denote the sequence $x_{1},x_{2},...,x_{k}$ by $x_{(1:k)}$.
-We can calculate the observed expected generalized performance, denoted as $V(x_{(1:k)})=\mathbb{E}_{x^{\prime}\in X}\left[U(x^{\prime};x_{(1:k)})\right]$.
-The state at each step $k$ is defined by the best performance for each task, achieved by models trained in earlier stages, represented as $U(x^{\prime};x_{(1:k-1)})$ for each target task $x^{\prime}$.
-The action at each step involves choosing a new task $x_{k}$, aimed at optimizing generalized performance based on current knowledge.
-The cumulative reward is quantified by the observed performance of the combined task sequence, $V(x_{(1:k)})$, providing a direct measure of the strategic value of each decision.
+[…]
+${J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x_{k}},x^{\prime};\pi_{1:k-1}})$ $=\max\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{k},x^{\prime})},{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1},x^{\prime}})\right)\quad\forall x^{\prime}\in X\quad\text{if }k>1.$ (2)
+And ${J}(\pi_{1:1},x)\equiv{J}(\pi_{1},x)$. Then, the overall sequential decision problem can be written as:
+$\max_{x_{k}}\ \ {V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})=\max_{x_{k}}\mathbb{E}_{x^{\prime}\sim\mathcal{U}(X)}\left[{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x_{k}},x^{\prime};\pi_{1:k-1}})\right]\quad\text{s.t.
+}x_{k}\in X\setminus x_{1:k-1}.$ (3)
+The state at each step $k$ is defined by the best generalization performance for each task, achieved by policies trained in earlier stages, represented as $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1},x^{\prime}})$ for each target task $x^{\prime}$.
+The action at each step is choosing a new task $x_{k}$.
 In general, SSTS exhibits stochastic transitions, for example due to randomness in RL training.
-For simplicity, in this work, we assume deterministic transitions; that is, training MDP $x$ will always yield the same performance $J(x)$ and generalization gap $\Delta J(x,x^{\prime}),\forall x^{\prime}\in X$.
-The problem’s horizon is defined by $|X|$, indicating that all models will be trained unless a specific termination condition is met.
-This termination condition is determined by reaching a predefined level of suboptimality.
-Model-based transfer learning (MBTL).
-All methods that solve SSTS are transfer learning methods, by definition.
-We additionally designate a method that solves SSTS to be model-based if an explicit structure is imposed upon $\Delta J(\cdot,\cdot)$.
-To the best of the authors’ knowledge, this is the first work to study such methods in the context of reinforcement learning.
-4 Methods for Model-Based Transfer Learning
-Algorithm 1 Model-based Transfer Learning (MBTL)
-0:CMDPs $c\mathcal{M}(x)$, Task $x\in X$, Transfer budget $K$
-0:$\boldsymbol{\pi}$ and $V$ Initialize : $J,V=0\ \forall x\in X$, $\boldsymbol{\pi}=\{\}$, $k=1$
-1:while $k\leq K$ do
-2:    $x_{k}\leftarrow\textbf{NextTask}(J_{k-1},V_{k-1})$
-3:    $\pi_{k}\leftarrow\text{Train}(\mathcal{M}(x_{k}))$
-4:    $\boldsymbol{\pi}\leftarrow\boldsymbol{\pi}\cup\{\pi_{k}\}$
-5:    Update $J(x),V(x)$
-6:    $k\leftarrow k+1$
-7:end while
-8:return  $\boldsymbol{\pi}$ and $V$
-(a) Pseudo-greedy strategy
-(b) Equidistant strategy
-Figure 3: Naive methods for solving SSTS. Note that the normalized performance of the trained tasks is assumed to always be 1 (known and constant).
+For simplicity, in this work, we assume deterministic transitions; that is, training context-MDP $x$ will always yield the same performance ${J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)$ and generalization gap $\Delta{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}}),\forall x^{\prime}\in X$.
+The problem’s maximum horizon is defined by $|X|$, but can be terminated early if conditions are met (e.g., performance level, training budget).
+4 Model-Based Transfer Learning (MBTL)
+In this section, we introduce an algorithm called Model-based Transfer Learning to solve the SSTS problem.
+MBTL models the generalization performance in two parts: 1) the performance set point, modeled using Gaussian processes, and 2) performance loss (generalization gap), modeled as a linear function of contextual similarity.
+MBTL combines these two pieces of information within a Bayesian optimization (BO) framework to sequentially select training tasks that maximize generalization performance.
+The high-level pseudo-code of the MBTL algorithm is given below.
+MBTL Pseudo-code 1:Input: Task set, Gaussian Process (GP) model, and training budget 2:while training budget remains do 3:    Select next source task $\leftarrow$ Task with the highest acquisition value based on estimated training performance, generalization gap, and generalization performance, determined by the current GP model and tailored acquisition function 4:    Train a policy on the selected source task using reinforcement learning, then evaluate its training performance 5:    Update the GP model with the selected source task training performance to estimate the training performance of tasks in the task set 6:end while 7:Output: Set of policies that maximizes generalized performance across task set
 4.1 Modeling assumptions
-We make key assumptions about the task space, performance functions, and the generalization gap.
-Assumption 1 (Continuity of the task space).
-The task set $X$ is continuous, meaning that for any two tasks $x,x^{\prime}\in X$ and any $\epsilon>0$, there exists a $\delta>0$ such that $\|x-x^{\prime}\|<\delta$ implies the corresponding tasks $x$ and $x^{\prime}$ are indistinguishably close.
-Formally,
-$\forall x,x^{\prime}\in X,\forall\epsilon>0,\exists\delta>0\text{ such that }\|x-x^{\prime}\|<\delta\implies\|J(x)-J(x^{\prime})\|<\epsilon.$
-Assumption 2 (Smoothness of the performance function).
-The performance function $J(x)$, $U(x)$, or $V(x)$ is smooth.
-This implies that these functions are continuously differentiable over $x$.
-Formally, $J(x),U(x),V(x)\in C^{1}(X)$, where $C^{1}(X)$ denotes the space of continuously differentiable functions on $X$.
+We consider a task set $X$ that is continuous and the performance function $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi},x),V(x)$ for a policy $\pi$ to be smooth over the task space $X$.
+In practice, such as control systems, tasks often vary continuously and smoothly rather than abruptly.
+For example, adjusting the angle of a robotic arm by a small amount typically results in a small change in the system and optimal action.
+Inspired by the empirical generalization gap performance as observed in Figure 1, we estimate the generalization gap $\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x^{\prime})$ with a linear function.
 Assumption 1 (Linear generalization gap).
-A linear model is used to model the generalization gap function, formally $\Delta\hat{J}(x,x^{\prime})\simeq\theta|x-x^{\prime}|$, where $\theta$ is the slope of the linear function and $x$ and $x^{\prime}$ are the context of the source task and target task, respectively.
-Assumption 1 and 2 are for implementing MBTL, which relies on the continuous task space and smooth performance function.
-Moreover, we simplify the generalization gap function as a linear function inspired by empirical observations such as shown in Figure 1.
-Such structure (Assumption 3) enables the development of provable algorithms for solving SSTS.
-Algorithm 1 provides a generic psuedocode for a MBTL algorithm.
-The generalization gap structure is leveraged within Line 2 (function NextTask).
-4.2 Naive strategies
-To understand what is desirable in a method for solving SSTS, it is instructive to consider some naive strategies.
-An Equidistant Strategy (ES) samples source tasks so that the tasks are equidistant along the full range of tasks we aim to solve.
-In this method, source tasks are selected to be equidistant within the context space based on the transfer budget $K$.
-Mathematically, the selected source tasks are chosen such that the distances between consecutive tasks are approximately equal, i.e., $x_{k}=\frac{2k-1}{2K}|X|\quad\text{for }k=1,2,\ldots,K.$ Note that this strategy requires advanced knowledge of the total number of training tasks (e.g., training budget).
-A Greedy Strategy (GS) greedily selects the source task that is predicted to have the highest generalized performance in each round, assuming known and constant training performance.
-Figure 3 depicts the difference between ES and GS; observe that the normalized performance of the trained tasks is assumed to be 1.
-However, a critical question emerges concerning scenarios where training performance is unknown or stochastic.
-Furthermore, ES and GS are open-loop strategies; that is, they do not incorporate feedback from actual training outcomes.
+A linear function is used to model the generalization gap function, formally $\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x^{\prime})\simeq\theta|x-x^{\prime}|$, where $\theta$ is the slope of the linear function and $x$ and $x^{\prime}$ are the context of the source task and target task, respectively.
+The relaxation of this assumption can yield additional efficiency benefits but also adds modeling complexity, and thus is an interesting direction for future work.
 4.2 Bayesian optimization
-Bayesian optimization (BO) is a powerful strategy for finding the global optimum of an objective function when obtaining the function is costly or the function itself lacks a simple analytical form [26, 7].
-BO integrates prior knowledge with observations to efficiently decide the next task to train, aiming to optimize the predicted value.
-Gaussian Process.
-Within the framework of BO, we model the source training performance $J_{\text{pred}}(x)$ using Gaussian Process (GP) regression.
-Specifically, the function $J_{\text{pred}}(x)$ is assumed to follow a GP $(J_{\text{pred}}(x)\sim\mathcal{GP}(m(x),k(x,x^{\prime})))$, where $m(x)$ represents the mean function, $\mathbb{E}[J_{\text{pred}}(x)]$, and $k(x,x^{\prime})$ is the covariance function, representing the expected product of deviations of $J_{\text{pred}}(x)$ and $J_{\text{pred}}(x^{\prime})$ from their respective means.
-Let $D_{k-1}$ denote the data observed up to iteration $k-1$, consisting of the pairs $\{(x_{i},V_{\text{obs}}(x_{i}))\}_{i=1,...,k-1}$, where $V_{\text{obs}}(x_{(1:i)})$ are the observed outcomes.
-The posterior prediction of $J_{\text{pred}}$ at a new point $x$, given the data $D_{k-1}$ and previous inputs $x_{(1:k-1)}$, is normally distributed as $P(J_{\text{pred},k}\mid D_{k-1})=\mathcal{N}(\mu_{k}(x),\sigma_{k}^{2}(x))$.
-$\mu_{k}(x)$ and $\sigma_{k}^{2}(x)$ are defined as $\mu_{k}(x)=m(x)+\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{y}$ and $\sigma_{k}^{2}(x)=k(x,x)-\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{k}$, with $\mathbf{k}$ being the vector of covariances between $x$ and each $x_{i}$ in the observed data, i.e., $\mathbf{k}=[k(x,x_{1}),\ldots,k(x,x_{k-1})]$, and $\mathbf{K}$ is the covariance matrix for the observed inputs, defined as $\mathbf{K}=[k(x_{i},x_{j})]_{1\leq i,j\leq k-1}$.
-This formulation enables the GP to update its beliefs about the posterior prediction with every new observation, progressively improving the estimation.
+Bayesian optimization (BO) is a powerful strategy for finding the global optimum of an objective function when obtaining the function is costly, or the function itself lacks a simple analytical form [31, 7].
+BO integrates prior knowledge with observations to efficiently decide the next task to train by using the acquisition function.
+MBTL is a BO method which optimizes for promising source tasks by leveraging Assumption 1 in its acquisition function.
+The role of BO is to approximate ${V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ (see Equation 3) using the data acquired thus far.
+The next source task $x_{k}$ is then selected using this estimate.
+Gaussian process (GP) regression.
+Within the framework of BO, we model the source training performance ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ \forall x\in X\setminus x_{1:k}}$ using Gaussian process (GP) regression.
+Specifically, the function ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ is assumed to follow a GP $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)\sim\mathcal{GP}(\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)],k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})))$, where $k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ is the covariance function, representing the expected product of deviations of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{\tilde{x}},}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ from their respective means.
+Let $D_{k-1}$ denote the data observed up to iteration $k-1$, consisting of the pairs $\{(x_{i},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{i},x_{i})})\}_{i=1,...,k-1}$.
+The estimated performance $\hat{J}_{k}$ after querying $k-1$ samples is updated as more samples are obtained.
+The posterior prediction of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}_{k}}$ at a new point $x$, given the data $D_{k-1}$ and previous inputs ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}}$, is normally distributed as $P(\hat{J}_{k}\mid D_{k-1})=\mathcal{N}(\mu_{k}(x),\sigma_{k}^{2}(x))$.
+$\mu_{k}(x)$ and $\sigma_{k}^{2}(x)$ are defined as $\mu_{k}(x)=\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)]+\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{y}$ and $\sigma_{k}^{2}(x)=k(x,x)-\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{k}$, with $\mathbf{k}$ being the vector of covariances between $x$ and each $x_{i}$ in the observed data, i.e., $\mathbf{k}=[k(x,x_{1}),\ldots,k(x,x_{k-1})]$, and $\mathbf{K}$ is the covariance matrix for the observed inputs, defined as $\mathbf{K}=[k(x_{i},x_{j})]_{1\leq i,j\leq k-1}$.
+This enables the GP to update its beliefs about the posterior prediction with every new observation, progressively improving the estimation.
 Acquisition function.
 The acquisition function plays a critical role in BO by guiding the selection of the next source training task.
-At each decision step $k$, the task $x_{k}$ is chosen by maximizing the acquisition function, as denoted by $x_{k}=\operatorname*{arg\,max}_{x}a(x;x_{(1:k-1)})$.
-Especially in our case, the acquisition function can be designed as the expected marginal improvement of generalized performance across all tasks.
-One effective strategy employed in the acquisition function is the upper confidence bound (UCB) acquisition function, which considers the trade-off between the expected performance of a task based on current models (exploitation) and the measure of uncertainty associated with the task’s outcome (exploration).
+At each decision step $k$, the task $x_{k}$ is chosen by maximizing the acquisition function, as denoted by $x_{k}=\operatorname*{arg\,max}_{x}a(x;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}})$.
+One effective strategy employed in the acquisition function is the upper confidence bound (UCB) acquisition function, which considers the trade-off between the expected performance of a task based on the current task (exploitation) and the measure of uncertainty associated with the task’s outcome (exploration) [40].
+Especially in our case, the acquisition function can be designed as UCB function subtracted by generalization gap and so-far best performance.
 It is defined as follows:
-$a(x;x_{(1:k-1)})=\mathbb{E}_{x^{\prime}\in X}[[\mu_{k-1}(x)+\beta_{k}^{1/2}\sigma_{k-1}(x)-\Delta J(x,x^{\prime})-U(x^{\prime};x_{(1:k-1)})]_{+}]$ (4)
-where $[\cdot]_{+}$ represents $\max(\cdot,0)$ and we can use various forms of $\beta_{k}$.
-Generalized performance.
-We focus on the generalized performance evaluated across all target tasks, defined as $g(x)=\mathbb{E}_{x^{\prime}\in X}[J_{\text{pred}}(x)-\Delta J(x,x^{\prime})]=J_{\text{pred}}(x)-\mathbb{E}_{x^{\prime}\in X}[\Delta J(x,x^{\prime})]$.
+$a(x;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}})=\mathbb{E}_{x^{\prime}\in X}[[\mu_{k-1}(x)+\beta_{k}^{1/2}\sigma_{k-1}(x)-\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})-{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{{1:k-1}},x^{\prime}})]_{+}]$ (4)
+where $[\cdot]_{+}$ represents $\max(\cdot,0)$ and we can use various forms of $\beta_{k}$, which is the trade-off parameter between exploitation and exploration.
 4.3 Regret analysis
 We use regret to quantify the effectiveness of our source task selection based on BO.
-Specifically, we define regret at iteration $k$ as $r_{k}=g(x^{*}_{k})-g(x_{k})$, where $g(x^{*}_{k})$ represents the maximum generalized performance achievable across all tasks, and $g(x_{k})$ is the generalized performance at the current task selection $x_{k}$.
+Specifically, we define regret at iteration $k$ as $r_{k}=\hat{V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})-\hat{V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$, where $\hat{V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ represents the maximum generalization performance achievable across all tasks, and $\hat{V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ is the performance at the current task selection $x_{k}$.
 Consequently, the cumulative regret after $K$ iterations is given by $R_{K}=\sum_{k=1}^{K}r_{k}$, summing the individual regrets over all iterations.
@@
 Mathematically, we aim to prove that $\lim_{K\rightarrow\infty}\frac{R_{K}}{K}=0$, indicating that, on average, the performance of our strategy approaches the optimal performance as the number of iterations increases.
-Regret of MBTL-GP.
-Having established the general framework for regret analysis, we now turn our attention to the specific regret properties of our MBTL-GP algorithm.
-To analyze the regret of MBTL-GP, consider the scaling factor for the UCB acquisition function given by $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$ in Equation 4.
+Regret of MBTL.
+Having established the general framework for regret analysis, we now turn our attention to the specific regret properties of our MBTL algorithm.
+To analyze the regret of MBTL, consider the scaling factor for the UCB acquisition function given by $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$ in Equation 4.
 It is designed to achieve sublinear regret with high probability, aligning with the theoretical guarantees outlined in Theorem 1 and 5 from [40].
@@
 In this section, we demonstrate that strategic reduction of the possible sets, guided by insights from previous task selections or source task training performance, leads to significantly tighter regret bounds than Theorem 1.
-By focusing on the most promising regions of the task space, our approach enhances learning efficiency and maximizes the model’s performance and applicability.
+By focusing on the most promising regions of the task space, our approach enhances learning efficiency and maximizes the policy’s performance and applicability.
 Given the generalization gap observed in Figure 1, we observe that performance loss decreases as the context similarity increases.
-We model the degradation from the source task using a simple linear function in Assumption 3.
+We model the degradation from the source task using a linear function in Assumption 1.
 Training on the source task can solve a significant portion of the remaining tasks.
@@
 If the source task selected in the previous steps could solve the remaining target task sufficiently, we can eliminate the search space at a desirable rate.
-Consequently, at each step, we effectively reduce the search space.
-Given the generalization gap observed in Figure 1, we model the degradation from the source task using a simple linear function in Assumption 3.
+Formally, we can define the search space at $k$-th iteration as follows:
+Definition 2 (Search space).
+We define the search space $X_{k}$ at iteration $k$ as a subset of $X$, with each element $x^{\prime}\in X_{k}$, such that $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1},x^{\prime}})\leq\hat{J}(\pi_{x_{k}},x^{\prime}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\Delta J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x_{k}},x^{\prime}})$.
+Given the generalization gap observed in Figure 1, we model the degradation from the source task using a linear function in Assumption 1.
 While the figure might not strictly appear linear, the linear approximation simplifies analysis and is supported by empirical observations.
@@
 Consequently, at each step, we effectively focus on a reduced search space.
-Definition 2 (Reduced search space).
-We define the reduced search space $X_{k}$ as a subset of $X$, with each element $x^{\prime}\in X_{k}$, such that $U(x^{\prime};x_{(1:k-1)})\leq\hat{J}(x_{k})+\Delta J(x^{\prime},x_{k})$.
 We leverage the reduced uncertainty in well-sampled regions to tighten the regret bound while slightly lowering the probability $\delta$ in Theorem 1.
@@
 Consider $|X_{k}|=\frac{1}{\sqrt{k}}|X|$. The regret bound would be $R_{K}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\log K}$ with a probability of at least $1-\delta^{\prime}$.
-Figure 4: Empirical results of the restriction of search space by MBTL-GP compared to two examples from Corollaries 2.1 and 2.2.
+Figure 3: Empirical results of the restriction of search space by MBTL compared to two examples from Corollaries 2.1 and 2.2.
 In cases where the search space is defined using MBTL-GS, the largest segment’s length would reduce geometrically, described by $|X_{k}|\leq 2^{-\lfloor\log_{2}k\rfloor}|X|$.
@@
 The regret bound for the $|X_{k}|\leq 2^{-\lfloor\log_{2}k\rfloor}|X|$ would be $R_{K}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\pi^{2}/6}$ with a probability of at least $1-\delta^{\prime}$.
-Proofs for Corollaries 2.1 and 2.2 are provided in Appendix A.1.3 and A.1.4, respectively.
-Based on our experiments presented in Section 5, the rate of elimination of the largest segment for MBTL-GP is shown in Figure 4.
+Proofs for Corollaries 2.1 and 2.2 are provided in Appendix A.3.2 and A.3.3, respectively.
+Based on our experiments presented in Section 5, the rate of elimination of the largest segment for MBTL is shown in Figure 3.
 5 Experiments and analysis
 5.1 Setup
-Our experiments span a variety of CMDP settings, focusing on transportation tasks as well as classic control benchmarks.
-We introduce three intelligent transportation tasks that utilize DRL but suffer from brittleness issues, as well as the classic control tasks.
+Our experiments consider CMDPs that span standard and real-world benchmarks.
+In particular, we consider standard continuous control benchmarks from the CARL library [6].
+In addition, we study problems from RL for intelligent transportation systems, using [47] to model the CMDPs.
+Surprisingly, despite the relatively low complexity of the CMDPs considered, standard deep RL algorithms appear to struggle to solve the tasks.
 Baselines.
-Our proposed methods are compared against a range of baselines including (1) Random selection, where each training task is chosen randomly; (2) Exhaustive training, which involves independent training separate models on all tasks exhaustively but is computationally expensive; (3) Multi-task RL, where a single context-conditioned model is trained for all tasks; (4) Oracle transfer, which assumes full knowledge of generalized performance for all policies and selects the best source task for each target task.
-Proposed methods.
-Our proposed methods are: (1) MBTL-ES, selecting source tasks for the context to be equidistance based on the transfer budget; (2) MBTL-GS, selecting the next source task greedily to maximize the marginal improvement of expected generalized performance based on linear generalization gap (Assumption 3); (3) MBTL-GP, selecting source task based on the BO using GP and UCB with scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$.
+We consider two types of baselines when evaluating our proposed algorithm: canonical and multi-policy.
+The canonical baselines are selected to validate multi-policy training; the multi-policy training baselines are heuristic methods designed to validate the Bayesian optimization approach.
+The canonical baselines include: (1) Independent training, which involves independently training separate models on each task; and (2) Multi-task RL, where a single “universal” context-conditioned policy is trained for all tasks.
+The multi-policy baselines include: (3) Random selection, where each training task is chosen uniformly at randomly; (4) Equidistant strategy, which selects training tasks by equally subdividing the context space based on a given training budget, and then trains them in lexicographical order; (5) Greedy strategy, which greedily selects the next source task based on Assumption 1 and fixed training performance; and (6) Sequential Oracle transfer, which has access to generalized performance corresponding to policies for all tasks (including those not yet selected) and uses that information to greedily select the best source task at each step.
+Proposed method. We evaluate MBTL with the scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$.
 DRL algorithms and performance measure.
-We utilize Deep Q-Networks (DQN) for discrete actions [24] and Proximal Policy Optimization (PPO) for continuous actions [30].
-We evaluate our methods by the average performance across all target tasks after training 15 source tasks and the number of trained models needed to achieve a certain level of suboptimality.
-We employ min-max normalization of the rewards for each task.
-5.2 Traffic tasks
-Table 1: Comparative performance of different methods on traffic CMDP tasks
-Benchmark Baselines MBTL (Ours) Oracle
-Task Context Variation Random Exhaustive Multitask MBTL-GS MBTL-ES MBTL-GP Oracle Transfer
-Number of Trained Models $k$ $N$ $1$ $k$ $K$ $k$ $N$
-Traffic Signal Road Length 0.9249 0.9409 0.8242 0.9278 0.9213 0.9371 0.9409
-Traffic Signal Inflow 0.8457 0.8646 0.8319 0.8496 0.8700 0.8673 0.8768
-Traffic Signal Speed Limit 0.8821 0.8857 0.6083 0.8862 0.8858 0.8854 0.8876
-Eco-Driving Penetration Rate 0.5959 0.5260 0.1945 0.5827 0.5934 0.6323 0.6660
-Eco-Driving Inflow 0.4774 0.4061 0.2229 0.4673 0.4705 0.5108 0.5528
-Eco-Driving Green Phase 0.4406 0.3850 0.4228 0.4431 0.4557 0.4700 0.5027
-AA-Ring-Acc Hold Duration 0.8924 0.8362 0.9209 0.8776 0.9057 0.9242 0.9552
-AA-Ring-Vel Hold Duration 0.9785 0.9589 0.9720 0.9807 0.9772 0.9816 0.9822
-AA-Ramp-Acc Hold Duration 0.6050 0.4276 0.5158 0.6143 0.5956 0.6318 0.7111
-AA-Ramp-Vel Hold Duration 0.6690 0.5473 0.5034 0.5907 0.6787 0.7182 0.7686
-Average 0.7312 0.6778 0.6017 0.7220 0.7354 0.7559 0.7844
-†Higher the better. The bold indicates the best for each benchmark, excluding the Oracle transfer. Detailed results with variance for each method are provided in Appendix A.2.2.
-‡AA: Advisory autonomy tasks, Ring: Single lane ring, Ramp: Highway ramp, Acc: Acceleration guidance, Vel: Speed guidance.
-We validate our methods on diverse simulated traffic scenarios, focusing on to what extent Our proposed methods can optimize the global objective only with a small number of trained models.First, while most traffic lights operate on fixed schedules, we can design learning-based adaptive (1) Traffic signal control to optimize the traffic [8, 21].
-However, considering that every intersection looks different, challenges persist in generalizing across various intersection configurations [18].
-Given the significant portion of greenhouse gas emissions in the United States coming from the transportation sector [11], the second traffic task is (2) Eco-driving Lagrangian control, which is critical for climate change mitigation.
-DRL-based eco-driving strategies have been developed [13, 38, 17] but still have some issues of difficulties in generalization.
-The last traffic task is (3) Advisory autonomy, which is a real-time speed advisory system that enables human drivers to emulate the system-level performance of autonomous vehicles in mixed autonomy systems [34, 15].
-Instead of direct and instantaneous control, human drivers receive periodic guidance, which varies based on the traffic states.
-In Appendix A.2, we provide comprehensive details about our experiments.
-Figure 5: Comparison of normalized generalized performance of all target tasks.
-Discussion.
-Table 1 compares the min-max normalized generalized performance across the full target task range.
-While exhaustive training is computationally expensive and exhibits suboptimal performance, the Oracle transfer, which assumes perfect knowledge of transferability, demonstrates the potential benefits of zero-shot transfer for solving the complete range of multiple tasks.
-Our proposed MBTL methods leverage this transferability by intelligently selecting the next source task to learn from.
-Notably, MBTL-GP shows superior performance, closely aligning with the Oracle transfer, which suggests the Gaussian Process effectively models the underlying task dynamics, facilitating better generalization from learned experiences.
-For example, MBTL-GP significantly improves generalized performance with a few source tasks for traffic signal control problems varying green phase durations (Figure 5).
+We utilize Deep Q-Networks (DQN) for discrete action spaces [29] and Proximal Policy Optimization (PPO) for continuous action spaces [35].
+We evaluate our method by the average performance across all $N$ target tasks after training up to $K=$15 source tasks or the number of source tasks needed to achieve a certain level of performance.
+We employ min-max normalization of the rewards for each task, and we provide comprehensive details about our model in Appendix A.4.1.
+5.2 Traffic benchmark experiments
+Table 1: Comparative performance of different methods on traffic CMDPs
+Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
+Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
+Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
+Traffic Signal Inflow 0.8646 0.8319 0.8457 0.8700 0.8496 0.8673 0.8768
+Traffic Signal Speed Limit 0.8857 0.6083 0.8821 0.8858 0.8862 0.8854 0.8876
+Eco-driving Penetration Rate 0.5260 0.1945 0.5959 0.5934 0.5827 0.6323 0.6660
+Eco-driving Inflow 0.4061 0.2229 0.4774 0.4705 0.4673 0.5108 0.5528
+Eco-driving Green Phase 0.3850 0.4228 0.4406 0.4557 0.4431 0.4700 0.5027
+AA-Ring-Acc Hold Duration 0.8362 0.9209 0.8924 0.9057 0.8776 0.9242 0.9552
+AA-Ring-Vel Hold Duration 0.9589 0.972 0.9785 0.9772 0.9807 0.9816 0.9822
+AA-Ramp-Acc Hold Duration 0.4276 0.5158 0.6050 0.5956 0.6143 0.6318 0.7111
+AA-Ramp-Vel Hold Duration 0.5473 0.5034 0.669 0.6787 0.5907 0.7182 0.7686
+Average 0.6778 0.6017 0.7312 0.7354 0.722 0.7559 0.7844
+†Higher the better.
+Bold values represent the highest value(s) within the statistically significant range for each CMDP, excluding the oracle.
+Detailed results with variance for each method are provided in Appendix A.4.2.
+‡AA: Advisory autonomy benchmark, Ring: Single lane ring, Ramp: Highway ramp, Acc: Acceleration guidance, Vel: Speed guidance.
+We consider three traffic control domains.
+First, while most traffic lights still operate on fixed schedules, RL can be used to design adaptive (1) Traffic signal control to optimize traffic [9, 23].
+However, considering that every intersection is different, challenges persist in generalizing across intersection configurations [19].
+Given the significant portion of greenhouse gas emissions in the United States due to transportation [12], the second traffic domain is (2) Dynamic eco-driving at signalized intersections, which concerns learning energy-efficient driving strategies in urban settings.
+DRL-based eco-driving strategies have been developed [14, 45, 18] but still experience difficulties in generalization.
+Our final traffic domain is (3) Advisory autonomy, in which real-time speed or acceleration advisories guide human drivers to act as vehicle-based traffic controllers in mixed traffic environments [39, 8, 16].
+The context space $X$ is discretized into $N=\{50,50,40\}$ contexts for the three domains, respectively.
+In Appendix A.4, we provide details about our experiments.
+Figure 4: Traffic CMDP results.
+Method comparison of normalized performance over $N$ tasks.
+MBTL efficiently selects source training tasks.
+The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 25x fewer samples needed.
+Results.
+Table 1 and Figure 4 summarize the results.
+Notably, the Oracle far outperforms the standard baselines (Independent and Multi-task), indicating the potential for transfer learning and intelligent training of multiple models, respectively.
+MBTL rapidly approaches the Oracle, within $\approx 10$ transfer steps, indicating that the Gaussian process effectively models the training performance and linear generalization gap models generalization performance.
+It is important to note that multi-task RL studies commonly consider Independent training as a strong baseline due to its avoidance of negative transfer and other training instability issues.
+Indeed, Independent often (but not always) outperforms Multi-task in our experiments.
+Yet, our experiments show that MBTL consistently outperforms both Independent and Multi-task and matches their performance with 5-25x better sample efficiency.
+Among the sequential baselines, MBTL consistently outperforms the heuristic baselines, indicating the value of adaptively selecting source tasks based on feedback.
+The sequential baselines, such as random, equidistant, and greedy strategy, also generally outperform Independent and Multi-task, indicating the general value of multi-policy training for solving CMDPs.
 More results are provided in Appendix A.4.
-5.3 Standard control benchmarks
-Table 2: Comparative performance of different methods on standard control CMDP tasks
-Benchmark Baselines MBTL (Ours) Oracle
-Task Context Variation Random Exhaustive Multitask MBTL-GS MBTL-ES MBTL-GP Oracle Transfer
-Number of Trained Models $k$ $N$ $1$ $k$ $K$ $k$ $N$
-Pendulum Length 0.7270 0.7383 0.6830 0.7327 0.7092 0.7697 0.7969
-Pendulum Mass 0.6329 0.6237 0.5793 0.6408 0.6092 0.6827 0.7132
-Pendulum Timestep 0.7989 0.8135 0.7247 0.8177 0.7541 0.8141 0.8801
-Cartpole Mass of Cart 0.7221 0.9466 0.7153 0.6501 0.7516 0.8212 0.9838
-Cartpole Length of Pole 0.8121 0.9110 0.5441 0.8217 0.8428 0.9124 0.9875
-Cartpole Mass of Pole 0.8858 0.9560 0.6073 0.8744 0.7909 0.9351 1.0000
-BipedalWalker Gravity 0.9330 0.9281 0.7898 0.9359 0.9494 0.9393 0.9674
-BipedalWalker Friction 0.9650 0.9317 0.9051 0.9645 0.9664 0.9713 0.9778
-BipedalWalker Scale 0.8605 0.8694 0.7452 0.8792 0.8496 0.8886 0.9107
-HalfCheetah Gravity 0.8542 0.6679 0.6292 0.8663 0.8634 0.9073 0.9544
-HalfCheetah Friction 0.8567 0.6693 0.7242 0.8591 0.8703 0.9274 0.9663
-HalfCheetah Stiffness 0.8533 0.6561 0.7007 0.8785 0.7817 0.9146 0.9674
-Average 0.8251 0.8093 0.6957 0.8267 0.8116 0.8736 0.9255
-†Higher the better. The bold indicates the best for each benchmark, excluding the Oracle transfer. Detailed results with variance for each method are provided in Appendix A.2.2.
-Building on the evaluation of a suite of traffic-related tasks, we extend our investigation to classic control tasks to assess the applicability of our methods across a broader spectrum of environments.
-We utilize context-extended versions of standard RL environments from CARL benchmark library [6] to rigorously test the robustness and adaptability of our methods under varied contexts.
-For Cartpole tasks, we explored contextual Markov decision processes (CMDPs) with varying cart masses, pole lengths, and pole masses.
-We vary the timestep duration, pendulum length, and pendulum mass in Pendulum.
+5.3 Continuous control benchmark experiments
+Table 2: Comparative performance of different methods on standard control CMDPs
+Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
+Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
+Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
+Pendulum Length 0.7383 0.6830 0.7270 0.7092 0.7327 0.7697 0.7969
+Pendulum Mass 0.6237 0.5793 0.6329 0.6092 0.6408 0.6827 0.7132
+Pendulum Timestep 0.8135 0.7247 0.7989 0.7541 0.8177 0.8141 0.8801
+Cartpole Mass of Cart 0.9466 0.7153 0.7221 0.7516 0.6501 0.8212 0.9838
+Cartpole Length of Pole 0.9110 0.5441 0.8121 0.8428 0.8217 0.9124 0.9875
+Cartpole Mass of Pole 0.9560 0.6073 0.8858 0.7909 0.8744 0.9351 1.0000
+BipedalWalker Gravity 0.9281 0.7898 0.9330 0.9494 0.9359 0.9393 0.9674
+BipedalWalker Friction 0.9317 0.9051 0.965 0.9664 0.9645 0.9713 0.9778
+BipedalWalker Scale 0.8694 0.7452 0.8605 0.8496 0.8792 0.8886 0.9107
+HalfCheetah Gravity 0.6679 0.6292 0.8542 0.8634 0.8663 0.9073 0.9544
+HalfCheetah Friction 0.6693 0.7242 0.8567 0.8703 0.8591 0.9274 0.9663
+HalfCheetah Stiffness 0.6561 0.7007 0.8533 0.7817 0.8785 0.9146 0.9674
+Average 0.8093 0.6957 0.8251 0.8116 0.8267 0.8736 0.9255
+†Higher the better.
+Bold values represent the highest value(s) within the statistically significant range for each CMDP, excluding the oracle.
+Detailed results with variance for each method are provided in Appendix A.4.2.
+Figure 5: Continuous control CMDP results.
+Method comparison of normalized performance over $N$ tasks.
+The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 50x fewer samples needed.
+To probe the generality of MBTL, we utilize context-extended versions of standard RL environments from CARL benchmark library [6] to evaluate our methods under varied contexts.
+For the Cartpole domain, we considered CMDPs with varying cart mass, pole length, and pole mass.
+In Pendulum, we vary the timestep duration, pendulum length, and pendulum mass.
 The BipedalWalker was tested under varying friction, gravity, and scale.
-In HalfCheetah tasks, we manipulated friction, gravity, and stiffness parameters.
+In HalfCheetah domain, we manipulated friction, gravity, and stiffness parameters.
 These variations critically influence the dynamics and physics of the environments.
-The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times, enabling a comprehensive analysis of transfer learning under drastically different conditions.
+The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times ($N=100$), enabling a comprehensive analysis of transfer learning under drastically different conditions.
 We provide more details in Appendix A.4.
-Discussion.
-Table 2 compares the performance across the full target task range.
-Our experiments demonstrate that MBTL significantly improves performance and sample efficiency in dynamic urban transportation systems compared to the baselines.
-In control tasks such as Cartpole, Pendulum, BipedalWalker, and HalfCheetah, where physical parameters like mass directly influence the dynamics, MBTL methods, particularly MBTL-GP, excel by effectively transferring knowledge between similar task variations.
-For example, in Figure 5, MBTL-GP outperforms other baselines, including multitask training.
-The results underscore the potential of MBTL in solving multiple related tasks.
-Figure 6: Sensitivity analysis on different DRL algorithms (Cartpole, length of pole).
-Figure 7: Sensitivity analysis on the different acquisition functions.
+Figure 6: The GP sequentially updates estimates of the performance function (blue) based on previously trained models.
+Then, MBTL selects the next source task that maximizes the acquisition function (red).
+(CMDP: Pendulum (Time step)).
+Results.
+The results summarized in Table 2 demonstrate MBTL’s sample efficiency and competitive performance across diverse control domains, often closely trailing the Oracle only with a small number of tasks.
+With the exception of the Cartpole domain, MBTL generally shows superior performance.
+Figure 5 shows that MBTL demonstrates rapid learning curves, reaching high normalized performance with relatively few samples.
+Specifically, Figure 6 illustrates the detailed process of how MBTL utilizes GP for performance estimation and chooses the next source task that maximizes the acquisition function that evaluates the expected improvement of generalized performance.
+MBTL achieves comparable performance to multi-task or independent baselines with 8-50x fewer samples, highlighting its effectiveness in reducing training requirements.
+Figure 7: Sensitivity analysis on the DRL algorithm underlying MBTL (DQN, PPO, and A2C), tested on Cartpole with varying length of pole. MBTL remains effective.
+Figure 8: Sensitivity analysis on acquisition functions.
 5.3.1 Sensitivity analysis
-DRL algorithms.
-Figure 6 shows an ablation study comparing different DRL algorithms—DQN, PPO, and Advantage Actor-Critic (A2C) [25].
-The heatmap results show differences in transferability, and the results underscore the robustness of MBTL to different algorithms.
+DRL algorithms. Figure 7 shows that MBTL remains effective with different underlying DRL algorithms—DQN, PPO, and Advantage Actor-Critic (A2C) [30]—used for single-task training.
 Acquisition functions.
@@
 While expected improvement (EI) focuses on promising marginal gains beyond the current best, UCB utilizes both mean and variance for balancing exploration and exploitation.
-For Cartpole and advisory autonomy tasks, UCB with decreasing $\beta_{k}$ as $k$ increases shows the best performance.
-However, for eco-driving control, the $\beta_{k}$ formulation in Section 4.4 exhibits the highest performance, though the difference is not statistically significant.
+Overall, we find that MBTL is not particularly sensitive to the choice of optimism representation in the acquisition function, which indicates that MBTL has a weak dependence on hyperparameters.
 6 Related work
+Contextual Reinforcement Learning.
 Robustness and generalization challenges in DRL are generally addressed by a few common techniques in the literature.
@@
 When they are visible, they can be directly used as side information by conditioning the policy on them [38].
-A common approach is multi-task learning, which aims to leverage a common structure between tasks.
+Multi-task learning.
+A common approach for solving CRL is multi-task learning, which aims to leverage common structure between tasks.
 The structure between tasks has been exploited using policy sketches for task structuring [2] and by sharing a distilled policy that captures common behavior across tasks [44].
-However, when the context features are not visible, the resulting CMDP becomes a partially observable CMDP [20, 9] and can be challenging for multi-task learning.
+One of the known challenges in multi-task learning is that when the context features are not visible, the resulting CMDP becomes a partially observable CMDP [22, 10] and can be challenging for multi-task learning.
+Moreover, another challenge is negative transfer, where training from tasks that are too different results in training instability or failure [21, 41, 42].
 In this work, we use multi-task learning as a baseline in evaluating our methods.
+Zero-shot transfer and policy reuse.
 Zero-shot transfer is another commonly used technique that focuses on adapting models trained in one environment to perform effectively in unseen settings without further training [22].
+Li et al.
+[25] introduces context-aware policy reuse, which adapts multiple source policies based on context, enhancing generalization in complex environments.
 Rezaei-Shoshtari et al.
@@
 Sinapov et al.
-[32] uses meta-data to learn inter-task transferability to learn the expected benefit of transfer given a source-target task pair.
+[37] use meta-data to learn inter-task transferability to learn the expected benefit of transfer given a source-target task pair.
+Bao et al.
+[3] propose a metric for evaluating transferability based on information-theoretic feature representations across tasks.
 Another popular class of techniques for robust and generalizable DRL is domain adaptation strategies.
@@
 However, our focus in this work centered around in-distribution generalization, whereas domain adaptation techniques are commonly used for out-of-distribution generalization.
+Source task selection.
+In the context of transfer learning, selecting appropriate source tasks is crucial.
+Li and Zhang [24] proposes an optimal online method for dynamically selecting the most relevant single source policy in reinforcement learning.
+Beyond RL, Meiseles and Rokach [28] emphasizes structural alignment in time-series source models to prevent performance degradation, while Poth et al.
+[33] finds that selecting aligned intermediate tasks in natural language processing boosts transfer effectiveness.
+Together, these studies show the importance of context-aligned source tasks selection to improve transferability and reduce negative transfer.
 7 Conclusion
-Our study refines the robustness of DRL by developing MBTL using zero-shot transfer.
-MBTL strategically selects a set of source tasks to maximize overall performance and minimize training costs.
-It models the generalization gap and uses BO to predict task performance efficiently.
-We theoretically show that the method exhibits regret that is sublinear in the number of training tasks.
-We validate MBTL with urban traffic and control benchmarks.
-Limitation is that MBTL relies highly on assumptions that may not hold in highly dynamic or complex environments, potentially limiting its applicability and leading to suboptimal policy convergence.
-Future work will focus on improving the modeling of the generalization gap and exploring scenarios with multiple context variations where simple strategies like MBTL-ES or MBTL-GS may not be applicable.
+This study introduces a method called Model-based Transfer Learning (MBTL), which layers on top of existing RL methods to effectively solve CMDPs.
+Rather than independent or multi-task training, which trains $N$ or 1 models, respectively, MBTL intelligently selects an intermediate number of models to train.
+MBTL has two key components: an explicit model of the generalization gap and a Gaussian process component to estimate training performance.
+MBTL achieves up to 50x improved sample efficiency on standard and real-world benchmarks.
+Furthermore, MBTL achieves sublinear regret in the number of training tasks.
+A limitation is that MBTL is designed for a single-dimensional context variation.
+Promising directions of future work include studying high-dimensional context spaces and out-of-distribution generalization, as well as the development of new real-world CMDP benchmarks.
+Acknowledgments and Disclosure of Funding
+The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the research results reported within this paper.
+This work was supported by the National Science Foundation (NSF) CAREER award (#2239566), the Kwanjeong Educational Foundation Ph.D.
+scholarship program, and an Amazon Robotics Ph.D.
+Fellowship.
+The authors would like to thank the anonymous reviewers for their valuable feedback.
 Appendix A Appendix
 […]
-A.1 Theoretical analysis
-A.1.1 Notation
+A.1 Notation
 Table 3 describes the notation used in this paper.
@@
 $x^{\prime}$ Target task ($x^{\prime}\in X$)
+${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}}$ Trained policy from source task ($x\in X$)
 $x_{k}$ Selected source task at transfer step $k$ ($k=1,...,K$)
-$\mathbf{x}$ Vector of source tasks ($\mathbf{x}=[x_{1},...,x_{k}]$ where $x_{i}\in X$ for $1\leq i\leq k$)
-$c\mathcal{M}(x)$ Contextual MDP parameterized by $x$
-$J(x)$ Performance of task $c\mathcal{M}(x)$
-$\Delta\hat{J}(x,x^{\prime})$ Generalization gap (source: $x$, target: $x^{\prime}$)
-$U(x^{\prime};x)$ Transferred performance (source: $x$ (or $\mathbf{x}$), target: $x^{\prime}$)
-$V(x)$ Expected generalized performance of source model $x$ evaluated on all $x^{\prime}\in X$
+$\mathcal{M}_{x}$ Contextual MDP parameterized by $x$
+$J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ Performance of task $\mathcal{M}_{x}$
+${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{x},x^{\prime})}$ Generalization performance (source: $x$ (or $\mathbf{x}$), target: $x^{\prime}$)
+$\Delta{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x^{\prime})$ Generalization gap (source: $x$, target: $x^{\prime}$)
+$V(x^{\prime};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}})$ Expected generalization performance of source model $x$ evaluated on all $x^{\prime}\in X$
 Figure 9 helps understand the discrepancy between the observed generalized performance and the predicted one.
-Figure 9 illustrates how to calculate the marginal improvement of expected generalized performance ($\hat{V}(x;x_{1},...,x_{k-1})-V(x_{1},...,x_{k-1})$).
+Figure 10 illustrates how to calculate the marginal improvement of expected generalized performance ($\hat{V}(x;x_{1:k-1})-V(x;x_{1:k-1})$).
 […]
-Figure 8: Illustration of the discrepancy between observed ($U$) and predicted ($\hat{U}$) generalized performance after training on source task $x_{1}$ and attempting zero-shot transfer to $x^{\prime}$.
+Figure 9: Illustration of the discrepancy between observed ($J$) and predicted ($\hat{J}$) generalized performance after training on source task $x_{1}$ and attempting zero-shot transfer to $x^{\prime}$.
 […]
-Figure 9: Step for choosing $x_{2}$ that maximizes the estimated marginal improvement ($\hat{V}(x;x_{1})-V(x_{1})$).
-$\hat{V}(x;x_{1})$ corresponds to the red area under the red line and $V(x_{1})$ as the area under $U(x^{\prime};x_{1})$.
-A.1.2 Proof of Theorem 2
+Figure 10: Step for choosing $x_{2}$ that maximizes the estimated marginal improvement ($\hat{V}(x;\pi_{1})-V(x_{1})$).
+$\hat{V}(x;\pi_{1})$ corresponds to the red area under the red line and $V(x_{1})$ as the area under $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1},x^{\prime}})$.
+A.2 Model-Based Transfer Learning (MBTL) Algorithm
+[…]
+A.3 Theoretical analysis
+A.3.1 Proof of Theorem 2
 See 2
@@
 $R_{K}=\sum_{k=1}^{K}r_{k}\leq\sqrt{\sum_{k=1}^{K}r_{k}\left(\frac{|X|}{|X_{k}|}\right)^{2}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}.$ (5)
-A.1.3 Proof of Corollary 2.1
+A.3.2 Proof of Corollary 2.1
 See 2.1
@@
 $R_{K}=\sum_{k=1}^{K}r_{k}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\log K}.$ (7)
-A.1.4 Proof of Corollary 2.2
+A.3.3 Proof of Corollary 2.2
 See 2.2
@@
 $R_{K}=\sum_{k=1}^{K}r_{k}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}\leq\sqrt{\frac{C_{1}\beta_{K}\gamma_{K}\pi^{2}}{6}}.$ (9)
-A.2 Experiment details
-A.2.1 Details about Gaussian process (GP)
+A.4 Experiment details
+A.4.1 Details about Gaussian process (GP)
 In our study, we conducted hyperparameter tuning experiments for Gaussian Process (GP) regression to optimize its performance.
-Specifically, we varied the noise standard deviation over the set $\{0.001,0.01,0.1,1\}$ and the number of restarts for the optimizer over the set $\{5,10,15\}$.
-We explored several kernel configurations, including the combination of a constant kernel (C) and a radial basis function kernel (RBF) with a length scale ranging from 0.01 to 100, as well as more complex kernels incorporating a white noise kernel and an exponential sine squared kernel with varying length scales and periodicities.
-Each GP model was instantiated with the specified kernel, and the noise level was set to the square of the noise standard deviation.
-This systematic exploration aimed to identify the optimal set of hyperparameters that could enhance the model’s predictive accuracy.
-A.2.2 Results of table with standard deviation
-Table 4: Comparative performance of different methods on context-variant traffic and benchmark tasks
-Benchmark Baselines Model-Based Transfer Learning Oracle
-Task Context Variation Random Exhaustive Multitask MBTL-PS MBTL-ES MBTL-GP Oracle Transfer
-Number of Trained Models $k$ $N$ $1$ $k$ $K$ $k$ $N$
+Specifically, we varied the noise standard deviation over the set $\{0.001,0.01,0.1,1\}$ and the number of restarts for the optimizer over the set $\{5,6,7,8,9,10,11,12,13,14,15\}$ with several kernel configurations.
+We set noise standard deviation as $0.1$, the number of restarts for the optimizer over the set as $9$, and the kernel as the combination of a constant kernel (C) with the bound from 0.001 to 1000 and a radial basis function kernel (RBF) with a length scale ranging from 0.01 to 100.
+The noise level was set to the square of the noise standard deviation ($0.1^{2}$).
+A.4.2 Results of table with standard deviation
+Table 4: Comparative performance of different methods on context-variant traffic and benchmark CMDPs
+Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
+Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
+Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
 Traffic Signal Road Length
-0.9249
 0.9409
 0.8242
+0.9249
+0.9213
 0.9278
-0.9213
 0.9371
@@
 Traffic Signal Inflow
-0.8457
 0.8646
 0.8319
+0.8457
+0.8700
 0.8496
-0.8700
 0.8673
@@
 Traffic Signal Speed Limit
-0.8821
 0.8857
 0.6083
+0.8821
+0.8858
 0.8862
-0.8858
 0.8854
@@
 Eco-Driving Penetration Rate
-0.5959
 0.5260
 0.1945
+0.5959
+0.5934
 0.5827
-0.5934
 0.6323
@@
 Eco-Driving Inflow
-0.4774
 0.4061
 0.2229
+0.4774
+0.4705
 0.4673
-0.4705
 0.5108
@@
 Eco-Driving Green Phase
-0.4406
 0.3850
 0.4228
+0.4406
+0.4557
 0.4431
-0.4557
 0.4700
@@
 AA-Ring-Acc Hold Duration
-0.8924
 0.8362
 0.9209
+0.8924
+0.9057
 0.8776
-0.9057
 0.9242
@@
 AA-Ring-Vel Hold Duration
-0.9785
 0.9589
 0.9720
+0.9785
+0.9772
 0.9807
-0.9772
 0.9816
@@
 AA-Ramp-Acc Hold Duration
-0.6050
 0.4276
 0.5158
+0.6050
+0.5956
 0.6143
-0.5956
 0.6318
@@
 AA-Ramp-Vel Hold Duration
-0.6690
 0.5473
 0.5034
+0.6690
+0.6787
 0.5907
-0.6787
 0.7182
@@
 Pendulum Length
-0.7270
 0.7383
 0.6830
+0.7270
+0.7092
 0.7327
-0.7092
 0.7697
@@
 Pendulum Mass
-0.6329
 0.6237
 0.5793
+0.6329
+0.6092
 0.6408
-0.6092
 0.6827
@@
 Pendulum Timestep
-0.7989
 0.8135
 0.7247
+0.7989
+0.7541
 0.8177
-0.7541
 0.8141
@@
 Cartpole Mass of Cart
-0.7221
 0.9466
 0.7153
+0.7221
+0.7516
 0.6501
-0.7516
 0.8212
@@
 Cartpole Length of Pole
-0.8121
 0.9110
 0.5441
+0.8121
+0.8428
 0.8217
-0.8428
+(0.0980 )
 0.9124
@@
 Cartpole Mass of Pole
-0.8858
 0.9560
 0.6073
+0.8858
+0.7909
 0.8744
-0.7909
 0.9351
@@
 BipedalWalker Gravity
-0.9330
 0.9281
 0.7898
+0.9330
+0.9494
 0.9359
-0.9494
 0.9393
@@
 BipedalWalker Friction
-0.9650
 0.9317
 0.9051
+0.9650
+0.9664
 0.9645
-0.9664
 0.9713
@@
 BipedalWalker Scale
-0.8605
 0.8694
 0.7452
+0.8605
+0.8496
 0.8792
-0.8496
 0.8886
@@
 HalfCheetah Gravity
-0.8542
 0.6679
 0.6292
+0.8542
+0.8634
 0.8663
-0.8634
 0.9073
@@
 HalfCheetah Friction
-0.8567
 0.6693
 0.7242
+0.8567
+0.8703
 0.8591
-0.8703
 0.9274
@@
 HalfCheetah Stiffness
-0.8533
 0.6561
 0.7007
+0.8533
+0.7817
 0.8785
-0.7817
 0.9146
@@
 Average
-0.8068
 0.7775
 0.6337
+0.8068
+0.7996
 0.8037
-0.7996
 0.8435
 0.8833
-* Note: The bold values indicate the maximum value for each task excluding the oracle. Standard deviation across multiple runs in the parenthesis.
-A.2.3 Details about traffic signal control
+* Note: Bold values represent the highest value(s) within the statistically significant range for each task, excluding the oracle. Standard deviation across multiple runs in the parenthesis.
+‡AA: Advisory autonomy tasks, Ring: Single lane ring, Ramp: Highway ramp, Acc: Acceleration guidance, Vel: Speed guidance.
+A.4.3 Detailed sample complexity comparison results
+Table 5 presents a comparison of sample complexity required for MBTL to perform as good as the best generalization performance of baselines (Independent training and multi-task training) across various tasks in the CMDP.
+Each row lists a different domain, the specific context variation applied (e.g., changes in physical properties or environmental parameters), and two key values: $k^{*}$ and $N$, where $k^{*}$ represents the number of models required by MBTL to reach a performance level comparable to the baseline.
+This value is shown as a range (e.g., $[3,5,3]$), indicating results from three random seeds.
+$N$ represents the total number of contexts.
+The value $\frac{N}{k^{*}}$ helps illustrate the sample efficiency of MBTL.
+Table 5: Sample complexity comparison to baseline performance on CMDP tasks
+Task Context Variation $k^{*}$ $N$
+Pendulum Length [3, 5, 3] 100
+Pendulum Mass [3, 3, 3] 100
+Pendulum Timestep [6, 5, 6] 100
+Cartpole Mass of Cart [9, 13, 6] 100
+Cartpole Length of Pole [4, 4, 3] 100
+Cartpole Mass of Pole [3, 3, 3] 100
+BipedalWalker Gravity [2, 2, 3] 100
+BipedalWalker Friction [2, 2, 2] 100
+BipedalWalker Scale [3, 3, 1] 100
+HalfCheetah Gravity [2, 2, 2] 100
+HalfCheetah Friction [1, 2, 2] 100
+HalfCheetah Stiffness [1, 2, 2] 100
+AA-Ring-Acc Hold Duration [3, 8, 2] 40
+AA-Ring-Vel Hold Duration [2, 1, 1] 40
+AA-Ramp-Acc Hold Duration [1, 2, 2] 40
+AA-Ramp-Vel Hold Duration [2, 2, 2] 40
+Traffic Signal Road Length [8, 5, 6] 50
+Traffic Signal Inflow [2, 2, 2] 50
+Traffic Signal Speed Limit [4, 4, 5] 50
+Eco-Driving Penetration Rate [2, 2, 1] 50
+Eco-Driving Inflow [2, 1, 1] 50
+Eco-Driving Green Phase [2, 2, 2] 50
+A.4.4 Details about traffic signal control
 Most traffic lights operate on fixed schedules, but adaptive traffic signal control using DRL can optimize the traffic flow using real-time information on the traffic [9, 23], though challenges persist in generalizing across various intersection configurations [19].
@@
 The plots display how different strategies adapt with increasing transfer steps:
-  • Inflow: Performance improves as the number of transfer steps increases, with MBTL-GP strategy consistently achieving the highest scores, demonstrating their effectiveness in adapting to changes in inflow conditions.
+  • Inflow: Performance improves as the number of transfer steps increases, with MBTL strategy consistently achieving the highest scores, demonstrating their effectiveness in adapting to changes in inflow conditions.
   • Speed Limit: Here, performance levels are relatively stable across all strategies except for the multitask training.
-  • Road Length: There is a general upward trend in performance for all strategies, particularly for MBTL-GP, indicating robustness in adapting to different road lengths.
-This data suggests that MBTL-GP and Oracle are particularly effective across varying conditions, maintaining higher levels of performance adaptability.
+  • Road Length: There is a general upward trend in performance for all strategies, particularly for MBTL, indicating robustness in adapting to different road lengths.
+This data suggests that MBTL and Oracle are particularly effective across varying conditions, maintaining higher levels of performance adaptability.
 Figure 13: Comparison of normalized generalized performance of all target tasks: Traffic signal control.
-A.2.4 Details about eco-driving Lagrangian control
+A.4.5 Details about eco-driving control
 Given the significant portion of greenhouse gas emissions in the United States coming from the transportation sector [12], eco-driving behaviors are critical for climate change mitigation.
-Deep reinforcement learning-based eco-driving strategies have been developed [13, 38, 17] but still have some issues of difficulties in generalization.
+Deep reinforcement learning-based eco-driving strategies have been developed [14, 45, 18, 20] but still have some issues of difficulties in generalization.
 We also extend to various intersection configurations with different traffic inflow rates, penetration rates of eco-driving systems, and durations of green phases at static traffic signals to optimize vehicle behaviors for reduced emissions.
-Figure 13 illustrates the traffic road network used in eco-driving Lagrangian control task.
-The road network is depicted with traffic flowing vertically and horizontally crossing the static phase traffic signal.
+Figure 14 illustrates the traffic road network used in the eco-driving control task.
+The road network is depicted as traffic flowing vertically and horizontally, crossing the static phase traffic signal.
 There are both guided and default vehicles in the system.
@@
 (c) Penetration rate variation
-Figure 14: Examples of transferability heatmap for eco-driving Lagrangian control.
+Figure 15: Examples of transferability heatmap for eco-driving control.
 Results
@@
 The graphs depict performance enhancement over transfer steps for different strategies.
-Notably, MBTL-GP consistently demonstrates superior performance across all variations, indicating robust adaptability to changing task parameters.
-Figure 15: Comparison of normalized generalized performance of all target tasks: Eco-driving Lagrangian control.
-A.2.5 Details about advisory autonomy task
-Advisory autonomy involves a real-time speed advisory system that enables human drivers to emulate the system-level performance of autonomous vehicles in mixed autonomy systems [34, 15].
+Notably, MBTL consistently demonstrates superior performance across all variations, indicating robust adaptability to changing task parameters.
+Figure 16: Comparison of normalized generalized performance of all target tasks: Eco-driving control.
+A.4.6 Details about advisory autonomy task
+Advisory autonomy involves a real-time speed advisory system that enables human drivers to emulate the system-level performance of autonomous vehicles in mixed autonomy systems [39, 8, 16].
 Instead of direct and instantaneous control, human drivers receive periodic guidance, which varies based on road type and guidance strategy.
@@
 Figure 19 illustrates the comparison of normalized generalized performance for advisory autonomy tasks, specifically acceleration and speed guidance in a ring and acceleration guidance on a ramp.
-The graphs demonstrate that MBTL-GP consistently exhibits higher performance across all tasks.
-Particularly, acceleration guidance in both ring and ramp scenarios shows significant performance improvements over transfer steps, with MBTL-GP closely matching in some instances.
+The graphs demonstrate that MBTL consistently exhibits higher performance across all tasks.
+Particularly, acceleration guidance in both ring and ramp scenarios shows significant performance improvements over transfer steps, with MBTL closely matching in some instances.
 Figure 19: Comparison of normalized generalized performance of all target tasks: Advisory autonomy.
-A.2.6 Details of control tasks
+A.4.7 Details of control tasks
 For this experimental phase, we selected context-extended versions of standard RL environments from the CARL benchmark library, including Cartpole, Pendulum, BipedalWalker, and Halfcheetah.
@@
 The range of context variations was established by scaling the default values specified in the CARL framework from 0.1 to 10 times, enabling a comprehensive examination of each model’s performance under drastically different conditions.
-A.2.7 Results of control tasks
-Figure 19 compares our proposed methods with baselines and oracle method in different variations of context in classic control tasks.
-Figure 19: Comparison on control benchmarks: Cartpole, Pendulum, BipedalWalker, and Halfcheetah.
-A.2.8 Details about Cartpole
+A.4.8 Details about Cartpole
 Transferability heatmap
@@
 Figure 21 presents a comparison of normalized generalized performance for the Cartpole task across different strategies when varying the mass of the cart, length of the pole, and mass of the pole.
-In the mass of cart variation, the performance generally increases with transfer steps, with MBTL-GP strategies achieving the highest scores, indicating robust adaptability to changes in cart mass.
+In the mass of cart variation, the performance generally increases with transfer steps, with MBTL strategies achieving the highest scores, indicating robust adaptability to changes in cart mass.
 Similar trends are observed with length variation and mass of pole variation.
-MBTL-GP shows close to oracle performance.
+MBTL shows close to oracle performance.
 Figure 21: Comparison of normalized generalized performance of all target tasks: Cartpole.
-A.2.9 Details about Pendulum
+A.4.9 Details about Pendulum
 Transferability heatmap
@@
 Figure 23 shows a comparison of normalized generalized performance for the Pendulum task across different strategies when varying the timestep, length of the pendulum, and mass of the pendulum.
-For the length of the pendulum variation and mass of the pendulum one, MBTL-GP strategies demonstrate the highest scores, suggesting robust adaptability to changes in pendulum dynamics.
-MBTL-GP shows performance close to that of the Oracle across all variations, indicating its effectiveness in handling dynamic changes in system parameters.
+For the length of the pendulum variation and mass of the pendulum one, MBTL strategies demonstrate the highest scores, suggesting robust adaptability to changes in pendulum dynamics.
+MBTL shows performance close to that of the Oracle across all variations, indicating its effectiveness in handling dynamic changes in system parameters.
 Figure 23: Comparison of normalized generalized performance of all target tasks: Pendulum.
-A.2.10 Details about BipedalWalker
+A.4.10 Details about BipedalWalker
 Transferability heatmap
@@
 Figure 25 shows the comparison of normalized generalized performance for all variations within the BipedalWalker task.
-There is no huge difference in performance for all three cases, but if we look into the tabualr results in Table 2, MBTL-GP shows the highest performance across varying conditions, indicating their robustness in adapting to changes in physical parameters of the model.
+There is no huge difference in performance for all three cases, but if we look into the tabualr results in Table 2, MBTL shows the highest performance across varying conditions, indicating their robustness in adapting to changes in physical parameters of the model.
 This suggests that these strategies are more effective in handling the complexities introduced by different frictions, gravities, and scales compared to other baselines.
 Figure 25: Comparison of normalized generalized performance of all target tasks: BipedalWalker.
-A.2.11 Details about HalfCheetah
+A.4.11 Details about HalfCheetah
 Transferability heatmap
@@
 Figure 27 presents a comparison of normalized generalized performance across various strategies for the HalfCheetah task with respect to the varied physical properties.
-The results indicate that the MBTL-GP generally outperforms others, particularly in managing variations in gravity and stiffness, suggesting the superior adaptability of these models to physical changes in the task environment.
+The results indicate that the MBTL generally outperforms others, particularly in managing variations in gravity and stiffness, suggesting the superior adaptability of these models to physical changes in the task environment.
 The trends across different parameters confirm the critical impact of task-specific dynamics on the effectiveness of the tested strategies.
 Figure 27: Comparison of normalized generalized performance of all target tasks: HalfCheetah.
-A.3 Potential impacts
+A.5 Potential impacts
 Our work has the potential to reduce the computational effort needed to solve complex real-world problems, offering scalable solutions for implementing deep reinforcement learning in dynamic environments.

### v3

@@
 We experimentally validate our methods using urban traffic and standard continuous control benchmarks.
-The experimental results suggest that MBTL can achieve up to 50x improved sample efficiency compared with canonical independent training and multi-task training.
+The experimental results suggest that MBTL can achieve up to 43x improved sample efficiency compared with canonical independent training and multi-task training.
 Further experiments demonstrate the efficacy of BO and the insensitivity to the underlying RL algorithm and hyperparameters.
@@
 In this work, we consider training an intermediate set of $K$ models, where $N>K>1$, in an effort to balance performance and efficiency; we refer to this strategy as multi-policy training.
+Figure 1: Normalized performance comparison across different problem variations in Eco-Driving benchmark.
+Traditional DRL approaches (e.g., Independent training or multi-task training) exhibit greater training instability, whereas Oracle Transfer, zero-shot transfer with full information, shows the potential for performance improvement by multi-policy training.
+We build upon zero-shot transfer, a widely-used practical technique which directly applies a policy trained in one context (source task) to another (target task) without adaptation.
+Figure 1 shows that multi-policy training with zero-shot transfer has the potential to improve the performance even over the independent training.
+In this article, we strategically select source tasks by explicitly modeling the generalization performance to estimate the value of training a new source task.
 A note on terminology.
@@
 This article is concerned with training reliability.
-Figure 1: Example generalization gap depicted for Cartpole CMDP.
-The solid lines show the true zero-shot transfer generalization performance across contexts.
-Source tasks are indicated by dotted lines.
-We build upon zero-shot transfer, a widely-used practical technique which directly applies a policy trained in one context (source task) to another (target task) without adaptation.
-This experiences performance degradation, also called generalization gap [17, 22].
-For instance, Figure 1 depicts that the performance degrades as the target task diverges from the source task, corresponding to an increasing generalization gap.
-Nonetheless, leveraging the notion that training is expensive but zero-shot transfer is cheap, we are interested in optimally selecting a set of source (training) tasks, such that the generalization performance on the target range of tasks is maximized.
-In this article, we strategically select source tasks by explicitly modeling the generalization performance to estimate the value of training a new source task.
 The main contributions of this work are:
-  • We introduce Model-Based Transfer Learning (MBTL), a novel framework for studying generalization in reinforcement learning (Figure 2).
-To the best of our knowledge, this is the first work to explicitly model generalization performance for contextual RL.
-As such, our work opens the door for further investigation into reliable model-based methodologies for contextual RL.
+  • We introduce Model-Based Transfer Learning (MBTL), a novel framework for solving CMDP sample efficiently (Figure 2).
+To the best of our knowledge, this is the first work to explicitly model generalization performance for contextual RL (CRL).
+As such, our work opens the door for further investigation into reliable model-based methodologies for CRL.
   • We provide theoretical analysis for the sublinear regret of MBTL and give conditions for achieving tighter regret bounds.
-  • We empirically validate our methods in urban traffic and standard continuous control benchmarks for contextual RL, observing up to 50x improvements in sample efficiency.
+  • We empirically validate our methods in urban traffic and standard continuous control benchmarks for contextual RL, observing up to 43x improvements in sample efficiency.
 We further include ablations on the components of the algorithm.
@@
 We distinguish between estimated values $\hat{J}(x)$ and observed outcomes $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)$, with the latter measured after training and evaluation.
+Figure 3: Example generalization gap depicted for Cartpole CMDP.
+The solid lines show the true zero-shot transfer generalization performance across contexts.
+Source tasks are indicated by dotted lines.
 Generalization gap via zero-shot transfer.
@@
 Zero-shot transfer involves applying a policy trained on a source task ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x}}$ to a different target task ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x^{\prime}}}$, with $x,x^{\prime}\in X$.
+This experiences performance degradation, also called generalization gap [17, 23].
+For instance, Figure 3 depicts that the performance degrades as the target task diverges from the source task, corresponding to an increasing generalization gap.
+Nonetheless, leveraging the notion that training is expensive but zero-shot transfer is cheap, we are interested in optimally selecting a set of source (training) tasks, such that the generalization performance on the target range of tasks is maximized.
 We observe the generalization performance, denoted by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{x},x^{\prime}})$, by evaluating the target task $x^{\prime}$ based on the policy trained using source task $x$ via zero-shot generalization.
 We define the generalization gap as the absolute performance difference in average reward when transferring from source task $x$ to target task $x^{\prime}$:
-$\underbrace{\Delta J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{Generalization gap}}=\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)}_{\text{Source task performance}}-\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Generalization} performance}}.$ (1)
+$\underbrace{\Delta J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{Generalization gap}}=|\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)}_{\text{Source task performance}}-\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Generalization} performance}}|.$ (1)
 3 Problem formulation
@@
 In this section, we introduce an algorithm called Model-based Transfer Learning to solve the SSTS problem.
-MBTL models the generalization performance in two parts: 1) the performance set point, modeled using Gaussian processes, and 2) performance loss (generalization gap), modeled as a linear function of contextual similarity.
+MBTL models the generalization performance in two parts: 1) the performance set point, modeled using Gaussian processes, and 2) generalization gap, modeled as a linear function of contextual similarity.
 MBTL combines these two pieces of information within a Bayesian optimization (BO) framework to sequentially select training tasks that maximize generalization performance.
-The high-level pseudo-code of the MBTL algorithm is given below.
-MBTL Pseudo-code 1:Input: Task set, Gaussian Process (GP) model, and training budget 2:while training budget remains do 3:    Select next source task $\leftarrow$ Task with the highest acquisition value based on estimated training performance, generalization gap, and generalization performance, determined by the current GP model and tailored acquisition function 4:    Train a policy on the selected source task using reinforcement learning, then evaluate its training performance 5:    Update the GP model with the selected source task training performance to estimate the training performance of tasks in the task set 6:end while 7:Output: Set of policies that maximizes generalized performance across task set
 4.1 Modeling assumptions
@@
 For example, adjusting the angle of a robotic arm by a small amount typically results in a small change in the system and optimal action.
-Inspired by the empirical generalization gap performance as observed in Figure 1, we estimate the generalization gap $\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x^{\prime})$ with a linear function.
+Inspired by the empirical generalization gap performance as observed in Figure 3, we estimate the generalization gap with a linear function of contextual similarity.
 Assumption 1 (Linear generalization gap).
@@
 We use regret to quantify the effectiveness of our source task selection based on BO.
-Specifically, we define regret at iteration $k$ as $r_{k}=\hat{V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})-\hat{V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$, where $\hat{V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ represents the maximum generalization performance achievable across all tasks, and $\hat{V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ is the performance at the current task selection $x_{k}$.
+Specifically, we define regret at iteration $k$ as $r_{k}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$, where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ represents the maximum generalization performance achievable across all tasks, and $\hat{V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ is the performance at the current task selection $x_{k}$.
 Consequently, the cumulative regret after $K$ iterations is given by $R_{K}=\sum_{k=1}^{K}r_{k}$, summing the individual regrets over all iterations.
@@
 We utilize Deep Q-Networks (DQN) for discrete action spaces [29] and Proximal Policy Optimization (PPO) for continuous action spaces [36].
+For statistical reliability, we run each experiment three times with different random seeds.
 We evaluate our method by the average performance across all $N$ target tasks after training up to $K=$15 source tasks or the number of source tasks needed to achieve a certain level of performance.
@@
 5.2 Traffic benchmark experiments
-Table 1: Comparative performance of different methods on traffic CMDPs
-Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
+Table 1: Comparative performance of different methods on traffic CMDPs ($K=15$)
+Benchmark (CMDP) Baselines Multi-policy Baselines MBTL Oracle
 Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
 Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
-Traffic Signal Inflow 0.8646 0.8319 0.8457 0.8700 0.8496 0.8673 0.8768
-Traffic Signal Speed Limit 0.8857 0.6083 0.8821 0.8858 0.8862 0.8854 0.8876
-Eco-driving Penetration Rate 0.5260 0.1945 0.5959 0.5934 0.5827 0.6323 0.6660
-Eco-driving Inflow 0.4061 0.2229 0.4774 0.4705 0.4673 0.5108 0.5528
-Eco-driving Green Phase 0.3850 0.4228 0.4406 0.4557 0.4431 0.4700 0.5027
-AA-Ring-Acc Hold Duration 0.8362 0.9209 0.8924 0.9057 0.8776 0.9242 0.9552
-AA-Ring-Vel Hold Duration 0.9589 0.972 0.9785 0.9772 0.9807 0.9816 0.9822
-AA-Ramp-Acc Hold Duration 0.4276 0.5158 0.6050 0.5956 0.6143 0.6318 0.7111
-AA-Ramp-Vel Hold Duration 0.5473 0.5034 0.669 0.6787 0.5907 0.7182 0.7686
-Average 0.6778 0.6017 0.7312 0.7354 0.722 0.7559 0.7844
+Traffic Signal Road Length 0.9409 0.8242 0.9366 0.9337 0.9349 0.9364 0.9409
+Traffic Signal Inflow 0.8646 0.8319 0.8699 0.8712 0.8682 0.8432 0.8768
+Traffic Signal Speed Limit 0.8857 0.6083 0.8872 0.8872 0.8874 0.8867 0.8876
+Eco-Driving Penetration Rate 0.526 0.1945 0.6212 0.6278 0.5992 0.6148 0.666
+Eco-Driving Inflow 0.4061 0.2229 0.5077 0.5513 0.5299 0.5172 0.5528
+Eco-Driving Green Phase 0.385 0.4228 0.4724 0.4697 0.4678 0.4985 0.5027
+AA-Ring-Acc Hold Duration 0.8362 0.9209 0.9306 0.9225 0.9136 0.9382 0.9552
+AA-Ring-Vel Hold Duration 0.9589 0.972 0.9819 0.9819 0.9819 0.9818 0.9822
+AA-Ramp-Acc Hold Duration 0.4276 0.5158 0.6750 0.6821 0.6882 0.6964 0.7111
+AA-Ramp-Vel Hold Duration 0.5473 0.5034 0.7170 0.7386 0.6918 0.7273 0.7686
+Average 0.6778 0.6017 0.7600 0.7666 0.7563 0.7641 0.7844
 †Higher the better.
@@
 ‡AA: Advisory autonomy benchmark, Ring: Single lane ring, Ramp: Highway ramp, Acc: Acceleration guidance, Vel: Speed guidance.
-We consider three traffic control domains.
+We consider three traffic control benchmarks.
 First, while most traffic lights still operate on fixed schedules, RL can be used to design adaptive (1) Traffic signal control to optimize traffic [8, 24].
@@
 MBTL efficiently selects source training tasks.
-The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 25x fewer samples needed.
+The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 30x fewer samples.
 Results.
 Table 1 and Figure 5 summarize the results.
-Notably, the Oracle far outperforms the standard baselines (Independent and Multi-task), indicating the potential for transfer learning and intelligent training of multiple models, respectively.
-MBTL rapidly approaches the Oracle, within $\approx 10$ transfer steps, indicating that the Gaussian process effectively models the training performance and linear generalization gap models generalization performance.
+Notably, the Oracle far outperforms the standard baselines (independent and multi-task training), indicating the potential for transfer learning and intelligent training of multiple models, respectively.
+MBTL rapidly approaches the Oracle within $\approx 10$ transfer steps, indicating that the GP effectively models the training performance and linear generalization gap models the generalization performance.
 It is important to note that multi-task RL studies commonly consider Independent training as a strong baseline due to its avoidance of negative transfer and other training instability issues.
-Indeed, Independent often (but not always) outperforms Multi-task in our experiments.
-Yet, our experiments show that MBTL consistently outperforms both Independent and Multi-task and matches their performance with 5-25x better sample efficiency.
-Among the sequential baselines, MBTL consistently outperforms the heuristic baselines, indicating the value of adaptively selecting source tasks based on feedback.
-The sequential baselines, such as random, equidistant, and greedy strategy, also generally outperform Independent and Multi-task, indicating the general value of multi-policy training for solving CMDPs.
+Indeed, independent training often (but not always) outperforms multi-task training in our experiments.
+Yet, our experiments show that MBTL outperforms both independent and multi-task baselines and matches their performance with up to 30x improved sample efficiency.
+Among the multi-policy baselines, MBTL often outperforms the heuristic multi-policy baselines, indicating the value of adaptively selecting source tasks based on feedback.
+The multi-policy baselines, such as random, equidistant, and greedy strategy, also generally outperform Independent and Multi-task, indicating the general value of multi-policy training for solving CMDPs.
 More results are provided in Appendix A.4.
 5.3 Continuous control benchmark experiments
-Table 2: Comparative performance of different methods on standard control CMDPs
-Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
+Table 2: Comparative performance of different methods on standard control CMDPs ($K=15$)
+Benchmark (CMDP) Baselines Multi-policy Baselines MBTL Oracle
 Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
 Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
-Pendulum Length 0.7383 0.6830 0.7270 0.7092 0.7327 0.7697 0.7969
-Pendulum Mass 0.6237 0.5793 0.6329 0.6092 0.6408 0.6827 0.7132
-Pendulum Timestep 0.8135 0.7247 0.7989 0.7541 0.8177 0.8141 0.8801
-Cartpole Mass of Cart 0.9466 0.7153 0.7221 0.7516 0.6501 0.8212 0.9838
-Cartpole Length of Pole 0.9110 0.5441 0.8121 0.8428 0.8217 0.9124 0.9875
-Cartpole Mass of Pole 0.9560 0.6073 0.8858 0.7909 0.8744 0.9351 1.0000
-BipedalWalker Gravity 0.9281 0.7898 0.9330 0.9494 0.9359 0.9393 0.9674
-BipedalWalker Friction 0.9317 0.9051 0.965 0.9664 0.9645 0.9713 0.9778
-BipedalWalker Scale 0.8694 0.7452 0.8605 0.8496 0.8792 0.8886 0.9107
-HalfCheetah Gravity 0.6679 0.6292 0.8542 0.8634 0.8663 0.9073 0.9544
-HalfCheetah Friction 0.6693 0.7242 0.8567 0.8703 0.8591 0.9274 0.9663
-HalfCheetah Stiffness 0.6561 0.7007 0.8533 0.7817 0.8785 0.9146 0.9674
-Average 0.8093 0.6957 0.8251 0.8116 0.8267 0.8736 0.9255
+Pendulum Length 0.7383 0.6830 0.7607 0.7601 0.7774 0.7755 0.7969
+Pendulum Mass 0.6237 0.5793 0.6647 0.6794 0.6887 0.6667 0.7132
+Pendulum Timestep 0.8135 0.7247 0.8331 0.8292 0.8497 0.8333 0.8801
+Cartpole Mass of Cart 0.9466 0.7153 0.8961 0.9044 0.8299 0.9356 0.9838
+Cartpole Length of Pole 0.9110 0.5441 0.9497 0.9484 0.9424 0.9488 0.9875
+Cartpole Mass of Pole 0.9560 0.6073 0.9870 0.9927 0.9916 0.9813 1.0000
+BipedalWalker Gravity 0.9281 0.7898 0.9654 0.9666 0.9656 0.9655 0.9674
+BipedalWalker Friction 0.9317 0.9051 0.9739 0.9738 0.9738 0.9725 0.9778
+BipedalWalker Scale 0.8694 0.7452 0.8910 0.8975 0.8990 0.8962 0.9107
+HalfCheetah Gravity 0.6679 0.6292 0.9086 0.9000 0.9089 0.9214 0.9544
+HalfCheetah Friction 0.6693 0.7242 0.9314 0.9457 0.9184 0.9225 0.9663
+HalfCheetah Stiffness 0.6561 0.7007 0.9191 0.9097 0.9295 0.9205 0.9674
+Average 0.8093 0.6957 0.8901 0.8923 0.8896 0.8950 0.9255
 †Higher the better.
@@
 Method comparison of normalized performance over $N$ tasks.
-The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 50x fewer samples needed.
+The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 43x improved sample efficiency.
 To probe the generality of MBTL, we utilize context-extended versions of standard RL environments from CARL benchmark library [5] to evaluate our methods under varied contexts.
@@
 The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times ($N=100$), enabling a comprehensive analysis of transfer learning under drastically different conditions.
-We provide more details in Appendix A.4.
+We provide more experimental details in Appendix A.4.
 Figure 7: The GP sequentially updates estimates of the performance function (blue) based on previously trained models.
@@
 Results.
-The results summarized in Table 2 demonstrate MBTL’s sample efficiency and competitive performance across diverse control domains, often closely trailing the Oracle only with a small number of tasks.
-With the exception of the Cartpole domain, MBTL generally shows superior performance.
-Figure 5 shows that MBTL demonstrates rapid learning curves, reaching high normalized performance with relatively few samples.
+The results summarized in Table 2 demonstrate sample efficiency and competitive performance of multi-policy training including MBTL across diverse control domains, often closely trailing the Oracle only with a small number of trained policies.
+Figure 6 shows that with the exception of a few context variations, MBTL generally shows superior performance.
 Specifically, Figure 7 illustrates the detailed process of how MBTL utilizes GP for performance estimation and chooses the next source task that maximizes the acquisition function that evaluates the expected improvement of generalized performance.
-MBTL achieves comparable performance to multi-task or independent baselines with 8-50x fewer samples, highlighting its effectiveness in reducing training requirements.
+MBTL achieves comparable performance to multi-task or independent baselines with 5-43x fewer samples, highlighting its effectiveness in reducing training requirements.
 Figure 8: Sensitivity analysis on the DRL algorithm underlying MBTL (DQN, PPO, and A2C), tested on Cartpole with varying length of pole. MBTL remains effective.
@@
 Robustness and generalization challenges in DRL are generally addressed by a few common techniques in the literature.
-The broader umbrella of such methods falls under CRL [6].
-CRL utilizes side information about the problem variations to improve the generalization and robustness.
+The broader umbrella of such methods falls under CRL, which utilizes side information about the problem variations to improve the generalization and robustness.
 In particular, CRL formalizes generalization in DRL using CMDPs [14, 32, 5], which incorporate context-dependent dynamics, rewards, and initial state distributions into the formalism of MDPs.
-This provides a structured framework to study generalization in DRL [15, 32].
 The contexts of CMDPs are not always visible during training [23].
 When they are visible, they can be directly used as side information by conditioning the policy on them [39].
-Multi-task learning.
-A common approach for solving CRL is multi-task learning, which aims to leverage common structure between tasks.
-The structure between tasks has been exploited using policy sketches for task structuring [2] and by sharing a distilled policy that captures common behavior across tasks [44].
-One of the known challenges in multi-task learning is that when the context features are not visible, the resulting CMDP becomes a partially observable CMDP [22, 10] and can be challenging for multi-task learning.
-Moreover, another challenge is negative transfer, where training from tasks that are too different results in training instability or failure [21, 41, 42].
-In this work, we use multi-task learning as a baseline in evaluating our methods.
+In this paper, we focus on a scenario where the learner can choose which context-MDP to train on.
+This contrasts with other CRL works that assume context-MDPs arrive from a fixed distribution.
+Multi-task training.
+Multi-task methods can help address CRL by exploiting shared structure across tasks.
+Prior work has leveraged techniques such as policy sketches for task decomposition [2] and distilled policies that capture common behaviors [46].
+However, a key limitation arises when the context is unobserved, effectively transforming the CMDP into a partially observable setting [23, 9], which complicates multi-task training.
+Another challenge is negative transfer, wherein training on tasks that are too dissimilar leads to instability or outright failure [22, 42, 44].
+Although more recent multi-task approaches such as MOORE [16] and PaCo [43] have shown promise, they often focus on discrete task sets and are thus less suited to CRL, where tasks span a broad continuum of contexts.
+In this work, we include multi-task learning as a baseline to benchmark our methods.
 Zero-shot transfer and policy reuse.
-Zero-shot transfer is another commonly used technique that focuses on adapting models trained in one environment to perform effectively in unseen settings without further training [22].
-Li et al.
-[25] introduces context-aware policy reuse, which adapts multiple source policies based on context, enhancing generalization in complex environments.
-Rezaei-Shoshtari et al.
-[34] utilize hypernetworks to approximate an RL algorithm as a mapping from parameterized CMDP to a family of near-optimal solutions to achieve zero-shot transfer across varying task conditions.
+Zero-shot transfer—where models trained for one environment directly perform in new, unseen settings without additional training [23]—is an important strategy in CRL settings.
+For solving CMDP problems, prior works attempted to utilize zero-shot transfer to solve CMDP problems by approximation on RL algorithm and hypernetworks that maps from parameterized CMDP to a family of near-optimal solutions [35].
 Sinapov et al.
@@
 [3] propose a metric for evaluating transferability based on information-theoretic feature representations across tasks.
-Another popular class of techniques for robust and generalizable DRL is domain adaptation strategies.
-It is often used in making learned policies adapt to new domains with varying dynamics, observations, or rewards [5].
-For instance, Xing et al.
-[46] first learn a latent unified state representation for different domains and subsequently train DRL agents in the source domain based on that.
-However, our focus in this work centered around in-distribution generalization, whereas domain adaptation techniques are commonly used for out-of-distribution generalization.
+Taken together, these approaches highlight the importance of policy reuse, where efficiently selecting or adjusting a pre-trained policy accelerates learning and improves robustness in new contexts.
 Source task selection.
@@
 [33] finds that selecting aligned intermediate tasks in natural language processing boosts transfer effectiveness.
-Together, these studies show the importance of context-aligned source tasks selection to improve transferability and reduce negative transfer.
+Building upon these insights, we formulate the source task selection problem for CRL, enabling zero-shot transfer by estimating training performance online and leveraging structural generalization across context variations.
 7 Conclusion
@@
 MBTL has two key components: an explicit model of the generalization gap and a Gaussian process component to estimate training performance.
-MBTL achieves up to 50x improved sample efficiency on standard and real-world benchmarks.
+MBTL achieves up to 43x improved sample efficiency on standard and real-world benchmarks.
 Furthermore, MBTL achieves sublinear regret in the number of training tasks.
-A limitation is that MBTL is designed for a single-dimensional context variation.
-Promising directions of future work include studying high-dimensional context spaces and out-of-distribution generalization, as well as the development of new real-world CMDP benchmarks.
+A limitation is that MBTL is designed for a single-dimensional context variation with a reliance on the explicit similarity of context variables.
+Promising directions of future work include studying high-dimensional context spaces and formalizing task similarity, as well as the development of new real-world CMDP benchmarks.
 Acknowledgments and Disclosure of Funding
@@
 Figure 10 helps understand the discrepancy between the observed generalized performance and the predicted one.
-Figure 10 illustrates how to calculate the marginal improvement of expected generalized performance ($\hat{V}(x;x_{1:k-1})-V(x;x_{1:k-1})$).
+Figure 11 illustrates how to calculate the marginal improvement of expected generalized performance ($\hat{V}(x;\pi_{1:k-1})-V(x_{1:k-1})$).
 […]
@@
 A.4 Experiment details
-A.4.1 Details about Gaussian process (GP)
-In our study, we conducted hyperparameter tuning experiments for Gaussian Process (GP) regression to optimize its performance.
-Specifically, we varied the noise standard deviation over the set $\{0.001,0.01,0.1,1\}$ and the number of restarts for the optimizer over the set $\{5,6,7,8,9,10,11,12,13,14,15\}$ with several kernel configurations.
-We set noise standard deviation as $0.1$, the number of restarts for the optimizer over the set as $9$, and the kernel as the combination of a constant kernel (C) with the bound from 0.001 to 1000 and a radial basis function kernel (RBF) with a length scale ranging from 0.01 to 100.
-The noise level was set to the square of the noise standard deviation ($0.1^{2}$).
-A.4.2 Results of table with standard deviation
-Table 4: Comparative performance of different methods on context-variant traffic and benchmark CMDPs
-Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
+A.4.1 Details about Gaussian process (GP) Regresstion
+We use the GaussianProcessRegressor implementation from scikit-learn, which follows Algorithm 2.1 of [48]. Specifically, we construct a kernel by multiplying a constant kernel
+$C(\theta)=1.0,\quad\theta\in(10^{-3},10^{3}),$
+by a radial basis function (RBF) kernel
+$k_{\mathrm{RBF}}(\mathbf{x},\mathbf{x}^{\prime};\ell)=\exp\Bigl{(}-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}{2\ell^{2}}\Bigr{)}$
+with an initial length scale $\ell=1.0$ (constrained to lie in the range $[10^{-2},10^{2}]$).
+To determine the hyperparameters, we begin by generating synthetic data that aligns with our modeling assumptions, including constant training performance and a linear generalization gap, while introducing noise to degrade generalization performance by up to 10%, sampled from a uniform distribution.
+We vary the GP hyperparameters, including noise standard deviation over the set $\{0.001,0.01,0.1,1\}$, the number of restarts for the optimizer over $\{5,6,\dots,15\}$, and explore several kernel configurations on the synthetic data.
+We then select the hyperparameter configuration that maximizes the average predictive performance.
+Specifically, we choose a noise standard deviation of $\sigma=0.001$ and perform $15$ random restarts of the hyperparameter optimizer to reduce the risk of convergence to poor local minima.
+We use the same GP hyperparameter configuration across all experiments and benchmarks.
+A.4.2 Accuracy of generalization gap assumption
+In Figure 12, we report the Pearson correlation between the observed generalization gap and the estimated gap under our linear assumption (Assumption 1).
+Each histogram shows how strongly the two measures align across various tasks in standard control (blue) and traffic (red) benchmarks.
+Many tasks cluster around moderate positive correlations (0.3–0.5), suggesting that a linear function of context similarity can reasonably capture the gap in most scenarios.
+However, certain tasks—such as Eco-driving—exhibit higher correlations (above 0.6), whereas others—such as HalfCheetah—are closer to 0, indicating that the assumption holds more effectively in some domains than in others.
+Figure 12: Accuracy of linear generalization gap assumption. Pearson correlation analysis on the observed generalization gap and the estimated gap under our linear assumption.
+A.4.3 Results of table with standard deviation
+Table 4: Comparative performance of different methods on context-variant traffic and benchmark CMDPs ($K=15$)
+Benchmark (CMDP) Baselines Multi-policy Baselines MBTL Oracle
 Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
@@
 0.8242
-0.9249
-0.9213
-0.9278
-0.9371
+0.9366
+0.9337
+0.9349
+0.9364
 0.9409
@@
 0.8319
-0.8457
-0.8700
-0.8496
-0.8673
+0.8699
+0.8712
+0.8682
+0.8432
 0.8768
@@
 0.6083
-0.8821
-0.8858
-0.8862
-0.8854
+0.8872
+0.8872
+0.8874
+0.8867
 0.8876
@@
 0.1945
-0.5959
-0.5934
-0.5827
-0.6323
+0.6212
+0.6278
+0.5992
+0.6148
 0.6660
@@
 0.2229
-0.4774
-0.4705
-0.4673
-0.5108
+0.5077
+0.5513
+0.5299
+0.5172
 0.5528
@@
 0.4228
-0.4406
-0.4557
-0.4431
-0.4700
+0.4724
+0.4697
+0.4678
+0.4985
 0.5027
@@
 0.9209
-0.8924
-0.9057
-0.8776
-0.9242
+0.9306
+0.9225
+0.9136
+0.9382
 0.9552
@@
 0.9720
-0.9785
-0.9772
-0.9807
-0.9816
+0.9819
+0.9819
+0.9819
+0.9818
 0.9822
@@
 0.5158
-0.6050
-0.5956
-0.6143
-0.6318
+0.6750
+0.6821
+0.6882
+0.6964
 0.7111
@@
 0.5034
-0.6690
-0.6787
-0.5907
-0.7182
+0.7170
+0.7386
+0.6918
+0.7273
 0.7686
@@
 0.6830
-0.7270
-0.7092
-0.7327
-0.7697
+0.7607
+0.7601
+0.7774
+0.7755
 0.7969
@@
 0.5793
-0.6329
-0.6092
-0.6408
-0.6827
+0.6647
+0.6794
+0.6887
+0.6667
 0.7132
@@
 0.7247
-0.7989
-0.7541
-0.8177
-0.8141
+0.8331
+0.8292
+0.8497
+0.8333
 0.8801
@@
 0.7153
-0.7221
-0.7516
-0.6501
-0.8212
+0.8961
+0.9044
+0.8299
+0.9356
 0.9838
@@
 0.5441
-0.8121
-0.8428
-0.8217
-(0.0980 )
-0.9124
+0.9497
+0.9484
+0.9424
+0.9488
 0.9875
@@
 0.6073
-0.8858
-0.7909
-0.8744
-0.9351
+0.9870
+0.9927
+0.9916
+0.9813
 1.0000
@@
 0.7898
-0.9330
-0.9494
-0.9359
-0.9393
+0.9654
+0.9666
+0.9656
+0.9655
 0.9674
@@
 0.9051
-0.9650
-0.9664
-0.9645
-0.9713
+0.9739
+0.9738
+0.9738
+0.9725
 0.9778
@@
 0.7452
-0.8605
-0.8496
-0.8792
-0.8886
+0.8910
+0.8975
+0.8990
+0.8962
 0.9107
@@
 0.6292
-0.8542
-0.8634
-0.8663
-0.9073
+0.9086
+0.9000
+0.9089
+0.9214
 0.9544
@@
 0.7242
-0.8567
-0.8703
-0.8591
-0.9274
+0.9314
+0.9457
+0.9184
+0.9225
 0.9663
@@
 0.7007
-0.8533
-0.7817
-0.8785
-0.9146
+0.9191
+0.9097
+0.9295
+0.9205
 0.9674
@@
 0.6337
-0.8068
-0.7996
-0.8037
-0.8435
+0.8309
+0.8352
+0.829
+0.8355
 0.8833
@@
 ‡AA: Advisory autonomy tasks, Ring: Single lane ring, Ramp: Highway ramp, Acc: Acceleration guidance, Vel: Speed guidance.
-A.4.3 Detailed sample complexity comparison results
+A.4.4 Detailed sample complexity comparison results
 Table 5 presents a comparison of sample complexity required for MBTL to perform as good as the best generalization performance of baselines (independent training and multi-task training) across various tasks in the CMDP.
@@
 $N$ represents the total number of contexts.
-The value $\frac{N}{k^{*}}$ helps illustrate the sample efficiency of MBTL.
+The value $\frac{N}{k^{*}}$ helps represent the sample efficiency of MBTL.
 Table 5: Sample complexity comparison to baseline performance on CMDP tasks
@@
 Eco-Driving Green Phase [3, 3, 5] 50
-A.4.4 Details about traffic signal control
+A.4.5 Details about traffic signal control benchmark
 Most traffic lights operate on fixed schedules, but adaptive traffic signal control using DRL can optimize the traffic flow using real-time information on the traffic [8, 24], though challenges persist in generalizing across various intersection configurations [19].
@@
 Different configurations of intersections (e.g., road length, inflow, speed limits) are modeled to represent varying real-world conditions.
+To represent various real-world conditions, we vary factors such as road length, inflow rate, and speed limits from 0.1 to 5 times; by default, the road length is 750 meters, the flow rate is 1000 vehicles per hour, and the speed limit is 13.89 m/s.
+Figure 13: Illustration of the traffic networks in traffic signal control task.
 Training configuration
 We used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0 and PPO for RL algorithm [36].
+We utilized the default implementation of the PPO algorithm with default hyperparameters provided by the Stable-Baselines3 library [34].
 All experiments are done on a distributed computing cluster equipped with 48 Intel Xeon Platinum 8260 CPUs.
-Figure 11: Illustration of the traffic networks in traffic signal control task.
+License
+Traffic signal control benchmark falls under MIT License.
+Potential of multi-policy training and zero-shot transfer
+Figure 14: Normalized performance of three DRL-based methods—Oracle Transfer (gray), independent training (orange), and multi-task training (green)—under different traffic-signal benchmarks: flow rate (left), speed limit (middle), and lane length (right).
+While independent and multi-task training approaches exhibit higher variance and reduced asymptotic performance, Oracle Transfer benefits from zero-shot transfer with full information, demonstrating more stable and generally higher performance.
+Figure 14 shows how each approach adapts to variations in flow rate, speed limit, and lane length for a traffic signal control benchmark.
+The y-axis shows normalized performance, with higher values indicating better control policies.
+Oracle Transfer consistently achieves superior performance across these different settings, owing to its ability to leverage full task information in a zero-shot transfer manner.
+By contrast, independent and multi-task training exhibit more pronounced performance drops and greater instability when faced with shifts in problem parameters, underscoring the challenges of generalizing policies in traditional DRL approaches.
 Transferability heatmap
@@
   • Inflow: Performance improves as the number of transfer steps increases, with MBTL strategy consistently achieving the highest scores, demonstrating their effectiveness in adapting to changes in inflow conditions.
-  • Speed Limit: Here, performance levels are relatively stable across all strategies except for the multitask training.
+  • Speed Limit: Here, performance levels are relatively stable across all strategies except for the multi-task training.
   • Road Length: There is a general upward trend in performance for all strategies, particularly for MBTL, indicating robustness in adapting to different road lengths.
@@
 Figure 16: Comparison of normalized generalized performance of all target tasks: Traffic signal control.
-A.4.5 Details about eco-driving control
+A.4.6 Details about eco-driving control benchmark
 Given the significant portion of greenhouse gas emissions in the United States coming from the transportation sector [11], eco-driving behaviors are critical for climate change mitigation.
@@
 The reward mechanism is designed to optimize the driving strategy by balancing the average speed of the vehicles against penalties for emissions, thereby promoting eco-friendly driving behaviors within the traffic system.
+Figure 17: Illustration of the traffic networks in eco-driving control task.
 Training configuration
 We also used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0 and PPO for RL algorithm [36].
-Figure 14: Illustration of the traffic networks in eco-driving control task.
+For detailed experimental details and RL hyperparameter configurations, please refer to [20, 21].
+License
+Eco-driving benchmark falls under MIT License [21].
+Potential of multi-policy training and zero-shot transfer
+Figure 1 shows how each RL training paradigm adapts to variations in green phase time, penetration rate, and inflow rate in the eco-driving control benchmark.
+Oracle Transfer remains the strongest method across all configurations, benefiting from zero-shot transfer.
+independent training shows unstable performance across different task variations, performance, while multi-task training lags behind.
+Overall, the trends highlight the advantage of leveraging zero-shot transfer in traffic CMDPs.
 Transferability heatmap
@@
 Figure 19: Comparison of normalized generalized performance of all target tasks: Eco-driving control.
-A.4.6 Details about advisory autonomy task
+A.4.7 Details about advisory autonomy benchmark
 Advisory autonomy involves a real-time speed advisory system that enables human drivers to emulate the system-level performance of autonomous vehicles in mixed autonomy systems [40, 7, 15].
@@
 The highway ramp scenario introduces a more complex dynamic, where vehicles not only travel along the highway but also merge from ramps, creating potential stop-and-go traffic patterns that challenge the adaptability of autonomous guidance systems.
-Problem Definition: In a single-lane scenario, the state space includes the speeds of the ego and leading vehicles, along with the headway.
+Problem Definition
+In a single-lane scenario, the state space includes the speeds of the ego and leading vehicles, along with the headway.
 For highway ramp scenarios, additional states cover the relative positions and speeds of adjacent vehicles.
@@
 Rewards are based on system throughput or average speed of all vehicles in the system.
-Context Variations: We explore different durations of coarse-grained guidance holds to test various levels of human compatibility, adjusting the model based on observed driver behaviors and system performance.
+Context Variations
+We explore different durations of coarse-grained guidance holds to test various levels of human compatibility, adjusting the model based on observed driver behaviors and system performance.
+License
+Advisory autonomy benchmark falls under MIT License [40].
 Figure 20: Illustration of the traffic networks in advisory autonomy task.
+Potential of multi-policy training and zero-shot transfer
+Figure 21: Normalized performance of Oracle Transfer, independent training, and multi-task training under Advisory Autonomy benchmark with human compatibility task variations.
+Figure 21 shows that ring-road networks tend to yield higher performance and smaller gaps compared to highway ramp scenarios.
+In addition, independent training exhibits greater performance drop and variability due to training instability.
+Oracle Transfer retains clear potential improvements over other baselines.
 Transferability heatmap
@@
 Figure 23: Comparison of normalized generalized performance of all target tasks: Advisory autonomy.
-A.4.7 Details of control tasks
+A.4.8 Details of control benchmarks
 For this experimental phase, we selected context-extended versions of standard RL environments from the CARL benchmark library, including Cartpole, Pendulum, BipedalWalker, and Halfcheetah.
@@
 The range of context variations was established by scaling the default values specified in the CARL framework from 0.1 to 10 times, enabling a comprehensive examination of each model’s performance under drastically different conditions.
-A.4.8 Details about Cartpole
+We utilized the default implementation of the DQN, A2C, and PPO algorithm with default hyperparameters provided by the Stable-Baselines3 library [34].
+A.4.9 Details about Cartpole benchmark
+Potential of multi-policy training and zero-shot transfer
+Cartpole may be considered a simpler benchmark than traffic benchmarks, yet independent and multi-task training methods still face notable difficulty when faced with context variations.
+Figure 24 shows that Oracle Transfer performs at the highest performance across problem variations, while independent training or multi-task training shows a larger variance in performance.
+Figure 24: Normalized performance of Oracle Transfer, independent training, and multi-task training in Cartpole benchmarks.
 Transferability heatmap
@@
 Figure 26 presents a comparison of normalized generalized performance for the Cartpole task across different strategies when varying the mass of the cart, length of the pole, and mass of the pole.
-In the mass of cart variation, the performance generally increases with transfer steps, with MBTL strategies achieving the highest scores, indicating robust adaptability to changes in cart mass.
+In the mass of cart variation, performance generally increases with transfer steps, with MBTL strategies achieving the highest scores.
+This indicates robust adaptability to changes in cart mass.
 Similar trends are observed with length variation and mass of pole variation.
@@
 Figure 26: Comparison of normalized generalized performance of all target tasks: Cartpole.
-A.4.9 Details about Pendulum
+A.4.10 Details about Pendulum benchmark
+Potential of multi-policy training and zero-shot transfer
+Pendulum is also one of the simplest benchmarks in classic control.
+In the Pendulum benchmark, we vary three key parameters: time step (left), pole length (middle), and ball mass (right).
+As shown in Figure 27, a few well-trained policies in specific contexts transfer effectively to new tasks, particularly under Oracle Transfer.
+For certain configurations (e.g., shorter poles, lighter balls), Independent and Oracle Transfer both excel, while multi-task struggles.
+These results suggest a remaining performance gap that multi-policy training and zero-shot transfer could help bridge.
+Figure 27: Normalized performance of Oracle Transfer, independent training, and multi-task training in Pendulum benchmarks.
+Figure 28 provides a visual mapping of which trained policy is being applied for each specific context in the Pendulum benchmark.
+CMDP with time step variation demonstrates that the tasks are covered by a few “good" policies nearby.
+In addition, contexts with shorter poles and lighter balls often gravitate toward a single high-performing policy, whereas more challenging configurations may require a specialized or distinct policy.
+This illustrates how Oracle Transfer can seamlessly select from a suite of learned policies, demonstrating robust zero-shot transfer and stronger adaptability.
+Figure 28: Visulazation on which policy is used to solve specific context-MDP in Pendulum CMDP.
 Transferability heatmap
@@
 Figure 30: Comparison of normalized generalized performance of all target tasks: Pendulum.
-A.4.10 Details about BipedalWalker
+A.4.11 Details about BipedalWalker benchmark
+Potential of multi-policy training and zero-shot transfer
+Figure 31 compares the performance of different RL training methods in the BipedalWalker benchmark.
+independent training typically performs nearly as well but suffers intermittent dips.
+Similarly, multi-task training experiences larger swings, occasionally collapsing to low performance in certain parameter regions.
+However, Oracle Transfer remains near-perfect across every setting.
+These patterns highlight how multi-policy training with zero-shot transfer and per-task training generally fare better than a single universal model when faced with diverse environment dynamics.
+Figure 31: Normalized performance of Oracle Transfer, independent training, and multi-task training in BipedalWalker benchmarks.
 Transferability heatmap
@@
 Figure 33: Comparison of normalized generalized performance of all target tasks: BipedalWalker.
-A.4.11 Details about HalfCheetah
+A.4.12 Details about HalfCheetah benchmark
+Potential of multi-policy training and zero-shot transfer
+In this HalfCheetah benchmark, each subplot examines how policies adapt to changing friction, gravity, and stiffness (Figure 34).
+Oracle Transfer maintains nearly perfect scores for all parameter ranges, indicating robust zero-shot adaptability.
+In contrast, independent training experiences larger fluctuations, while multi-task training remains consistent yet at a lower performance plateau.
+The clear gap between Oracle Transfer and the other methods highlights the advantage of leveraging specialized multi-policy training solutions that effectively transfer across diverse dynamics.
+Figure 34: Normalized performance of Oracle Transfer, independent training, and multi-task training in HalfCheetah benchmarks.
 Transferability heatmap
@@
 Figure 36: Comparison of normalized generalized performance of all target tasks: HalfCheetah.
+A.4.13 Implementation of the recent multi-task baselines
+In Figure 37, we compare two recent multi-task algorithms—PaCo [43] and MOORE [16]—with several our baselines tested on Cartpole CMDP benchmark.
+Although MOORE underperforms relative to our baseline multitask implementation, PaCo achieves competitive or even higher performance at certain points, demonstrating its potential to generalize across multiple tasks.
+Also, it is important to note that those algorithms are naively implemented without thorough investigation.
+These trends show that enhanced multi-task strategies can be beneficial in some CMDP settings, whereas not all multi-task methods readily adapt to broader parameter variations.
+Figure 37: Normalized performance comparison of PaCo [43] and MOORE [16] on Cartpole benchmark.
 A.5 Potential impacts

### v4

@@
 Assumption 1 (Linear generalization gap).
-A linear function is used to model the generalization gap function, formally $\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x^{\prime})\simeq\theta|x-x^{\prime}|$, where $\theta$ is the slope of the linear function and $x$ and $x^{\prime}$ are the context of the source task and target task, respectively.
+[…]
 The relaxation of this assumption can yield additional efficiency benefits but also adds modeling complexity, and thus is an interesting direction for future work.
@@
 Within the framework of BO, we model the source training performance ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\forall x\in X\setminus x_{1:k}}$ using Gaussian process (GP) regression.
-Specifically, the function ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ is assumed to follow a GP $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)\sim\mathcal{GP}(\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)],k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})))$, where $k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ is the covariance function, representing the expected product of deviations of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{\tilde{x}},}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ from their respective means.
+Specifically, the function ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ is assumed to follow a GP $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)\sim\mathcal{GP}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m(x)},k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})))$, where $m(x)$ is mean and $k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ is the covariance function, representing the expected product of deviations of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{\tilde{x}},}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ from their respective means.
 Let $D_{k-1}$ denote the data observed up to iteration $k-1$, consisting of the pairs $\{(x_{i},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{i},x_{i})})\}_{i=1,...,k-1}$.
@@
 It is defined as follows:
-$a(x;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}})=\mathbb{E}_{x^{\prime}\in X}[[\mu_{k-1}(x)+\beta_{k}^{1/2}\sigma_{k-1}(x)-\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})-{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{{1:k-1}},x^{\prime}})]_{+}]$ (4)
+$a(x;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}})=\mathbb{E}_{x^{\prime}\in X}[[\mu_{k-1}(x)+\beta_{k}^{1/2}\sigma_{k-1}(x)-\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{{1:k-1}},x^{\prime}})]_{+}]$ (4)
 where $[\cdot]_{+}$ represents $\max(\cdot,0)$ and we can use various forms of $\beta_{k}$, which is the trade-off parameter between exploitation and exploration.
@@
 Definition 2 (Search space).
-We define the search space $X_{k}$ at iteration $k$ as a subset of $X$, with each element $x^{\prime}\in X_{k}$, such that $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1},x^{\prime}})\leq\hat{J}(\pi_{x_{k}},x^{\prime}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\Delta J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x_{k}},x^{\prime}})$.
+We define the search space $X_{k}$ at iteration $k$ as a subset of $X$, with each element $x^{\prime}\in X_{k}$, such that $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1},x^{\prime}})\leq\hat{J}(\pi_{x_{k}},x^{\prime}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x_{k}},x^{\prime}})$.
 Given the generalization gap observed in Figure 3, we model the degradation from the source task using a linear function in Assumption 1.
@@
 Theorem 2.
-For a given $\delta^{\prime}\in(0,1)$ and scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$, the cumulative regret $R_{K}$ is bounded by $\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}$ with probability at least $1-\delta^{\prime}$.
+For a given $\delta^{\prime}\in(0,1)$ and scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta^{\prime})$, the cumulative regret $R_{K}$ is bounded by $\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}$ with probability at least $1-\delta^{\prime}$.
 Here, $|X|$ denotes the cardinality of the set $X$, the number of elements in $X$.
@@
 The canonical baselines include: (1) Independent training, which involves independently training separate models on each task; and (2) Multi-task RL, where a single “universal” context-conditioned policy is trained for all tasks.
-The multi-policy baselines include: (3) Random selection, where each training task is chosen uniformly at randomly; (4) Equidistant strategy, which selects training tasks by equally subdividing the context space based on a given training budget, and then trains them in lexicographical order; (5) Greedy strategy, which greedily selects the next source task based on Assumption 1 and fixed training performance; and (6) Sequential Oracle transfer, which has access to generalized performance corresponding to policies for all tasks (including those not yet selected) and uses that information to greedily select the best source task at each step.
-Proposed method. We evaluate MBTL with the scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$.
+The multi-policy baselines include: (3) Random selection, where each training task is chosen uniformly at randomly; (4) Greedy strategy, which greedily selects the next source task based on Assumption 1 and fixed training performance; and (5) Sequential Oracle transfer, which has access to generalized performance corresponding to policies for all tasks (including those not yet selected) and uses that information to greedily select the best source task.
+Proposed method.
+In early iterations of BO, GP lacks sufficient observations (often just one or two) and thus relies heavily on its prior.
+To mitigate this, we incorporate a brief warm-up phase in MBTL.
+Specifically, we collect three additional data points using a simple greedy approach that selects tasks with the worst observed performance so far (inspired by [7]).
+After this initialization, the method switches to full BO for source-task selection.
+We use the scaling factor of $\beta_{k}=2\log(|X|k^{2})$.
 DRL algorithms and performance measure.
-We utilize Deep Q-Networks (DQN) for discrete action spaces [29] and Proximal Policy Optimization (PPO) for continuous action spaces [36].
+We utilize Deep Q-Networks (DQN) for discrete action spaces [29] and Trust Region Policy Optimization (TRPO) [36] or Proximal Policy Optimization (PPO) [37] for continuous action spaces.
 For statistical reliability, we run each experiment three times with different random seeds.
@@
 Benchmark (CMDP) Baselines Multi-policy Baselines MBTL Oracle
-Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
-Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
-Traffic Signal Road Length 0.9409 0.8242 0.9366 0.9337 0.9349 0.9364 0.9409
-Traffic Signal Inflow 0.8646 0.8319 0.8699 0.8712 0.8682 0.8432 0.8768
-Traffic Signal Speed Limit 0.8857 0.6083 0.8872 0.8872 0.8874 0.8867 0.8876
-Eco-Driving Penetration Rate 0.526 0.1945 0.6212 0.6278 0.5992 0.6148 0.666
-Eco-Driving Inflow 0.4061 0.2229 0.5077 0.5513 0.5299 0.5172 0.5528
-Eco-Driving Green Phase 0.385 0.4228 0.4724 0.4697 0.4678 0.4985 0.5027
-AA-Ring-Acc Hold Duration 0.8362 0.9209 0.9306 0.9225 0.9136 0.9382 0.9552
-AA-Ring-Vel Hold Duration 0.9589 0.972 0.9819 0.9819 0.9819 0.9818 0.9822
-AA-Ramp-Acc Hold Duration 0.4276 0.5158 0.6750 0.6821 0.6882 0.6964 0.7111
-AA-Ramp-Vel Hold Duration 0.5473 0.5034 0.7170 0.7386 0.6918 0.7273 0.7686
-Average 0.6778 0.6017 0.7600 0.7666 0.7563 0.7641 0.7844
+Domain Context Variation Independent Multi-task Random Greedy Ours Sequential
+Number of Trained Models $N$ $1$ $K$ $K$ $K$ $N$
+Traffic Signal Road Length 0.9409 0.8242 0.9366 0.9349 0.9409 0.9432
+Traffic Signal Inflow 0.8646 0.8319 0.8699 0.8682 0.8729 0.8773
+Traffic Signal Speed Limit 0.8857 0.6083 0.8872 0.8874 0.8866 0.8877
+Eco-Driving Penetration Rate 0.5260 0.1945 0.6212 0.5992 0.6519 0.6668
+Eco-Driving Inflow 0.4061 0.2229 0.5077 0.5299 0.5356 0.5531
+Eco-Driving Green Phase 0.3850 0.4228 0.4724 0.4678 0.4932 0.5058
+AA-Ring-Acc Hold Duration 0.8362 0.9219 0.9307 0.9021 0.9329 0.9567
+AA-Ring-Vel Hold Duration 0.9589 0.9688 0.9820 0.9819 0.9820 0.9822
+AA-Ramp-Acc Hold Duration 0.4276 0.5374 0.6599 0.6570 0.6282 0.7120
+AA-Ramp-Vel Hold Duration 0.5473 0.5257 0.7210 0.6461 0.7426 0.7691
+Average 0.6778 0.6059 0.7589 0.7474 0.7667 0.7854
 †Higher the better.
@@
 ‡AA: Advisory autonomy benchmark, Ring: Single lane ring, Ramp: Highway ramp, Acc: Acceleration guidance, Vel: Speed guidance.
-We consider three traffic control benchmarks.
+We consider three traffic benchmarks.
 First, while most traffic lights still operate on fixed schedules, RL can be used to design adaptive (1) Traffic signal control to optimize traffic [8, 24].
@@
 Among the multi-policy baselines, MBTL often outperforms the heuristic multi-policy baselines, indicating the value of adaptively selecting source tasks based on feedback.
-The multi-policy baselines, such as random, equidistant, and greedy strategy, also generally outperform Independent and Multi-task, indicating the general value of multi-policy training for solving CMDPs.
+The multi-policy baselines, such as random and greedy strategy, also generally outperform Independent and Multi-task, indicating the general value of multi-policy training for solving CMDPs.
 More results are provided in Appendix A.4.
 5.3 Continuous control benchmark experiments
-Table 2: Comparative performance of different methods on standard control CMDPs ($K=15$)
+To probe the generality of MBTL, we utilize context-extended versions of standard RL environments from CARL benchmark library [5] to evaluate our methods under varied contexts.
+For the Cartpole benchmark, we considered CMDPs with varying cart mass, pole length, and pole mass.
+In Pendulum, we vary the timestep duration, pendulum length, and pendulum mass.
+The BipedalWalker was tested under varying friction, gravity, and scale.
+In HalfCheetah domain, we manipulated friction, gravity, and stiffness parameters.
+These variations critically influence the dynamics and physics of the environments.
+The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times ($N=100$), enabling a comprehensive analysis of transfer learning under drastically different conditions.
+We provide more experimental details in Appendix A.4.
+Table 2: Comparative performance of different methods on continuous control CMDPs ($K=15$)
 Benchmark (CMDP) Baselines Multi-policy Baselines MBTL Oracle
-Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
-Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
-Pendulum Length 0.7383 0.6830 0.7607 0.7601 0.7774 0.7755 0.7969
-Pendulum Mass 0.6237 0.5793 0.6647 0.6794 0.6887 0.6667 0.7132
-Pendulum Timestep 0.8135 0.7247 0.8331 0.8292 0.8497 0.8333 0.8801
-Cartpole Mass of Cart 0.9466 0.7153 0.8961 0.9044 0.8299 0.9356 0.9838
-Cartpole Length of Pole 0.9110 0.5441 0.9497 0.9484 0.9424 0.9488 0.9875
-Cartpole Mass of Pole 0.9560 0.6073 0.9870 0.9927 0.9916 0.9813 1.0000
-BipedalWalker Gravity 0.9281 0.7898 0.9654 0.9666 0.9656 0.9655 0.9674
-BipedalWalker Friction 0.9317 0.9051 0.9739 0.9738 0.9738 0.9725 0.9778
-BipedalWalker Scale 0.8694 0.7452 0.8910 0.8975 0.8990 0.8962 0.9107
-HalfCheetah Gravity 0.6679 0.6292 0.9086 0.9000 0.9089 0.9214 0.9544
-HalfCheetah Friction 0.6693 0.7242 0.9314 0.9457 0.9184 0.9225 0.9663
-HalfCheetah Stiffness 0.6561 0.7007 0.9191 0.9097 0.9295 0.9205 0.9674
-Average 0.8093 0.6957 0.8901 0.8923 0.8896 0.8950 0.9255
+Domain Context Variation Independent Multi-task Random Greedy Ours Sequential
+Number of Trained Models $N$ $1$ $K$ $K$ $K$ $N$
+Pendulum Length 0.7383 0.6830 0.7607 0.7774 0.7749 0.8073
+Pendulum Mass 0.6237 0.5793 0.6647 0.6887 0.6933 0.7168
+Pendulum Timestep 0.8135 0.7247 0.8331 0.8497 0.8310 0.8880
+Cartpole Mass of Cart 0.9466 0.7153 0.8961 0.8299 0.9154 0.9998
+Cartpole Length of Pole 0.9110 0.5441 0.9497 0.9424 0.9717 0.9995
+Cartpole Mass of Pole 0.9560 0.6073 0.9870 0.9916 0.9941 1.0000
+BipedalWalker Gravity 0.9281 0.7898 0.9654 0.9656 0.9669 0.9721
+BipedalWalker Friction 0.9317 0.9051 0.9739 0.9738 0.9714 0.9779
+BipedalWalker Scale 0.8694 0.7452 0.8910 0.8990 0.8864 0.9155
+HalfCheetah Gravity 0.6679 0.6292 0.9086 0.9089 0.9308 0.9544
+HalfCheetah Friction 0.6693 0.7242 0.9314 0.9184 0.9404 0.9663
+HalfCheetah Stiffness 0.6561 0.7007 0.9191 0.9295 0.9214 0.9677
+Average 0.8093 0.6957 0.8901 0.8896 0.8998 0.9304
 †Higher the better.
@@
 The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 43x improved sample efficiency.
-To probe the generality of MBTL, we utilize context-extended versions of standard RL environments from CARL benchmark library [5] to evaluate our methods under varied contexts.
-For the Cartpole domain, we considered CMDPs with varying cart mass, pole length, and pole mass.
-In Pendulum, we vary the timestep duration, pendulum length, and pendulum mass.
-The BipedalWalker was tested under varying friction, gravity, and scale.
-In HalfCheetah domain, we manipulated friction, gravity, and stiffness parameters.
-These variations critically influence the dynamics and physics of the environments.
-The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times ($N=100$), enabling a comprehensive analysis of transfer learning under drastically different conditions.
-We provide more experimental details in Appendix A.4.
 Figure 7: The GP sequentially updates estimates of the performance function (blue) based on previously trained models.
@@
 Specifically, Figure 7 illustrates the detailed process of how MBTL utilizes GP for performance estimation and chooses the next source task that maximizes the acquisition function that evaluates the expected improvement of generalized performance.
-MBTL achieves comparable performance to multi-task or independent baselines with 5-43x fewer samples, highlighting its effectiveness in reducing training requirements.
+MBTL achieves comparable performance to multi-task or independent baselines with up to 43x fewer samples, highlighting its effectiveness in reducing training requirements.
 Figure 8: Sensitivity analysis on the DRL algorithm underlying MBTL (DQN, PPO, and A2C), tested on Cartpole with varying length of pole. MBTL remains effective.
@@
 A.3 Theoretical analysis
+This section provides detailed proofs of the regret bounds introduced in Theorem 2, Corollary 2.1, and Corollary 2.2 from the main text.
+Our analysis adapts key results from [42] to settings where the search space is restricted at each iteration.
 A.3.1 Proof of Theorem 2
@@
 Proof.
-The following Lemma 3 and 4 is basically considering the cardinality of restricted search space of $X_{k}$ instead of $X$ upon the lemmas in literature [41].
+We begin by introducing two lemmas (Lemmas 3 and 4) that extend the results in [42] to handle the restricted search space $X_{k}\subseteq X$ at each iteration.
 Lemma 3.
@@
 Lemma 4.
-Setting $\delta\in(0,1)$, $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$, and $C_{1}:=\frac{8}{\log(1+\sigma^{-2})}\geq 8\sigma^{2}$, we have $Pr\left[\sum_{k=1}^{K}r_{k}\left(\frac{|X|}{|X_{k}|}\right)^{2}\leq C_{1}\beta_{K}\gamma_{K}\quad\forall K\geq 1\right]\geq 1-\delta$.
-Using Lemma 3, Lemma 5.3 in [41], Lemma 4, and Cauchy–Schwarz inequality, we can bound the cumulative regret as:
+Setting $\delta^{\prime}\in(0,1)$, $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta^{\prime})$, and $C_{1}:=\frac{8}{\log(1+\sigma^{-2})}\geq 8\sigma^{2}$, we have $Pr\left[\sum_{k=1}^{K}r_{k}\left(\frac{|X|}{|X_{k}|}\right)^{2}\leq C_{1}\beta_{K}\gamma_{K}\quad\forall K\geq 1\right]\geq 1-\delta^{\prime}$.
+Using Lemma 3, we can bound each $r_{t}$ in terms of the restricted search space $X_{k}$.
+Then, applying Lemma 5.3 from [42] (which controls the deviation of the GP-UCB estimator) together with Lemma 4 (which sums these instantaneous regrets under restricted search spaces), and finally invoking the Cauchy–Schwarz inequality, we derive a bound on the cumulative regret.
+Specifically, with probability at least $1-\delta^{\prime}$, we have:
 $R_{K}=\sum_{k=1}^{K}r_{k}\leq\sqrt{\sum_{k=1}^{K}r_{k}\left(\frac{|X|}{|X_{k}|}\right)^{2}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}\leq\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}.$ (5)
@@
 Calculating the sum of squares for the reduced segments, we have:
-$\sum_{k=1}^{K}|X_{k}|^{2}=\sum_{k=1}^{K}2^{-2\lfloor\log_{2}k\rfloor}|X|^{2}\leq\frac{1}{k^{2}}|X|^{2}\leq\frac{\pi^{2}}{6}|X|^{2}$ (8)
+$\sum_{k=1}^{K}|X_{k}|^{2}=\sum_{k=1}^{K}2^{-2\lfloor\log_{2}k\rfloor}|X|^{2}\leq\sum_{k=1}^{K}\frac{1}{k^{2}}|X|^{2}\leq\frac{\pi^{2}}{6}|X|^{2}$ (8)
 The cumulative regret can be bounded as below:
@@
 A.4.3 Results of table with standard deviation
-Table 4: Comparative performance of different methods on context-variant traffic and benchmark CMDPs ($K=15$)
+Table 4: Comparative performance of different methods on context-variant traffic and control CMDPs ($K=15$)
 Benchmark (CMDP) Baselines Multi-policy Baselines MBTL Oracle
-Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
-Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
+Domain Context Variation Independent Multi-task Random Greedy Ours Sequential
 Traffic Signal Road Length
@@
 0.9366
-0.9337
 0.9349
-0.9364
 0.9409
+0.9432
 Traffic Signal Inflow
@@
 0.8699
-0.8712
 0.8682
-0.8432
-0.8768
+0.8729
+0.8773
 Traffic Signal Speed Limit
@@
 0.8872
-0.8872
 0.8874
-0.8867
-0.8876
+0.8866
+0.8877
 Eco-Driving Penetration Rate
@@
 0.6212
-0.6278
 0.5992
-0.6148
-0.6660
+0.6519
+0.6668
 Eco-Driving Inflow
@@
 0.5077
-0.5513
 0.5299
-0.5172
-0.5528
+0.5356
+0.5531
 Eco-Driving Green Phase
@@
 0.4724
-0.4697
 0.4678
-0.4985
-0.5027
+0.4932
+0.5058
 AA-Ring-Acc Hold Duration
 0.8362
-0.9209
-0.9306
-0.9225
-0.9136
-0.9382
-0.9552
+0.9219
+0.9307
+0.9021
+0.9329
+0.9567
 AA-Ring-Vel Hold Duration
 0.9589
-0.9720
+0.9688
+0.9820
 0.9819
-0.9819
-0.9819
-0.9818
+0.9820
 0.9822
@@
 0.4276
-0.5158
-0.6750
-0.6821
-0.6882
-0.6964
-0.7111
+0.5374
+0.6599
+0.6570
+0.6282
+0.7120
 AA-Ramp-Vel Hold Duration
 0.5473
-0.5034
-0.7170
-0.7386
-0.6918
-0.7273
-0.7686
+0.5257
+0.7210
+0.6461
+0.7426
+0.7691
 Pendulum Length
@@
 0.7607
-0.7601
 0.7774
-0.7755
-0.7969
+0.7749
+0.8073
 Pendulum Mass
@@
 0.6647
-0.6794
 0.6887
-0.6667
-0.7132
+0.6933
+0.7168
 Pendulum Timestep
@@
 0.8331
-0.8292
 0.8497
-0.8333
-0.8801
+0.8310
+0.8880
 Cartpole Mass of Cart
@@
 0.8961
-0.9044
 0.8299
-0.9356
-0.9838
+0.9154
+0.9998
 Cartpole Length of Pole
@@
 0.9497
-0.9484
 0.9424
-0.9488
-0.9875
+0.9717
+0.9995
 Cartpole Mass of Pole
@@
 0.9870
-0.9927
 0.9916
-0.9813
+0.9941
 1.0000
@@
 0.9654
-0.9666
 0.9656
-0.9655
-0.9674
+0.9669
+0.9721
 BipedalWalker Friction
@@
 0.9738
-0.9738
-0.9725
-0.9778
+0.9714
+0.9779
 BipedalWalker Scale
@@
 0.8910
-0.8975
 0.8990
-0.8962
-0.9107
+0.8864
+0.9155
 HalfCheetah Gravity
@@
 0.9086
-0.9000
 0.9089
-0.9214
+0.9308
 0.9544
@@
 0.9314
-0.9457
 0.9184
-0.9225
+0.9404
 0.9663
@@
 0.9191
-0.9097
 0.9295
-0.9205
-0.9674
-Average
-0.7775
-0.6337
-0.8309
-0.8352
-0.829
-0.8355
-0.8833
+0.9214
+0.9677
 * Note: Bold values represent the highest value(s) within the statistically significant range for each task, excluding the oracle. Standard deviation across multiple runs in the parenthesis.
@@
 Table 5: Sample complexity comparison to baseline performance on CMDP tasks
-Task Context Variation $k^{*}$ $N$
-Pendulum Length [4, 5, 5] 100
-Pendulum Mass [3, 3, 3] 100
-Pendulum Timestep [8, 5, 14] 100
-Cartpole Mass of Cart [15, 25, 18] 100
-Cartpole Length of Pole [10, 6, 14] 100
-Cartpole Mass of Pole [9, 9, 12] 100
-BipedalWalker Gravity [2, 3, 4] 100
-BipedalWalker Friction [3, 2, 2] 100
-BipedalWalker Scale [7, 6, 1] 100
-HalfCheetah Gravity [10, 2, 5] 100
-HalfCheetah Friction [1, 2, 13] 100
-HalfCheetah Stiffness [1, 2, 5] 100
-AA-Ring-Acc Hold Duration [5, 7, 30] 40
-AA-Ring-Vel Hold Duration [2, 1, 1] 40
-AA-Ramp-Acc Hold Duration [1, 2, 2] 40
-AA-Ramp-Vel Hold Duration [6, 4, 4] 40
-Traffic Signal Road Length [50, 50, 11] 50
-Traffic Signal Inflow [2, 44, 41] 50
-Traffic Signal Speed Limit [10, 4, 8] 50
-Eco-Driving Penetration Rate [3, 5, 1] 50
-Eco-Driving Inflow [2, 1, 1] 50
-Eco-Driving Green Phase [3, 3, 5] 50
+Task Variation $k^{*}$ $N$ $k^{*}$ average $N/k^{*}$ average
+Pendulum Length [3, 5, 3] 100 3.67 27.27
+Pendulum Mass [4, 4, 3] 100 3.67 27.27
+Pendulum Timestep [8, 9, 10] 100 9 11.11
+Cartpole Mass of Cart [18, 14, 22] 100 18 5.56
+Cartpole Length of Pole [13, 12, 10] 100 11.67 8.57
+Cartpole Mass of Pole [5, 5, 4] 100 4.67 21.43
+BipedalWalker Gravity [3, 3, 9] 100 5 20
+BipedalWalker Friction [2, 4, 2] 100 2.67 37.5
+BipedalWalker Scale [3, 13, 1] 100 5.67 17.65
+HalfCheetah Gravity [2, 2, 2] 100 2 50
+HalfCheetah Friction [1, 3, 3] 100 2.33 42.86
+HalfCheetah Stiffness [1, 3, 3] 100 2.33 42.86
+AA-Ring-Acc Hold Duration [3, 3, 4] 40 3.33 12
+AA-Ring-Vel Hold Duration [1, 3, 5] 40 3 13.33
+AA-Ramp-Acc Hold Duration [32, 3, 4] 40 13 3.08
+AA-Ramp-Vel Hold Duration [3, 2, 2] 40 2.33 17.14
+Traffic Signal Road Length [19, 15, 10] 50 14.67 3.41
+Traffic Signal Inflow [3, 2, 2] 50 2.33 21.43
+Traffic Signal Speed Limit [13, 8, 5] 50 8.67 5.77
+Eco-Driving Penetration Rate [2, 2, 1] 50 1.67 30
+Eco-Driving Inflow [3, 1, 1] 50 1.67 30
+Eco-Driving Green Phase [2, 2, 3] 50 2.33 21.43
 A.4.5 Details about traffic signal control benchmark
@@
 Different configurations of intersections (e.g., road length, inflow, speed limits) are modeled to represent varying real-world conditions.
-To represent various real-world conditions, we vary factors such as road length, inflow rate, and speed limits from 0.1 to 5 times; by default, the road length is 750 meters, the flow rate is 1000 vehicles per hour, and the speed limit is 13.89 m/s.
+We vary factors such as road length, inflow rate, and speed limits from 0.1 to 5 times; by default, the road length is 750 meters, the flow rate is 500 vehicles per hour, and the speed limit is 13.89 m/s.
 Figure 13: Illustration of the traffic networks in traffic signal control task.
 Training configuration
-We used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0 and PPO for RL algorithm [36].
-We utilized the default implementation of the PPO algorithm with default hyperparameters provided by the Stable-Baselines3 library [34].
+We used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0.
+For reinforcement learning, the Deep Q-Network (DQN) algorithm was employed with a neural network architecture comprising four hidden layers, each with 400 units.
+The learning rate was set to 0.001, and training was conducted over 800 epochs.
+The discount factor ($\gamma$) was configured at 0.75 to balance short-term and long-term rewards effectively.
 All experiments are done on a distributed computing cluster equipped with 48 Intel Xeon Platinum 8260 CPUs.
@@
 The reward mechanism is designed to optimize the driving strategy by balancing the average speed of the vehicles against penalties for emissions, thereby promoting eco-friendly driving behaviors within the traffic system.
+The traffic simulation used a default inflow of 400 vehicles per hour, a CAV penetration rate of 0.2, and a green phase time of 35 seconds to simulate realistic urban traffic conditions, with parameters varied from 0.1 to 5 times for CMDP.
 Figure 17: Illustration of the traffic networks in eco-driving control task.
@@
 We also used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0 and PPO for RL algorithm [37].
-For detailed experimental details and RL hyperparameter configurations, please refer to [20, 21].
+The Proximal Policy Optimization (PPO) algorithm was configured with a policy clipping parameter of 0.03 and an initial KL divergence coefficient of 0.1, targeting a KL divergence of 0.02 during training.
+The value function was clipped with a value clip of 3, and the value loss coefficient was set to 1.
+Entropy regularization was applied with a coefficient of 0.005 to encourage exploration.
+Gradient updates were performed over 10 steps per epoch, with training spanning 5000 epochs and 10 episodes per epoch.
+Each episode had a horizon of 1500 steps, and mini-batches of size 40 were used.
+The Adam optimizer was employed with a learning rate of 0.0001, weight decay of 0.97, and betas set to (0.9, 0.999).
+A neural network with four hidden layers, each 256 units wide, used the tanh activation function and orthogonal weight initialization.
+The simulation warmup steps were set to 50, and the simulation step size was 0.5 seconds.
+The discount factor ($\gamma$) was 0.99.
+For detailed experimental details and RL hyperparameter configurations, please refer to [20].
 License
@@
 The highway ramp scenario introduces a more complex dynamic, where vehicles not only travel along the highway but also merge from ramps, creating potential stop-and-go traffic patterns that challenge the adaptability of autonomous guidance systems.
+The road network consists of a pre-merge distance of 400 m, a merge distance of 100 m, and a post-merge distance of 30 m.
+Traffic inflow rates were set to 2000 vehicles per hour on the highway and 300 vehicles per hour on the ramp.
 Problem Definition
-In a single-lane scenario, the state space includes the speeds of the ego and leading vehicles, along with the headway.
+In a single-lane ring scenario, the state space includes the speeds of the ego and leading vehicles, along with the headway.
+Vehicle dynamics incorporated acceleration and deceleration limits of 0.5 m/s2.
 For highway ramp scenarios, additional states cover the relative positions and speeds of adjacent vehicles.
@@
 We explore different durations of coarse-grained guidance holds to test various levels of human compatibility, adjusting the model based on observed driver behaviors and system performance.
+Training configuration
+Advisory autonomy experiments utilized Trust Region Policy Optimization (TRPO) [36], with a discount factor ($\gamma$) of 0.999, a learning rate of $10^{-3}$, and the Adam optimizer configured with weight decay of 0.97 and betas (0.9, 0.999).
+Policies were modeled with a four-layer neural network, each with 256 units and tanh activation, using orthogonal weight initialization.
+The KL divergence constraint was set to 0.02, with an initial KL coefficient of 0.1 dynamically adjusted during training.
+Rewards were normalized and centered, and regularization penalties were applied to ensure stable and robust policy optimization.
 License
@@
 The range of context variations was established by scaling the default values specified in the CARL framework from 0.1 to 10 times, enabling a comprehensive examination of each model’s performance under drastically different conditions.
-We utilized the default implementation of the DQN, A2C, and PPO algorithm with default hyperparameters provided by the Stable-Baselines3 library [34].
+Implementation Details.
+We employed the PPO algorithm from stable_baselines3 (v1.5.0) [34] with its default hyperparameters, including a learning rate of $3\times 10^{-4}$, n_steps$=2048$, batch size $64$, discount factor $\gamma=0.99$, GAE parameter $\lambda=0.95$, clipping parameter $0.2$, entropy coefficient 00, and a value function loss coefficient of $0.5$.
 A.4.9 Details about Cartpole benchmark
+Environment Details.
+We utilized the CARL benchmark library’s default environment parameters for the Cartpole task, training for five million total timesteps.
+Specifically, Cartpole used the default length of a pole of $0.5$, the mass of the cart of $1.0$, and the mass of a pole of $0.1$.
 Potential of multi-policy training and zero-shot transfer
@@
 A.4.10 Details about Pendulum benchmark
+Environment Details.
+We utilized the CARL benchmark library’s default environment parameters for the Pendulum task, training for a million total timesteps.
+Specifically, we used the default length of $1.0$, the mass of $1.0$, and the simulation timestop of $0.05$.
 Potential of multi-policy training and zero-shot transfer
@@
 A.4.11 Details about BipedalWalker benchmark
+Environment Details.
+We utilized the CARL benchmark library’s default environment parameters for the BipedalWalker task, training for five million total timesteps.
+Specifically, BipedalWalker used the default friction of $2.5$, scale of $30$, and gravity of $10$.
 Potential of multi-policy training and zero-shot transfer
@@
 A.4.12 Details about HalfCheetah benchmark
+Environment Details (HalfCheetah).
+We utilized the CARL benchmark library’s default environment parameters for the HalfCheetah task, training for five million total timesteps.
+Specifically, HalfCheetah used the default joint stiffness of $15000$, gravity of $9.8$, and friction of $0.6$.
 Potential of multi-policy training and zero-shot transfer